### v1

# Initialization using Update Approximation is a Silver Bullet for Extremely Efficient Low-Rank Fine-Tuning

Abstract
Low-rank adapters have become a standard approach for efficiently fine-tuning large language models (LLMs), but they often fall short of achieving the performance of full fine-tuning. We propose a method, LoRA Silver Bullet or LoRA-SB, that approximates full fine-tuning within low-rank subspaces using a carefully designed initialization strategy. We theoretically demonstrate that the architecture of LoRA-XS—which inserts a trainable $r\times r$ matrix between $B$ and $A$ while keeping other matrices fixed—provides the precise conditions needed for this approximation. We leverage its constrained update space to achieve optimal scaling for high-rank gradient updates while removing the need for hyperparameter tuning. We prove that our initialization offers an optimal low-rank approximation of the initial gradient and preserves update directions throughout training. Extensive experiments across mathematical reasoning, commonsense reasoning, and language understanding tasks demonstrate that our approach exceeds the performance of standard LoRA while using 27-90x fewer parameters, and comprehensively outperforms LoRA-XS. Our findings establish that it is possible to simulate full fine-tuning in low-rank subspaces, and achieve significant efficiency gains without sacrificing performance. We have released our code publicly at https://github.com/RaghavSinghal10/lora-sb.
Machine Learning, ICML
1 Introduction
Pre-trained language models have become central to natural language processing, achieving state-of-the-art performance across diverse tasks (Radford et al., 2021; Kirillov et al., 2023; Achiam et al., 2023). While these models excel at general-purpose capabilities (Bubeck et al., 2023; Hao et al., 2022), adapting them to specific downstream tasks often requires fine-tuning. Although in-context learning (Brown et al., 2020; Radford et al., 2019) has gained popularity for its simplicity, it falls short in both performance and efficiency compared to fine-tuning (Liu et al., 2022). At the same time, full fine-tuning, while highly effective, is computationally expensive and impractical at scale, highlighting the need for more efficient adaptation techniques.
Parameter-efficient fine-tuning (PEFT) has emerged as a crucial approach for adapting large language models (LLMs) while addressing computational constraints. While full fine-tuning (FFT) typically achieves optimal performance, it requires updating billions of parameters, making it computationally intensive and often impractical for most applications. Low-rank decomposition methods, beginning with LoRA (Hu et al., 2021), have shown particular promise by significantly reducing trainable parameters through learning low-rank updates. This has sparked numerous advances in low-rank methods that either enhance performance through better optimization techniques and initialization strategies, or improve parameter efficiency through structured matrices and adaptive rank selection (Zhang et al., 2023; Wang et al., 2024b, a). However, these methods still face fundamental trade-offs, they must either maintain a relatively large number of parameters to match FFT performance, or accept a performance degradation when pursuing extreme parameter efficiency (Hu et al., 2021; Ding et al., 2023; Wang et al., 2024b). This raises an important question: can we design low-rank methods that maintain FFT-competitive performance while drastically reducing the parameter count beyond current approaches?
Figure 1: LoRA-SB Illustration. LoRA-XS (Bałazy et al., 2024) reduces parameter count compared to LoRA (Hu et al., 2021) by inserting a trainable $r\times r$ matrix $R$ between $B$ and $A$, while keeping other matrices fixed, leading to $W=W_{0}+sBRA$. Our method, LoRA-SB, leverages the same architecture. We find that updating $R$ using its gradients $g^{R}$ is equivalent to updating the full-finetuning matrix $W$ with an equivalent gradient $\tilde{g}_{SB}=sBg^{R}A$. We initialize $B$, $R$, and AA such that the equivalent gradient $\tilde{g}_{SB}$ optimally approximates the full fine-tuning gradient $g$ in low rank subspaces at each training step. In essence, we simulate the entire full fine-tuning process optimally within low-rank subspaces by utilizing only the first gradient $g_{1}$ (shown in green) from full fine-tuning.
Low-rank decomposition methods operate on a fundamental premise: fine-tuning requires learning only a low-rank update to the pre-trained weights. Some theoretical work extends this hypothesis, suggesting that methods like LoRA can learn any low-rank approximation of the full fine-tuning gradient. However, the gradients computed by these methods to update their trainable adapters do not inherently possess this property. For instance, LoRA’s gradients need explicit optimization at each step to better approximate the full fine-tuning gradient (Wang et al., 2024b). Additionally, initialization has emerged as a critical factor in low-rank adaptation, as highlighted by recent works like Pissa-LoRA (Meng et al., 2024) and LoRA-GA (Wang et al., 2024a), which address this challenge through various low-rank approximation techniques.
We formally analyze these limitations in the context of the architecture used in LoRA-XS (Bałazy et al., 2024)—which inserts a trainable $r\times r$ matrix between $B$ and $A$ while keeping other matrices fixed—and demonstrate that these challenges are even more pronounced in its framework. While exploring solutions inspired by existing LoRA-based methods, we discover a remarkable property unique to LoRA-XS: through careful initialization of matrices $A$ and $B$, we can approximately simulate the full fine-tuning optimization in low rank subspaces throughout the entire training, as shown in Figure 1. Our initialization strategy provides optimal scaling for approximating high-rank full fine-tuning gradients and eliminates the need for the scaling hyperparameter $\alpha$, results which we prove theoretically. Our key contributions include:
  • We theoretically formalize the limitations of LoRA-XS, showing how its constrained update space leads to suboptimal gradient approximation, initialization sensitivity, and hyperparameter dependence.
  • We propose a principled initialization strategy derived from approximating the first step of full fine-tuning, proving it provides optimal low-rank approximation of the initial gradient and preserves update directions throughout training.
  • We prove that our initialization makes gradient optimization hyperparameter-independent and guarantees convergence by maintaining orthonormal bases, eliminating the need for any tuning of the scaling factor.
  • Through extensive experiments on $4$ models across $16$ datasets covering mathematical reasoning, commonsense reasoning, and language understanding tasks, we demonstrate that our method surpasses the performance of LoRA while using 27-90x less parameters, and comprehensively outperforms LoRA-XS.
2 Related Work
Parameter-Efficient Fine-Tuning. Parameter-Efficient Fine-Tuning (PEFT) methods have become essential for adapting large pre-trained models under computational constraints. Early techniques like AdapterFusion (Pfeiffer et al., 2021) and Prefix-Tuning (Li & Liang, 2021) enabled task-specific adaptation with minimal parameter updates. Advances like soft prompts (Lester et al., 2021) further reduced trainable parameter counts while maintaining strong performance. Recent approaches have explored operating directly on model representations (Wu et al., 2024), offering different trade-offs between efficiency and performance.
Low-Rank Decomposition Methods. LoRA (Hu et al., 2021) demonstrated that weight updates during fine-tuning could be efficiently approximated using low-rank matrices, drastically reducing parameter counts while freezing pre-trained weights. Building on this insight, variants such as QLoRA (Dettmers et al., 2023) and AdaLoRA (Zhang et al., 2023) extended the paradigm through quantization and adaptive allocation strategies. The applicability of low-rank techniques has also been explored in pretraining with GaLore (Zhao et al., 2024) and ReLoRA (Lialin et al., 2023), highlighting the versatility of low-rank adaptation in both fine-tuning and pre-training contexts. LoRA-based methods have also been applied in other domains, such as efficient federated fine-tuning (Sun et al., 2024; Singhal et al., 2024).
Enhancing LoRA Performance. Recent efforts have focused on optimizing LoRA’s performance. PiSSA-LoRA (Meng et al., 2024) demonstrated faster convergence and improved task performance by initializing matrices with principal components of pre-trained weights. LoRA-Pro (Wang et al., 2024b) and LoRA-GA (Wang et al., 2024a) improved gradient approximation, aligning low-rank updates more closely with full fine-tuning. Other methods like DoRA (Liu et al., 2024) and rsLoRA (Kalajdzievski, 2023) introduced decomposition-based and scaling stabilization techniques to enhance learning stability and expand LoRA’s utility.
Improving Efficiency in LoRA Variants. Efficiency-focused innovations have pushed LoRA toward even greater parameter savings. LoRA-XS (Bałazy et al., 2024) achieves parameter reduction by inserting a small trainable weight matrix into frozen low-rank matrices. VeRA (Kopiczko et al., 2024) shares low-rank matrices across layers, relying on scaling vectors for task-specific adaptation. Tied-LoRA (Renduchintala et al., 2024) leverages weight tying to reduce parameter usage at higher ranks, while HydraLoRA (Tian et al., 2024) introduces an asymmetric architecture to improve adaptation in complex domains without domain-specific expertise.
3 Methodology
3.1 Preliminaries
In standard fine-tuning, a pre-trained weight matrix $W\in\mathbb{R}^{m\times n}$ is updated as:
$W=W_{0}+\Delta W$ (1)
where $W_{0}$ is the pre-trained weight. This requires updating $mn$ parameters per layer.
LoRA hypothesizes that these updates lie in a low-dimensional subspace and parameterizes the update matrix $\Delta W$ as:
$W=W_{0}+sBA$ (2)
where $B\in\mathbb{R}^{m\times r}$ and $A\in\mathbb{R}^{r\times n}$ are trainable low-rank matrices, with rank $r\ll\min(m,n)$, and $s$ is a scaling factor to stabilize training. This reduces parameters from $mn$ to $r(m+n)$.
LoRA-XS introduces a more efficient parameterization:
$W=W_{0}+sBRA$ (3)
where $B\in\mathbb{R}^{m\times r}$ and $A\in\mathbb{R}^{r\times n}$ are fixed matrices, and only $R\in\mathbb{R}^{r\times r}$ is trainable, reducing parameters to $r^{2}$.
We denote full fine-tuning gradient as $g=\frac{\partial L}{\partial W}$, and LoRA-XS gradient as $g^{R}_{\text{LoRA-XS}}=\frac{\partial L}{\partial R}$.
3.2 Motivation
LoRA-XS (Bałazy et al., 2024), although having significantly fewer learnable parameters compared to LoRA, exhibits suboptimal performance. The difference in architecture causes constraints on the type of updates LoRA-XS can learn. The subspace of learned updates is characterized in Lemma 1.
Lemma 1.
Let $\Delta W$ be an update learned by LoRA-XS . Then, the set of all $\Delta W$, say $\mathcal{W}_{LoRA-XS}$, can be given by $\{M\in\mathbb{R}^{m\times n}|\text{Col}(W)\subseteq\text{Col}(M)$  $\wedge\text{ Row}(M)\subseteq\text{Row}(A)\}$
Proof.
Appendix A.2
The above lemma implies that while $\Delta W$ is constrained to be $r$ ranked or lower, it also needs to have row and column spaces defined by those of $B$ and $A$ respectively. As against this, LoRA can learn any update $\Delta W$ as long as $\text{rank}(\Delta W)\leq r$. Thus the low expressivity of LoRA-XS as compared to LoRA can account for the performance drop. We identify three key limitations which arise due to this and otherwise:
1) Inadequate Gradient Approximation: LoRA optimization is mathematically equivalent to full fine-tuning using a constrained low-rank gradient. The gradient of LoRA does not optimally approximate the full gradient, and needs to be tuned at each step. LoRA-Pro (Wang et al., 2024b) finds that this results in suboptimal performances, and provides a closed form solution to optimize the gradients. In LoRA-XS , the gradient updates are restricted to an even more constrained low-rank space since the matrices $A$ and $B$ are fixed. We posit that the limitation becomes particularly severe when the ideal updates lie outside the space spanned by fixed $A$ and $B$, and consequently has a larger impact on performance.
2) Suboptimal Initialization: While initialization impacts all low-rank methods, it becomes critical in LoRA-XS where $A$ and $B$ are frozen. Unlike LoRA where poor initialization can be compensated through training, LoRA-XS must rely entirely on its initial subspace defined by $A$ and $B$. Consider the zero initialization of the $B$ matrix, for example. While LoRA may experience some performance degradation in this case (Wang et al., 2024a) (Meng et al., 2024), the ideal low-rank update $\Delta W$ can still be reached through gradient descent. In fact, zero initialization for the $B$ matrix is a commonly used choice, including in the original LoRA implementation (Hu et al., 2021). However, in the case of LoRA-XS , this would result in no learning, as the product $BRA$ would remain zero. The current LoRA-XS method leverages the most significant subspaces spanned by the columns of pre-trained weights $W_{0}$, inspired by (Meng et al., 2024). This initialization is inadequate because it fails to capture the specific subspaces relevant to the fine-tuning task. This issue, known to affect LoRA ’s performance (Wang et al., 2024a), could have an even greater impact on LoRA-XS for the reasons discussed above.
3) Hyperparameter Sensitivity: The scaling factor $\alpha$, present in almost every LoRA based fine-tuning method requires careful tuning to maintain stability during training. This hyperparameter acts as a bridge between the low-rank and full-rank spaces, compensating for the dimensional mismatch in gradients. Poor tuning of $\alpha$ can lead to unstable training or slow convergence (Kalajdzievski, 2023), adding complexity to the adaptation process and potentially limiting practical deployment.
3.3 Approximation of the full fine-tuning gradient
As mentioned above, LoRA optimization is mathematically equivalent to full fine-tuning using a constrained low-rank gradient. However the update generated using the gradients of LoRA does not result in the same update which the low-rank gradient would have generated. The following holds true for LoRA-XS as well.
To understand this, let us look at the change in weight $W$ and its relationship with changing of low-rank matrix $R$. The relationship can be simply given by $dW=\frac{\partial W}{\partial R}g^{R}$. This relationship implies that updating matrix $R$ with gradient $g^{R}$ is equivalent to updating $W$ with low rank equivalent gradient $\tilde{g}$ in full fine-tuning as described in Definition 1.
Definition 1 (Equivalent Gradient).
In the context of LoRA-XS optimization, we define the equivalent gradient as: $\tilde{g}=sBg^{R}A$ where $s$ is the scaling factor, and $g^{R}$ is the gradient with respect to matrix $R$.
The equivalent gradient describes the virtual low-rank gradient of matrix $W$ in LoRA-XS optimization process, despite $W$ not being directly trainable. This gradient determines how updates to $R$ affect the overall weight matrix. To bridge the performance gap between LoRA-XS and full fine-tuning, our goal is to minimize the discrepancy between the equivalent gradient $\tilde{g}$ and the full gradient $g$. First, we establish the relationship between gradients in LoRA-XS optimization in Lemma 2.
Lemma 2.
During LoRA-XS optimization, the gradient of the loss with respect to matrix $R$ can be expressed in terms of the gradient with respect to the weight matrix $W$ as: $g^{R}_{LoRA-XS}=sB^{T}gA^{T}$
Proof.
Appendix A.2
Leveraging this relationship, we can formulate our objective to minimize the distance between the equivalent gradient and the full gradient. We do not have access to the full fine-tuning gradient $g$ during LoRA-XS based fine-tuning. Thus we need to find the ideal gradient with respect to $R$, given by $g^{R}$, and subsequently the optimal approximation $\tilde{g}$, in terms of the gradient which is available to us during training: $g^{R}_{LoRA-XS}$. Fortunately, this optimization problem admits a closed-form solution independent of $g$ as described in Theorem 3.
Theorem 3.
[…]
Proof.
Appendix A.3. ∎
[…]
Theorem 4.
Consider the update for matrix $R$ using the closed-form solution derived in Theorem 3: $R\leftarrow R-\eta g^{R}$ where $\eta\geq 0$ is the learning rate. This update guarantees a reduction in the loss function, similar to traditional gradient descent methods. Specifically, the change in loss $\mathrm{d}L$ is given by: $\mathrm{d}L=-\eta\langle g^{R}_{LoRA-XS},g^{R}\rangle\leq 0$
Proof.
Appendix A.4. ∎
3.4 Initialization using update approximation
Let us first consider the objective in fine-tuning, the primary goal is to update weights to better suit the target task and domain. The initial gradient steps are particularly informative, as they indicate the direction of desired adaptation. We propose leveraging this insight by using the first update step from full fine-tuning to initialize our low-rank matrices.
This approach offers two key advantages: First, it ensures the low-rank space captures the most relevant subspace for the target task rather than relying on pre-trained weight properties. Second, since $A$ and $B$ matrices will remain fixed in LoRA-XS, initializing them to span the subspace of early adaptation increases the likelihood of capturing useful updates throughout training. This would also ensure that the final update is learnt in the correct subspace, of which we have no apriori information besides the first step of full fine-tuning.
In essence, our method can be summarized as follows: Set an initialization which best approximates the first step of full fine-tuning. Formally, given a full fine-tuning update $\Delta W$, our initialization should satisfy:
$sB_{init}R_{init}A_{init}\approx\Delta W_{first-step}$ (5)
The first step of full fine-tuning, for Adam-based optimizers such as AdamW, for a sample $x_{i}$ is:
$\Delta W_{first-step}=-\eta\times\textbf{sign}(\nabla_{W}\mathcal{L}(W_{0},x_{i}))$ (6)
Using a single sample may lead to noisy estimates. Instead, we compute a more stable initialization by averaging gradients over a subset of the training data:
$\Delta W_{avg}=-\eta\textbf{sign}(\sum_{i=0}^{n\leq|\mathbb{X}|}\nabla_{W}\mathcal{L}(W_{0},x_{i}))\dots x_{i}\in\mathbb{X}$ (7)
This averaged gradient better captures the general direction of adaptation required for the target task while being less sensitive to individual sample variations. We can then use truncated SVD to obtain a low-rank approximation of $\Delta W_{\text{avg}}$, and express it as $sBRA$. There exist infinite combinations of $B$ and $A$ which can obey the above relationship. For instance, we can initialize $B$ and $A$ as $US$ and $V^{T}$ and keep $R$ as $I/s$. This initialization is equivalent to the $B$ and $A$ initialization in LoRA-XS but by approximating the update rather than the pre-trained matrix. We note that the above process can be computed for any optimizer, by approximating the corresponding first step. We compute this specifically for AdamW since we use it.
3.5 Hyperparameter independence
The hyperparameter $\alpha$ is used in LoRA and subsequent decomposition based models to tackle the issue of instability caused to improper scaling of the updates. The gradient scaling is accounted for, by adding a hyperparameter to normalize the updates. The importance of scaling is well documented in methods like rank stabilization (Kalajdzievski, 2023). However, the full fine-tuning gradient $g$ needs no such tuning. We claim that approximating the full fine-tuning gradient removes the need for introducing a scaling factor.
Theorem 5.
The equivalent gradient $\tilde{g}$ is hyperparameter $s$ independent for $\tilde{g}=sBg^{R}A$, but not for $\tilde{g}=sBg^{R}_{LoRA-XS}A$
Proof.
Appendix A.5
The hyperparameter independence of the equivalent gradient eliminates the need for manual gradient scaling. Updates to weight matrix $W$ depend solely on this gradient (modulo learning rate), making any additional scaling redundant.
This can be understood by examining the relationship with full fine-tuning gradient $g$. Since $g$ is naturally scaled for optimal weight updates, and our method approximates $g$ in a constrained subspace, the equivalent gradient inherits appropriate scaling automatically. Notably, this property is unique to our gradient approximation approach and does not hold for standard LoRA-XS training.
3.6 LoRA-SB: Update approximation initialization is a silver bullet
The solutions discussed above independently address the gradient approximation and initialization problems, while also providing hyperparameter independence. Our proposed method, LoRA-SB, elegantly combines these solutions through a simple initialization strategy.
Our initialization strategy is derived from approximating the first step of full fine-tuning:
$U,S,V^{T}\leftarrow\textbf{SVD}(\Delta W_{avg})$ (8)
$B_{init}\leftarrow U[1:r]$ (9)
$A_{init}\leftarrow V[1:r]$ (10)
$R_{init}\leftarrow\dfrac{1}{s}S[1:r,1:r]$ (11)
By the Eckart-Young theorem (Eckart & Young, 1936; Mirsky, 1960), this gives the optimal rank-$r$ approximation of the initial full fine-tuning update. where $U$, $S$, $V$ are obtained from truncated SVD of the averaged first update $\Delta W_{\text{avg}}$.
This initialization leads to several key advantages that address the problems identified earlier.
Simplified Gradient Optimization. This initialization ensures $B_{\text{init}}$ and $A_{\text{init}}$ form orthonormal bases in $\mathbb{R}^{m}$ and $\mathbb{R}^{n}$ respectively, leading to $B^{T}B=AA^{T}=I$. With fixed $B$ and $A$ matrices being orthonormal, the optimal update step derived in Equation 3 simplifies to:
$g^{R}=\dfrac{1}{s^{2}}(B^{T}B)^{-1}g^{R}_{LoRA-XS}(AA^{T})^{-1}=\frac{1}{s^{2}}g^{R}_{LoRA-XS}$ (12)
This eliminates the need for complex matrix inversions during training.
Optimum Update Approximation. Our initialization guarantees that the first update optimally approximates the full fine-tuning weight updates:
$sB_{\text{init}}R_{\text{init}}A_{\text{init}}\approx\Delta W_{avg}$ (13)
By the Eckart-Young theorem, this gives the optimal rank-$r$ approximation of the initial full fine-tuning update.
Hyperparameter Independence. As shown in Theorem 5, when gradient approximation is applied with orthonormal $B$ and $A$, the hyperparameter $s$ can be set to 1, resulting in the following:
$\boxed{g^{R}=g^{R}_{\text{LoRA-XS}}}$ (14)
This demonstrates that our initialization guarantees optimal gradient approximation at every step, without requiring any scaling factor!
Guaranteed Loss Reduction. Since $B$ is a tall orthonormal matrix and $A$ is a wide orthonormal matrix, they remain full rank throughout training. This ensures that $dL$ remains negative 4, guaranteeing stable optimization and convergence.
Another heuristic which might lead to a good initialization is setting the weights $B$ and $A$, such that they match the first update also approximately matches the direction of $\Delta W$.
$\Delta(sB_{init}R_{init}A_{init})\approx\gamma\Delta W$ (15)
Thankfully, we don’t have to choose between the two. For SGD, we prove that setting $B_{init}$ and $A_{init}$ using Equations 8-11, results in the first update of LoRA-XS to best approximate the direction of the update of full fine-tuning.
Theorem 6.
If $A_{init}$ and $B_{init}$ are initialized using LoRA-SB for the first SGD optimizer, then $\Delta(B_{init}R_{init}A_{init})\approx\Delta W$
Proof.
Appendix A.6
LoRA-SB Advantages over LoRA. Notably, many of the properties described above are not achievable with standard LoRA methods. Even if $B$ and $A$ are initialized as orthonormal in LoRA, subsequent updates do not preserve this property because $B$ and $A$ are trainable. This results in several challenges:
  • Potential instability of $(B^{T}B)^{-1}$ and $(AA^{T})^{-1}$, as they are not guaranteed to remain non-singular during training.
  • Inability to ensure consistent loss reduction due to potential rank deficiency—$B$ and $A$ may not remain full-rank throughout training.
  • Necessity to fine-tune the hyperparameter $\alpha$.
  • Repeated re-computation of $B^{T}B$ and $AA^{T}$ is required at each optimizer step for accurate gradient approximation.
LoRA-SB’s simple initialization strategy solves multiple problems simultaneously, offering a robust and efficient approach to low-rank adaptation.
Algorithm. We present a PyTorch-like implementation of our method in Algorithm 1. To optimize GPU memory usage during initialization, we hook into the backward pass of PyTorch and compute the gradient for one layer at a time, immediately discarding the computed gradients (Lv et al., 2024; Wang et al., 2024a). This ensures that memory usage remains at $O(1)$, independent of the number of layers. This approach keeps memory consumption under control, ensuring it never exceeds the requirements of subsequent LoRA-SB fine-tuning. For very large batch sizes, one can also use gradient accumulation and quantization to further manage memory usage within the specified limits.
Algorithm 1 LoRA-SB, PyTorch-like
def initSB(model, D)
   # Estimate gradient with n samples
   $\mathbf{\Delta W_{avg}}\leftarrow$ est_grad(model, D, n)
   # Initialize B, R, A
   $\mathbf{B,R,A}\leftarrow$ trunc_SVD($\mathbf{\Delta W_{avg}}$)
   # Convert to LoRA-SB model
   $\text{{sb\_model}}\leftarrow$ lora_SB(model, B, R, A)
   return sb_model
# Load pre-trained model
$\text{{model}}\leftarrow$ AutoModel(base_model)
# Initialize LoRA-SB with dataset D
$\text{{sb\_model}}\leftarrow$ initSB(model, D)
# Standard training, only R trainable
$\text{{trainer}}\leftarrow$ Trainer(sb_model, ...)
$\text{{trainer}}.$train()
4 Experiments
We evaluate our method over $16$ different datasets on three widely-used NLP benchmarks, using models ranging from the 355 M-parameter RoBERTa-large model to the 9 B-parameter Gemma-2 model. Our setup spans both masked and autoregressive architectures, allowing us to comprehensively assess the effectiveness of our approach. We fine-tune RoBERTa-large (Liu et al., 2019), Llama-3.2 3B (Dubey et al., 2024), Mistral-7B (Jiang et al., 2023), and Gemma-2 9B (Team et al., 2024), showcasing our method’s adaptability across a variety of tasks and model architectures.
We implement all models using PyTorch (Paszke et al., 2019) and the popular HuggingFace Transformers library (Wolf et al., 2020). We run all experiments on a single NVIDIA A6000 GPU and report results as the average of three random seeds. To save memory, we initialize base models in torch.bfloat16 precision. Appendix D provides detailed information on the datasets used.
We compute the update approximation using only $\mathbf{1/1000}$ of each dataset’s total number of samples. This ensures that the additional training time overhead is minimal and has a negligible effect on overall efficiency. We provide a benchmark of this overhead in Section 5. These samples are randomly selected from the dataset’s training set in each run.
4.1 Arithmetic Reasoning
Details. We fine-tune Mistral-7B (Jiang et al., 2023) and Gemma-2 9B (Team et al., 2024) on 50K samples from the MetaMathQA (Yu et al., 2024) dataset and evaluate them on the GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) benchmarks. Evaluation focuses solely on the final numeric answer. We apply LoRA modules to the key, value, query, attention output, and all fully connected weight matrices, training with ranks $r=\{32,64,96\}$. Detailed hyperparameter settings are provided in Appendix C.
Results. We present the results for both models in Table 1. LoRA-SB significantly outperforms LoRA-XS across all parameter settings. Notably, LoRA-SB surpasses the performance of LoRA ($r=32$) while using 40x fewer trainable parameters for Mistral-7B and 90x fewer for Gemma-2 9B at ranks $r=96$ and $r=64$, respectively.
We present training loss curves comparing LoRA-SB and LoRA-XS in Figure 2. Thanks to superior initialization, LoRA-SB starts with a lower initial loss compared to LoRA-XS. Additionally, due to optimal gradient approximation, LoRA-SB maintains a consistently better loss curve throughout training and converges to a superior final value.
Table 1: Accuracy comparison of fine-tuning methods on Mistral-7B and Gemma-2 9B across the arithmetic reasoning benchmarks GSM8K and MATH, after training on MetaMathQA. # Params denotes the number of trainable parameters. The best results among PEFT methods are highlighted in bold.
Model Method Rank # Params Accuracy ($\uparrow$)
GSM8K MATH
Mistral-7B Full FT - $7.24$ B $63.87$ $17.65$
LoRA $32$ $83.88$ M $61.94$ $15.98$
LoRA-XS $32$ $0.23$ M $54.28$ $13.36$
LoRA-XS $64$ $0.92$ M $57.08$ $15.62$
LoRA-XS $96$ $2.06$ M $58.53$ $16.42$
LoRA-SB $32$ $0.23$ M $58.91$ $15.28$
LoRA-SB $64$ $0.92$ M $60.73$ $16.28$
LoRA-SB $96$ $2.06$ M $\mathbf{63.38}$ $\mathbf{17.44}$
Gemma-2 9B Full FT - $9.24$ B $79.23$ $38.02$
LoRA $32$ $108.04$ M $76.19$ $36.56$
LoRA-XS $32$ $0.30$ M $74.07$ $34.62$
LoRA-XS $64$ $1.20$ M $75.02$ $36.46$
LoRA-XS $96$ $2.71$ M $75.21$ $36.98$
LoRA-SB $32$ $0.30$ M $75.44$ $36.66$
LoRA-SB $64$ $1.20$ M $76.65$ $37.14$
LoRA-SB $96$ $2.71$ M $\mathbf{78.40}$ $\mathbf{37.70}$
(a) Mistral-7B
(b) Gemma-2 9B
Figure 2: Training loss curves for Mistral-7B and Gemma-2 9B on MetaMathQA, comparing LoRA-SB and LoRA-XS.
4.2 Commonsense Reasoning
Details. We fine-tune Llama-3.2 3B (Dubey et al., 2024) on CommonSense170K, a dataset that aggregates eight commonsense reasoning tasks (Hu et al., 2023). We evaluate the model’s performance on each dataset individually, which include BoolQ (Clark et al., 2019), SIQA (Sap et al., 2019), PIQA (Bisk et al., 2020), ARC-Challenge (Clark et al., 2018), ARC-Easy (Clark et al., 2018), OBQA (Mihaylov et al., 2018), WinoGrande (Sakaguchi et al., 2021), and HellaSwag (Zellers et al., 2019). Each example is framed as a multiple-choice question where the model generates the correct answer without explanations. We use the prompt template from (Hu et al., 2023). LoRA modules are applied to the key, value, query, attention output, and all fully connected weight matrices, training with ranks $r=\{32,64,96\}$. Hyperparameter details are provided in Appendix C.
Results. We present the results in Table 2. LoRA-SB consistently outperforms LoRA-XS across all parameter settings. Additionally, LoRA-SB ($r=96$) exceeds the performance of LoRA ($r=32$) with 27x fewer trainable parameters.
Table 2: Accuracy comparison of fine-tuning methods on Llama-3.2 3B across eight commonsense reasoning datasets. # Params denotes the number of trainable parameters. The best results among PEFT methods are highlighted in bold.
Model Method Rank # Params Accuracy ($\uparrow$)
BoolQ PIQA SIQA HellaS. WinoG. ARC-e ARC-c OBQA Avg.
Llama-3.2 3B Full FT - $3.21$ B $70.43$ $85.64$ $80.45$ $91.92$ $85.02$ $88.52$ $75.29$ $81.88$ $82.39$
LoRA $32$ $48.63$ M $70.03$ $\mathbf{85.20}$ $79.12$ $90.71$ $82.24$ $86.91$ $74.32$ $\mathbf{81.87}$ $81.44$
LoRA-XS $32$ $0.20$ M $65.01$ $82.87$ $76.17$ $87.32$ $80.12$ $84.78$ $70.31$ $75.71$ $77.79$
LoRA-XS $64$ $0.80$ M $66.53$ $83.12$ $77.98$ $88.53$ $81.76$ $85.15$ $72.04$ $77.14$ $79.03$
LoRA-XS $96$ $1.81$ M $67.28$ $83.35$ $78.66$ $88.99$ $82.08$ $85.18$ $72.61$ $78.88$ $79.63$
LoRA-SB $32$ $0.20$ M $66.33$ $84.06$ $78.91$ $89.04$ $81.37$ $86.62$ $72.44$ $76.97$ $79.47$
LoRA-SB $64$ $0.80$ M $68.35$ $84.55$ $79.94$ $\mathbf{91.68}$ $83.03$ $87.84$ $74.83$ $80.12$ $81.29$
LoRA-SB $96$ $1.81$ M $\mathbf{70.34}$ $84.76$ $\mathbf{80.19}$ $91.62$ $\mathbf{84.61}$ $\mathbf{87.92}$ $\mathbf{74.74}$ $81.20$ $\mathbf{81.92}$
4.3 Natural Language Understanding
Details. We fine-tune RoBERTa-large (Liu et al., 2019) on GLUE, a sequence classification benchmark that contains several datasets, covering domains like sentiment analysis and natural language inference. The datasets we evaluate on are: CoLA (Warstadt et al., 2019), RTE, MRPC (Dolan & Brockett, 2005), SST-2 (Socher et al., 2013), QNLI (Rajpurkar et al., 2018), and STS-B (Cer et al., 2017). LoRA modules are applied only to the self-attention layers, following the configuration in the original LoRA paper (Hu et al., 2021), with ranks $r=\{8,16,24\}$. Detailed experimental settings are available in Appendix C.
Results. The results are shown in Table 3. LoRA-SB consistently outperforms LoRA-XS across all parameter configurations. Notably, LoRA-SB ($r=24$) outperforms LoRA ($r=8$) with 39x lesser trainable parameters.
Table 3: Comparison of fine-tuning methods on RoBERTa-large across the GLUE benchmark datasets. # Params denotes the number of trainable parameters. The best results among PEFT methods are highlighted in bold. We use Pearson correlation for STS-B, Matthew’s correlation for CoLA, and accuracy for others. Higher is better for each metric.
Model Method Rank # Params CoLA RTE MRPC STS-B QNLI SST-2 All
Mcc $\uparrow$ Acc $\uparrow$ Acc $\uparrow$ Corr $\uparrow$ Acc $\uparrow$ Acc $\uparrow$ Avg. $\uparrow$
RoBERTa-large Full FT - $355.36$ M $68.44$ $83.42$ $90.21$ $91.76$ $93.92$ $96.21$ $87.33$
LoRA 88 $2162.69$ K $68.02$ $82.98$ $90.05$ $91.43$ $93.42$ $95.98$ $86.98$
LoRA-XS 88 $6.14$ K $61.07$ $75.23$ $86.21$ $89.29$ $92.44$ $94.72$ $83.16$
LoRA-XS $16$ $24.57$ K $63.32$ $79.06$ $86.28$ $90.36$ $93.89$ $95.76$ $84.70$
LoRA-XS $24$ $55.20$ K $66.27$ $80.14$ $88.48$ $90.77$ $93.21$ $95.89$ $85.79$
LoRA-SB 88 $6.14$ K $63.57$ $78.43$ $88.72$ $90.59$ $92.95$ $95.07$ $84.88$
LoRA-SB $16$ $24.57$ K $64.36$ $82.31$ $89.71$ $91.24$ $93.89$ $95.87$ $86.23$
LoRA-SB $24$ $55.20$ K $\mathbf{68.28}$ $\mathbf{83.03}$ $\mathbf{90.12}$ $\mathbf{91.65}$ $\mathbf{93.75}$ $\mathbf{96.11}$ $\mathbf{87.16}$
5 Analysis
Optimal Initialization is Important!
To isolate the impact of initialization on training, we perform truncated SVD on various matrices, including Kaiming initialization and $\Delta W_{avg}$ with varying levels of added Gaussian noise, as shown in Table 4. By applying truncated SVD, we ensure optimal gradient approximation, leading to initialization matrices $B_{\text{init}}$ and $A_{\text{init}}$ that form orthonormal bases in $\mathbb{R}^{m}$ and $\mathbb{R}^{n}$, respectively. This results in $B^{T}B=AA^{T}=I$, allowing us to isolate the effect of initialization. The results clearly demonstrate the significance of initialization—our approach consistently outperforms other variants.
Table 4: Comparison of initialization strategies using Mistral-7B on GSM8K and MATH. All methods ensure optimal gradient approximation, with differences arising solely from the initialization.
Initialization Method Accuracy ($\uparrow$)
GSM8K MATH
trunc_SVD (Kaiming) $00.00$ $00.00$
trunc_SVD ($\Delta W_{avg}+\mathcal{N}_{\mu=10^{-2}}$) $00.00$ $00.00$
trunc_SVD ($\Delta W_{avg}+\mathcal{N}_{\mu=10^{-3}}$) $58.83$ $14.76$
trunc_SVD ($\Delta W_{avg}+\mathcal{N}_{\mu=10^{-4}}$) $60.19$ $15.96$
trunc_SVD ($\Delta W_{avg}+\mathcal{N}_{\mu=10^{-5}}$) $60.65$ $15.98$
LoRA-SB; trunc_SVD ($\Delta W_{avg}$) $\mathbf{63.38}$ $\mathbf{17.44}$
Optimal Gradient Approximation is Important!
We aim to examine the effect of optimal gradient approximation on training. Specifically, we want $B_{\text{init}}R_{\text{init}}A_{\text{init}}\approx\Delta W_{avg}$ without enforcing $B^{T}B=AA^{T}=I$. We achieve this through the following steps:
$U,S,V^{T}\leftarrow\textbf{SVD}(\Delta W_{avg})$ (16)
$B_{\text{init}}\leftarrow U[1:r]S[1:r,1:r]$ (17)
$A_{\text{init}}\leftarrow V[1:r]$ (18)
$R_{\text{init}}\leftarrow I$ (19)
This construction ensures that $B_{\text{init}}R_{\text{init}}A_{\text{init}}\approx\Delta W_{avg}$, but only $AA^{T}=I$, while $B^{T}B\neq I$. This setup is suboptimal for gradient approximation since we do not explicity use the closed-form solution derived in Theorem 3.
We compare the resulting loss curves against LoRA-SB (which uses optimal gradient approximation) for Mistral-7B on MetaMathQA, as shown in Figure 3. Although both methods start similarly due to effective initialization, LoRA-SB converges to significantly better values, demonstrating the advantage of optimal gradient approximation. Furthermore, LoRA-SB achieves higher accuracies on the GSM8K and MATH benchmarks, with scores of $63.38$ and $17.44$ compared to $55.87$ and $12.74$, respectively.
Figure 3: Training loss for Mistral-7B on MetaMathQA, highlighting the impact of optimal gradient approximation.
Training Time Overhead vs LoRA-XS
As previously mentioned, we compute the update approximation using only $1/1000$ of the total training samples for each dataset. Table 5 presents the associated training time overhead for these computations, compared to LoRA-XS. The results show that the additional overhead is negligible, adding just 242–4 minutes compared to the total training time of 353–5 hours per epoch ($\approx 1.1\%$ to $1.3\%$). Additionally, the update computation is performed only once, at the beginning of the first epoch, prior to training.
Table 5: Training time overhead due to the initialization for various models on their respective tasks.
Model Overhead Training Time/Epoch
Mistral 7B 0:02:01 3:03:57
Gemma-2 9B 0:03:46 4:13:24
Llama-3.2 3B 0:03:54 4:54:31
Inference Overhead vs LoRA
LoRA-SB introduces a minimal inference cost overhead due to the insertion of the $r\times r$ matrix $R$ between $B$ and $A$, and the need for higher ranks to achieve comparable performance to LoRA. We benchmark the inference-time FLOPs and MACs across various models and find that the overhead is negligible. A detailed analysis can be found in Table 6 of Appendix B.
6 Conclusion
In this work, we introduced LoRA-SB, a novel framework that bridges the gap between low-rank parameter-efficient fine-tuning and full fine-tuning. This is enabled by our carefully designed initialization strategy, which approximates the first step of full fine-tuning and ensures that the most relevant subspaces for task-specific adaptation are captured from the outset. Our theoretical analysis uncovered critical limitations in LoRA-XS: suboptimal gradient approximation and reliance on fixed initialization that fails to adapt to task-specific nuances. By addressing both issues through a unified approach, we demonstrate that LoRA-SB achieves optimal scaling and preserves critical update directions. Furthermore, our approach ensures hyperparameter independence by directly approximating the full fine-tuning gradient, thereby eliminating instability and convergence issues associated with scaling factors.
Through extensive experiments on 4 models across 16 datasets covering mathematical reasoning, commonsense reasoning, and language understanding tasks, we demonstrate that our method exceeds the performance of LoRA while using upto 90x less parameters, and comprehensively outperforms LoRA-XS. Our work advances the state of PEFT while laying the groundwork for further innovations in low-rank adaptations for neural networks. Future work includes exploring adaptive layer-wise rank settings and evaluating the approach on other architectures, such as Vision Language Models (VLMs) and Vision Transformers (ViTs).
7 Acknowledgements
This research was supported by funding from Mohamed bin Zayed University of Artificial Intelligence (MBZUAI) and ADIA Lab.
Appendix A Mathematical Proofs
In all the proofs below, we will use the notations defined in Section 3.
A.1 Proof of Lemma 1
Lemma.
Let $\Delta W$ be an update learned by LoRA-XS . Then, the set of all $\Delta W$, say $\mathcal{W}_{LoRA-XS}$, can be given by $\mathcal{W}_{LoRA-XS}:=\{M\in\mathbb{R}^{m\times n}|\text{Col}(M)\subseteq\text{Col}(B)\wedge\text{ Row}(M)\subseteq\text{Row}(A)\}$
Proof.
The column space of $\Delta W$ is given by:
$\text{Col}(\Delta W)$ $=\{y\in\mathbb{R}^{m}\mid y=\Delta Wx,\,x\in\mathbb{R}^{n}\}$ (20)
Now since $\Delta W=BRA$:
$\text{Col}(\Delta W)=\{y\in\mathbb{R}^{m}\mid y=BRAx,\,x\in\mathbb{R}^{n}\}\implies\text{Col}(\Delta W)=\{y\in\mathbb{R}^{m}\mid y=Bz,\,z\in\text{Col}(RA)\}\dots\text{ for }z=RAx$ (21)
Now:
$\text{Col}(B)$ $=\{y\in\mathbb{R}^{m}\mid y=Bx,\,x\in\mathbb{R}^{r}\}$ (22)
Now since $\text{Col}(RA)\subseteq\mathbb{R}^{r}$:
$\text{Col}(\Delta W)$ $\subseteq\text{Col}(B)$ (23)
A symmetric proof can now be given for row space of $\Delta W$. ∎
A.2 Proof of Lemma 2
Lemma.
During LoRA-XS optimization, the gradient of the loss with respect to matrix $R$ can be expressed in terms of the gradient with respect to the weight matrix $W$ as: $g^{R}_{LoRA-XS}=sB^{T}gA^{T}$
Proof.
Let $L$ be the loss function. We have already defined $g$ and $g^{R}_{\text{LoRA-XS}}$ as:
$g:=\frac{\partial L}{\partial W}\quad\&\quad g^{R}_{\text{LoRA-XS}}:=\frac{\partial L}{\partial R}$ (24)
Now by chain rule:
$\frac{\partial L}{\partial R}=\frac{\partial L}{\partial W}\frac{\partial W}{\partial R}\implies\frac{\partial L}{\partial R}=\frac{\partial L}{\partial W}\frac{\partial W}{\partial X}\frac{\partial X}{\partial R}\dots\text{ for }X=RA$ (25)
We know that for $W=sBX$:
$\frac{\partial L}{\partial W}\frac{\partial W}{\partial X}=sB^{T}g\implies\frac{\partial L}{\partial R}=sB^{T}g\frac{\partial X}{\partial R}$ (26)
Let $sB^{T}g=y$. We know that when $X=RA$:
$y\frac{\partial X}{\partial R}=yA^{T}\implies\frac{\partial L}{\partial R}=yA^{T}=sB^{T}gA^{T}\$(27)
$\text{Therefore, \quad}\boxed{g^{R}_{\text{LoRA-XS}}=sB^{T}gA^{T}}$ (28)
A.3 Proof of Theorem 3
Theorem.
[…]
Proof.
Since we already defined the equivalent gradient $\tilde{g}:=sBg^{R}A$, the minimization problem can be formally denoted as:
$\operatorname*{arg\,min}_{g^{R}}F$ $=\|sBg^{R}A-g\|_{F}^{2}$ (29)
For differentiable $F$,
$\frac{\partial F}{\partial g^{R}}$ $=0\implies 2(\tilde{g}-g)\cdot\frac{\partial\tilde{g}}{\partial g^{R}}=0\implies 2(sBg^{R}A-g)\cdot\frac{\partial(sBg^{R}A)}{\partial g^{R}}=0$ (30)
Using the same trick from before and substituting $g^{R}A=X$, we get:
$2sB^{T}(sBg^{R}A-g)A^{T}=0\implies B^{T}(sBg^{R}A-g)A^{T}=0\implies B^{T}sBg^{R}AA^{T}=B^{T}gA^{T}$ (31)
From Lemma 2, we get:
$B^{T}gA^{T}=g^{R}_{\text{LoRA-XS}}/s\implies B^{T}sBg^{R}AA^{T}=g^{R}_{\text{LoRA-XS}}/s\implies B^{T}Bg^{R}AA^{T}=g^{R}_{\text{LoRA-XS}}/s^{2}$ (32)
Now since $B$ and $A$ are full rank, multiplying both sides by $(B^{T}B)^{-1}$ and $(AA^{T})^{-1}$ on the left and right side respectively gives:
$(B^{T}B)^{-1}(B^{T}Bg^{R}AA^{T})(AA^{T})^{-1}$ $=(B^{T}B)^{-1}g^{R}_{\text{LoRA-XS}}(AA^{T})^{-1}/s^{2}$ (33)
$\text{Therefore,}\quad\boxed{g^{R}=\dfrac{1}{s^{2}}(B^{T}B)^{-1}g^{R}_{\text{LoRA-XS}}(AA^{T})^{-1}}$ (34)
A.4 Proof of Theorem 4
Theorem.
Consider the update for matrix $R$ using the closed-form solution derived in Theorem 3: $R\leftarrow R-\eta g^{R}$ where $\eta\geq 0$ is the learning rate. This update guarantees a reduction in the loss function, similar to traditional gradient descent methods. Specifically, the change in loss $\mathrm{d}L$ is given by: $\mathrm{d}L=-\eta{\langle g^{R}_{LoRA-XS},g^{R}\rangle}=-\eta{\langle g^{R}_{LoRA-XS},\frac{1}{s^{2}}(B^{T}B)^{-1}g^{R}_{LoRA-XS}(AA^{T})^{-1}\rangle}\leq 0$
Proof.
We establish that during optimization, the differential change in the loss function, $\mathrm{d}L$, can be expressed as:
$\mathrm{d}L=-\eta\langle g^{R}_{\text{LoRA-XS}},\frac{1}{s^{2}}(B^{T}B)^{-1}g^{R}_{\text{LoRA-XS}}(AA^{T})^{-1}\rangle_{F}$ (35)
Let us derive this expression step by step. First, the differential change in loss can be written as:
$\mathrm{d}L=\langle\frac{\partial L}{\partial R},\mathrm{d}R\rangle_{F}$ (36)
The update equation for $R$ follows:
$\mathrm{d}R=-\eta g^{R}$ (37)
Given that $\frac{\partial L}{\partial R}=g^{R}_{\text{LoRA-XS}}$, we obtain:
$\mathrm{d}L$ $=-\eta\langle g^{R}_{\text{LoRA-XS}},g^{R}\rangle_{F}$ (38)
$\mathrm{d}L$ $=-\eta\langle g^{R}_{\text{LoRA-XS}},\frac{1}{s^{2}}(B^{T}B)^{-1}g^{R}_{\text{LoRA-XS}}(AA^{T})^{-1}\rangle_{F}$
To prove that $\mathrm{d}L\leq 0$, we demonstrate that:
$\langle g^{R}_{\text{LoRA-XS}},\frac{1}{s^{2}}(B^{T}B)^{-1}g^{R}_{\text{LoRA-XS}}(AA^{T})^{-1}\rangle_{F}\geq 0$ (39)
We first establish that $(B^{T}B)^{-1}$ is positive definite. For any non-zero vector $x$, since $B$ is full-rank:
$\langle x,B^{T}Bx\rangle=\langle Bx,Bx\rangle=\|Bx\|^{2}>0$ (40)
This proves $B^{T}B$, and consequently, $(B^{T}B)^{-1}$ are positive definite. Through Cholesky decomposition, we can express $(B^{T}B)^{-1}=VV^{T}$. Similarly, $(AA^{T})^{-1}$ is positive-definite and can be decomposed as $(AA^{T})^{-1}=UU^{T}$.
Therefore:
$\langle g^{R}_{\text{LoRA-XS}},\frac{1}{s^{2}}(B^{T}B)^{-1}g^{R}_{\text{LoRA-XS}}(AA^{T})^{-1}\rangle_{F}$ $=\frac{1}{s^{2}}\langle g^{R}_{\text{LoRA-XS}},VV^{T}g^{R}_{\text{LoRA-XS}}UU^{T}\rangle_{F}$ (41)
$=\frac{1}{s^{2}}\langle V^{T}g^{R}_{\text{LoRA-XS}}U,V^{T}g^{R}_{\text{LoRA-XS}}U\rangle_{F}$
$=\frac{1}{s^{2}}\|V^{T}g^{R}_{\text{LoRA-XS}}U\|_{F}^{2}\geq 0$
Thus, we have proven:
$\mathrm{d}L=-\eta\langle g^{R}_{\text{LoRA-XS}},\frac{1}{s^{2}}(B^{T}B)^{-1}g^{R}_{\text{LoRA-XS}}(AA^{T})^{-1}\rangle_{F}\leq 0$ (42)
For our specific initialization where $(B^{T}B)=I$, $(AA^{T})=I$, and $s=1$, the result simplifies to:
$\mathrm{d}L$ $=-\eta\langle g^{R}_{\text{LoRA-XS}},g^{R}_{\text{LoRA-XS}}\rangle_{F}\leq 0$ (43)
A.5 Proof of Theorem 5
Theorem.
The equivalent gradient $\tilde{g}$ is hyperparameter $s$ independent when $\tilde{g}=sBg^{R}A\dots\text{ but not when }\tilde{g}=sBg^{R}_{LoRA-XS}A$
Proof.
Let $g$ be the full fine-tuning gradient. We want to prove that $\tilde{g}$ does not depend on $s$, so we try to express it in terms of $g$ which does not depend on the LoRA-XS training process or reparameterization
1) For $\tilde{g}=sBg^{R}A$:
$g^{R}=\frac{1}{s^{2}}(B^{T}B)^{-1}g^{R}_{\text{LoRA-XS}}(AA^{T})^{-1}\implies\tilde{g}=\frac{s}{s^{2}}B(B^{T}B^{-1})g^{R}_{\text{LoRA-XS}}(AA^{T})^{-1}A$ (44)
Now since $g^{R}_{\text{LoRA-XS}}=sB^{T}gA^{T}$:
$\tilde{g}=\frac{1}{s}B(B^{T}B^{-1})sB^{T}gA^{T}(AA^{T})^{-1}A=B(B^{T}B^{-1})B^{T}gA^{T}(AA^{T})^{-1}A$ (45)
which is $s$ independent!
2) For $\tilde{g}=sBg^{R}_{\text{LoRA-XS}}A$
$g^{R}_{\text{LoRA-XS}}=sB^{T}gA^{T}\implies\tilde{g}=sB(sB^{T}gA^{T})A\implies\tilde{g}=s^{2}BB^{T}gA^{T}A$ (46)
which is not $s$ independent! ∎
A.6 Proof of Theorem 6
Theorem.
If $A_{init}$ and $B_{init}$ are initialized using LoRA-SB for the first SGD optimizer, then $\Delta(B_{init}R_{init}A_{init})\approx\Delta W$
Proof.
Consider a gradient descent step with learning rate $\eta$ and updates for $R$:
$\Delta R=-\eta\nabla_{R}\mathcal{L}(R)\implies B\Delta RA=-\eta B\nabla_{R}\mathcal{L}(R)A$ (47)
To measure its approximation quality of update of the weights in full finetuning:
$\Delta W$ $=-\eta\nabla_{W}\mathcal{L}(W_{0})$ (48)
We use Frobenius norm of the difference between these two updates as a criterion:
$\|B\Delta RA-\eta\mathcal{L}_{W}(W_{0})\|_{F}$ $=\eta\|B\nabla_{R}\mathcal{L}(R)A-\mathcal{L}_{W}(W_{0})\|_{F}$ (49)
We have shown before that:
$\nabla_{R}\mathcal{L}$ $=B^{T}\nabla_{W}\mathcal{L}A^{T}$ (50)
The problem becomes:
$\min_{A_{\text{init}},B_{\text{init}}}\|B^{T}(B^{T}\nabla_{W}\mathcal{L}A^{T})A-\nabla_{W}\mathcal{L}\|_{F}\dots\text{ where }\nabla_{W}\mathcal{L}=USV^{T}$ (51)
Using our initialization, we get:
$\|BB^{T}\nabla_{W}\mathcal{L}A^{T}A-\nabla_{W}\mathcal{L}\|_{F}$ $=\|U_{IR}U_{IR}^{T}UsV^{T}V_{IR}V_{IR}^{T}-UsV^{T}\|_{F}$ (52)
Now:
$U_{IR}U_{IR}^{T}UsV^{T}V_{IR}V_{IR}^{T}$ $=\sum_{i=1}^{r}\sigma_{i}u_{i}v_{i}^{T}$ (53)
Now rank of $W^{\prime}$ where:
$W^{\prime}$ $=U_{IR}U_{IR}^{T}UsV^{T}V_{IR}V_{IR}^{T}$ (54)
is $\leq r$ since rank of $B_{\text{init}}$ and $A_{\text{init}}$ is $r$
By Eckart-Young Theorem, the optimal low-rank optimization w.r.t Frobenius norm is:
$W^{\prime*}$ $=\underset{\text{rank}(W^{\prime})=r}{\text{arg min}}\|W^{\prime}-\nabla_{W}\mathcal{L}\|_{F}=\sum_{i=1}^{r}\sigma_{i}u_{i}v_{i}^{T}$ (55)
Since we get an identical expression, our solution is optimal. ∎
Appendix B Inference Costs
As discussed in Section 5, LoRA-SB introduces minimal inference cost overhead due to the insertion of the $r\times r$ matrix $R$ between $B$ and $A$, as well as the requirement for higher ranks to match LoRA’s performance. This comparison is presented in Table 6, showing that the additional overhead of LoRA-SB is negligible.
Table 6: Inference cost comparison between LoRA-SB and LoRA across various models for a sequence length of 256. The minimum rank at which LoRA-SB matches or exceeds LoRA’s performance is highlighted in bold.
Model Method Rank MACs FLOPs
RoBERTa-large LoRA 88 $77.86$ G $155.79$ G
LoRA-SB $16$ $78.42$ G $156.91$ G
LoRA-SB $24$ $78.97$ G $158.01$ G
LlaMA-3.2 3B LoRA $32$ $0.84$ T $1.67$ T
LoRA-SB $64$ $0.85$ T $1.70$ T
LoRA-SB $96$ $0.86$ T $1.72$ T
Mistral 7B LoRA $32$ $1.84$ T $3.69$ T
LoRA-SB $64$ $1.86$ T $3.73$ T
LoRA-SB $92$ $1.88$ T $3.77$ T
Gemma-2 9B LoRA $32$ $3.89$ T $7.77$ T
LoRA-SB $64$ $3.93$ T $7.86$ T
LoRA-SB $96$ $3.97$ T $7.94$ T
Appendix C Experiment Details
We ran our experiments on a single NVIDIA A6000 GPU, averaging results over three independent trials. We trained all models using the AdamW optimizer (Loshchilov & Hutter, 2019).
For arithmetic and commonsense reasoning tasks, we set up Mistral-7B, Gemma-2 9B, and Llama-3.2 3B with hyperparameters and configurations listed in Table 7. We adopted most settings from previous studies (Hu et al., 2023) but conducted our own learning rate sweep. Following LoRA-XS guidelines, we set $\alpha=r$ for their baseline configuration.
For the GLUE benchmark using RoBERTa-large, you can find the hyperparameter details in Table 8. We mostly adhered to the original configurations from the LoRA paper (Hu et al., 2021) but adjusted the learning rate through a sweep. In line with LoRA-XS settings, we fixed $\alpha$ at $16$ for their baseline.
Mistral-7B / Gemma-2 9B Llama-3.2 3B
Optimizer AdamW AdamW
Batch size 11 66
Max. Seq. Len $512$ $256$
Grad Acc. Steps $32$ $24$
Epochs 11 22
Dropout 0 $0.05$
Learning Rate $1\times 10^{-4}$ $2\times 10^{-3}$
LR Scheduler Cosine Linear
Warmup Ratio $0.02$ $0.02$
Table 7: Hyperparameter settings for training Mistral-7B and Gemma-2 9B on MetaMathQA, and Llama-3.2 3B on Commonsense170K.
CoLA RTE MRPC SST-2 QNLI STS-B
Optimizer AdamW
Batch size 128
Max Seq. Len. $256$
Epochs $30$ $30$ $30$ $15$ $15$ $30$
Dropout 0
Learning Rate $1\times 10^{-3}$
LR Scheduler Linear
Warmup Ratio $0.06$
Table 8: hyperparameter settings for RoBERTa-large on GLUE.
Appendix D Dataset Details
The MetaMathQA dataset (Yu et al., 2024) creates mathematical questions by rephrasing existing ones from different viewpoints, without adding new information. We assess this dataset using two benchmarks: GSM8K (Cobbe et al., 2021), which consists of grade-school math problems requiring multi-step reasoning, and MATH (Hendrycks et al., 2021), which presents difficult, competition-level math problems.
CommonSense170K is a comprehensive dataset that consolidates eight commonsense reasoning datasets (Hu et al., 2023), as described below:
  1. 1.
    HellaSwag (Zellers et al., 2019) challenges models to select the most plausible continuation of a given scenario from multiple possible endings.
  2. 2.
    ARC Easy (or ARC-e) (Clark et al., 2018) includes basic science questions at a grade-school level, offering simpler tasks to assess fundamental reasoning abilities.
  3. 3.
    PIQA (Bisk et al., 2020) evaluates physical commonsense reasoning, where models must choose the best action to take in a hypothetical scenario.
  4. 4.
    SIQA (Sap et al., 2019) tests social commonsense reasoning by asking models to predict the social consequences of human actions.
  5. 5.
    WinoGrande (Sakaguchi et al., 2021) presents sentence completion tasks requiring commonsense reasoning to select the correct binary option.
  6. 6.
    ARC Challenge (or ARC-c) (Clark et al., 2018) consists of more complex science questions designed to challenge models with sophisticated reasoning, beyond simple co-occurrence patterns.
  7. 7.
    OBQA (Mihaylov et al., 2018) features open-book, knowledge-intensive QA tasks that require multi-hop reasoning across multiple information sources.
  8. 8.
    BoolQ (Clark et al., 2019) involves answering yes/no questions based on real-world, naturally occurring queries.
The GLUE Benchmark is a comprehensive collection of tasks designed to evaluate natural language understanding (NLU) abilities. It included various datasets, including STS-B for measuring semantic textual similarity (Cer et al., 2017), RTE for recognizing textual entailment, MRPC for detecting paraphrases (Dolan & Brockett, 2005), CoLA for assessing linguistic acceptability (Warstadt et al., 2019), SST-2 for sentiment analysis (Socher et al., 2013), and QNLI for question-answer inference (Rajpurkar et al., 2018). GLUE’s broad scope makes it a standard benchmark for evaluating models like RoBERTa.

### v3

@@
 Abstract
-Low-rank adapters have become a standard approach for efficiently fine-tuning large language models (LLMs), but they often fall short of achieving the performance of full fine-tuning.
+Low-rank adapters have become standard for efficiently fine-tuning large language models (LLMs), but they often fall short of achieving the performance of full fine-tuning.
 We propose a method, LoRA Silver Bullet or LoRA-SB, that approximates full fine-tuning within low-rank subspaces using a carefully designed initialization strategy.
-We theoretically demonstrate that the architecture of LoRA-XS—which inserts a trainable $r\times r$ matrix between $B$ and $A$ while keeping other matrices fixed—provides the precise conditions needed for this approximation.
+We theoretically demonstrate that the architecture of LoRA-XS—which inserts a learnable $r\times r$ matrix between $B$ and $A$ while keeping other matrices fixed—provides the precise conditions needed for this approximation.
 We leverage its constrained update space to achieve optimal scaling for high-rank gradient updates while removing the need for hyperparameter tuning.
 We prove that our initialization offers an optimal low-rank approximation of the initial gradient and preserves update directions throughout training.
-Extensive experiments across mathematical reasoning, commonsense reasoning, and language understanding tasks demonstrate that our approach exceeds the performance of standard LoRA while using 27-90x fewer parameters, and comprehensively outperforms LoRA-XS.
+Extensive experiments across mathematical reasoning, commonsense reasoning, and language understanding tasks demonstrate that our approach exceeds the performance of standard LoRA while using 27-90 times fewer learnable parameters, and comprehensively outperforms LoRA-XS.
 Our findings establish that it is possible to simulate full fine-tuning in low-rank subspaces, and achieve significant efficiency gains without sacrificing performance.
-We have released our code publicly at https://github.com/RaghavSinghal10/lora-sb.
+Our code is publicly available at https://github.com/RaghavSinghal10/lora-sb.
 Machine Learning, ICML
@@
 Pre-trained language models have become central to natural language processing, achieving state-of-the-art performance across diverse tasks (Radford et al., 2021; Kirillov et al., 2023; Achiam et al., 2023).
-While these models excel at general-purpose capabilities (Bubeck et al., 2023; Hao et al., 2022), adapting them to specific downstream tasks often requires fine-tuning.
-Although in-context learning (Brown et al., 2020; Radford et al., 2019) has gained popularity for its simplicity, it falls short in both performance and efficiency compared to fine-tuning (Liu et al., 2022).
-At the same time, full fine-tuning, while highly effective, is computationally expensive and impractical at scale, highlighting the need for more efficient adaptation techniques.
-Parameter-efficient fine-tuning (PEFT) has emerged as a crucial approach for adapting large language models (LLMs) while addressing computational constraints.
-While full fine-tuning (FFT) typically achieves optimal performance, it requires updating billions of parameters, making it computationally intensive and often impractical for most applications.
-Low-rank decomposition methods, beginning with LoRA (Hu et al., 2021), have shown particular promise by significantly reducing trainable parameters through learning low-rank updates.
-This has sparked numerous advances in low-rank methods that either enhance performance through better optimization techniques and initialization strategies, or improve parameter efficiency through structured matrices and adaptive rank selection (Zhang et al., 2023; Wang et al., 2024b, a).
-However, these methods still face fundamental trade-offs, they must either maintain a relatively large number of parameters to match FFT performance, or accept a performance degradation when pursuing extreme parameter efficiency (Hu et al., 2021; Ding et al., 2023; Wang et al., 2024b).
-This raises an important question: can we design low-rank methods that maintain FFT-competitive performance while drastically reducing the parameter count beyond current approaches?
-Figure 1: LoRA-SB Illustration.
-LoRA-XS (Bałazy et al., 2024) reduces parameter count compared to LoRA (Hu et al., 2021) by inserting a trainable $r\times r$ matrix $R$ between $B$ and $A$, while keeping other matrices fixed, leading to $W=W_{0}+sBRA$.
-Our method, LoRA-SB, leverages the same architecture.
-We find that updating $R$ using its gradients $g^{R}$ is equivalent to updating the full-finetuning matrix $W$ with an equivalent gradient $\tilde{g}_{SB}=sBg^{R}A$.
-We initialize $B$, $R$, and AA such that the equivalent gradient $\tilde{g}_{SB}$ optimally approximates the full fine-tuning gradient $g$ in low rank subspaces at each training step.
-In essence, we simulate the entire full fine-tuning process optimally within low-rank subspaces by utilizing only the first gradient $g_{1}$ (shown in green) from full fine-tuning.
-Low-rank decomposition methods operate on a fundamental premise: fine-tuning requires learning only a low-rank update to the pre-trained weights.
-Some theoretical work extends this hypothesis, suggesting that methods like LoRA can learn any low-rank approximation of the full fine-tuning gradient.
-However, the gradients computed by these methods to update their trainable adapters do not inherently possess this property.
-For instance, LoRA’s gradients need explicit optimization at each step to better approximate the full fine-tuning gradient (Wang et al., 2024b).
-Additionally, initialization has emerged as a critical factor in low-rank adaptation, as highlighted by recent works like Pissa-LoRA (Meng et al., 2024) and LoRA-GA (Wang et al., 2024a), which address this challenge through various low-rank approximation techniques.
-We formally analyze these limitations in the context of the architecture used in LoRA-XS (Bałazy et al., 2024)—which inserts a trainable $r\times r$ matrix between $B$ and $A$ while keeping other matrices fixed—and demonstrate that these challenges are even more pronounced in its framework.
-While exploring solutions inspired by existing LoRA-based methods, we discover a remarkable property unique to LoRA-XS: through careful initialization of matrices $A$ and $B$, we can approximately simulate the full fine-tuning optimization in low rank subspaces throughout the entire training, as shown in Figure 1.
-Our initialization strategy provides optimal scaling for approximating high-rank full fine-tuning gradients and eliminates the need for the scaling hyperparameter $\alpha$, results which we prove theoretically.
-Our key contributions include:
+While these models excel at general-purpose capabilities (Bubeck et al., 2023; Hao et al., 2022), adapting them to specific downstream tasks often requires fine-tuning (FT).
+Although in-context learning (Brown et al., 2020; Radford et al., 2019) has gained popularity for its simplicity, it falls short in both performance and efficiency compared to FT (Liu et al., 2022).
+At the same time, full FT, while highly effective, is computationally expensive and impractical at scale.
+Parameter-efficient fine-tuning (PEFT) has become vital for adapting large language models (LLMs) under computational constraints.
+Low-rank methods like LoRA (Hu et al., 2021) address this by reducing learnable parameters via low-rank updates, sparking advancements in optimization, initialization, structured matrices, and adaptive rank selection (Zhang et al., 2023; Wang et al., 2024b, a).
+However, these methods face trade-offs: either retain many parameters to match full FT or sacrifice performance for extreme efficiency (Hu et al., 2021; Ding et al., 2023; Wang et al., 2024b).
+This raises a critical question: Can we design low-rank methods that achieve full FT-level performance while drastically reducing parameter counts?
+Figure 1: LoRA-SB.
+LoRA-XS (Bałazy et al., 2024) reduces parameters compared to LoRA (Hu et al., 2021) by inserting a learnable $r\times r$ matrix $R$ between $B$ and $A$, while keeping other matrices fixed, leading to $W=W_{0}+sBRA$.
+Our method, LoRA-SB, uses the same architecture.
+We find that updating $R$ using its gradients $g^{R}$ is equivalent to updating the full FT matrix $W$ with an equivalent gradient $\tilde{g}_{SB}=sBg^{R}A$.
+We initialize $B$, $R$, and $A$ such that the equivalent gradient $\tilde{g}_{SB}$ provably best approximates the full FT gradient $g$ in low rank subspaces at each step.
+In essence, we simulate the entire full FT process optimally within low-rank subspaces by utilizing only the first gradient $g_{1}$ (shown in green) from full FT.
+Low-rank decomposition methods operate on a fundamental premise: FT requires learning only a low-rank update to the pre-trained weights.
+Some theoretical work extends this hypothesis, suggesting that methods like LoRA can learn any low-rank approximation of the full FT gradient.
+However, the gradients computed by these methods do not inherently possess this property.
+For instance, LoRA’s gradients need explicit optimization at each step to better approximate the full FT gradient (Wang et al., 2024b).
+Additionally, initialization has emerged as a critical factor in low-rank adaptation, as highlighted by recent works like Pissa-LoRA (Meng et al., 2024) and LoRA-GA (Wang et al., 2024a).
+We analyze these limitations in the context of the architecture of LoRA-XS (Bałazy et al., 2024)—which inserts a learnable $r\times r$ matrix between $B$ and $A$ while keeping other matrices fixed—and demonstrate that these challenges are even more pronounced.
+While exploring solutions inspired by LoRA-based methods, we discover a remarkable property unique to LoRA-XS: through careful initialization of $A$ and $B$, we can simulate the full FT optimization in low rank subspaces through entire training, as shown in Figure 1.
+Our initialization provides optimal scaling for approximating high-rank full FT gradients and eliminates need for scaling the hyperparameter $\alpha$.
+Our key contributions are:
   • We theoretically formalize the limitations of LoRA-XS, showing how its constrained update space leads to suboptimal gradient approximation, initialization sensitivity, and hyperparameter dependence.
-  • We propose a principled initialization strategy derived from approximating the first step of full fine-tuning, proving it provides optimal low-rank approximation of the initial gradient and preserves update directions throughout training.
+  • We propose an initialization strategy derived from approximating the first step of full FT, proving it provides optimal low-rank approximation of the initial gradient and preserves update directions throughout training.
   • We prove that our initialization makes gradient optimization hyperparameter-independent and guarantees convergence by maintaining orthonormal bases, eliminating the need for any tuning of the scaling factor.
-  • Through extensive experiments on $4$ models across $16$ datasets covering mathematical reasoning, commonsense reasoning, and language understanding tasks, we demonstrate that our method surpasses the performance of LoRA while using 27-90x less parameters, and comprehensively outperforms LoRA-XS.
+  • Through extensive experiments on $4$ models across $16$ datasets covering mathematical reasoning, commonsense reasoning, and language understanding tasks, we demonstrate that our method surpasses LoRA’s performance while using 27-90x less learnable parameters, and comprehensively outperforms LoRA-XS.
 2 Related Work
-Parameter-Efficient Fine-Tuning.
-Parameter-Efficient Fine-Tuning (PEFT) methods have become essential for adapting large pre-trained models under computational constraints.
+Parameter-Efficient Fine-Tuning (PEFT).
+PEFT methods have become essential for adapting large pre-trained models under computational constraints.
 Early techniques like AdapterFusion (Pfeiffer et al., 2021) and Prefix-Tuning (Li & Liang, 2021) enabled task-specific adaptation with minimal parameter updates.
 Advances like soft prompts (Lester et al., 2021) further reduced trainable parameter counts while maintaining strong performance.
-Recent approaches have explored operating directly on model representations (Wu et al., 2024), offering different trade-offs between efficiency and performance.
+Recent approaches have explored operating directly on model representations (Wu et al., 2024).
 Low-Rank Decomposition Methods.
-LoRA (Hu et al., 2021) demonstrated that weight updates during fine-tuning could be efficiently approximated using low-rank matrices, drastically reducing parameter counts while freezing pre-trained weights.
+LoRA (Hu et al., 2021) demonstrated that weight updates during FT could be efficiently approximated using low-rank matrices, drastically reducing parameter counts.
 Building on this insight, variants such as QLoRA (Dettmers et al., 2023) and AdaLoRA (Zhang et al., 2023) extended the paradigm through quantization and adaptive allocation strategies.
-The applicability of low-rank techniques has also been explored in pretraining with GaLore (Zhao et al., 2024) and ReLoRA (Lialin et al., 2023), highlighting the versatility of low-rank adaptation in both fine-tuning and pre-training contexts.
-LoRA-based methods have also been applied in other domains, such as efficient federated fine-tuning (Sun et al., 2024; Singhal et al., 2024).
+The applicability of low-rank techniques has also been explored in pretraining with GaLore (Zhao et al., 2024) and ReLoRA (Lialin et al., 2023), highlighting the versatility of low-rank adaptation methods.
+LoRA-based methods have also been applied in other domains, such as efficient federated FT (Sun et al., 2024; Singhal et al., 2025).
 Enhancing LoRA Performance.
 Recent efforts have focused on optimizing LoRA’s performance.
-PiSSA-LoRA (Meng et al., 2024) demonstrated faster convergence and improved task performance by initializing matrices with principal components of pre-trained weights.
-LoRA-Pro (Wang et al., 2024b) and LoRA-GA (Wang et al., 2024a) improved gradient approximation, aligning low-rank updates more closely with full fine-tuning.
-Other methods like DoRA (Liu et al., 2024) and rsLoRA (Kalajdzievski, 2023) introduced decomposition-based and scaling stabilization techniques to enhance learning stability and expand LoRA’s utility.
+PiSSA (Meng et al., 2024) demonstrated improvements by initializing matrices with principal components of pre-trained weights.
+LoRA-Pro (Wang et al., 2024b) and LoRA-GA (Wang et al., 2024a) improved gradient approximation, aligning low-rank updates more closely with full FT.
+Methods like DoRA (Liu et al., 2024) and rsLoRA (Kalajdzievski, 2023) introduced decomposition-based and scaling stabilization techniques to enhance learning stability and expand LoRA’s utility.
 Improving Efficiency in LoRA Variants.
-Efficiency-focused innovations have pushed LoRA toward even greater parameter savings.
-LoRA-XS (Bałazy et al., 2024) achieves parameter reduction by inserting a small trainable weight matrix into frozen low-rank matrices.
+Efficiency-focused innovations have pushed LoRA toward more parameter savings.
+LoRA-XS (Bałazy et al., 2024) achieves this by inserting a small trainable weight matrix into frozen low-rank matrices.
 VeRA (Kopiczko et al., 2024) shares low-rank matrices across layers, relying on scaling vectors for task-specific adaptation.
-Tied-LoRA (Renduchintala et al., 2024) leverages weight tying to reduce parameter usage at higher ranks, while HydraLoRA (Tian et al., 2024) introduces an asymmetric architecture to improve adaptation in complex domains without domain-specific expertise.
+Tied-LoRA (Renduchintala et al., 2024) leverages weight tying to reduce parameter usage at higher ranks, while HydraLoRA (Tian et al., 2024) introduces an asymmetric architecture for improvement.
 3 Methodology
 3.1 Preliminaries
-In standard fine-tuning, a pre-trained weight matrix $W\in\mathbb{R}^{m\times n}$ is updated as:
+In standard FT, a pre-trained weight matrix $W\in\mathbb{R}^{m\times n}$ is updated using the update matrix $\Delta W$ as:
 $W=W_{0}+\Delta W$ (1)
-where $W_{0}$ is the pre-trained weight. This requires updating $mn$ parameters per layer.
-LoRA hypothesizes that these updates lie in a low-dimensional subspace and parameterizes the update matrix $\Delta W$ as:
+where $W_{0}$ is the pre-trained weight. This requires updating $mn$ parameters per layer. LoRA posits that updates lie in a low-dimensional subspace, parameterizing $\Delta W$ as:
 $W=W_{0}+sBA$ (2)
-where $B\in\mathbb{R}^{m\times r}$ and $A\in\mathbb{R}^{r\times n}$ are trainable low-rank matrices, with rank $r\ll\min(m,n)$, and $s$ is a scaling factor to stabilize training.
-This reduces parameters from $mn$ to $r(m+n)$.
-LoRA-XS introduces a more efficient parameterization:
+where $B\in\mathbb{R}^{m\times r}$ and $A\in\mathbb{R}^{r\times n}$ are trainable low-rank matrices with rank $r\ll\min(m,n)$, and $s$ is a scaling factor ($\alpha/r$) to stabilize training.
+This reduces the number of parameters from $mn$ to $r(m+n)$.
+LoRA-XS efficiently parameterizes as:
 $W=W_{0}+sBRA$ (3)
-where $B\in\mathbb{R}^{m\times r}$ and $A\in\mathbb{R}^{r\times n}$ are fixed matrices, and only $R\in\mathbb{R}^{r\times r}$ is trainable, reducing parameters to $r^{2}$.
-We denote full fine-tuning gradient as $g=\frac{\partial L}{\partial W}$, and LoRA-XS gradient as $g^{R}_{\text{LoRA-XS}}=\frac{\partial L}{\partial R}$.
+where $B$ and $A$ are fixed, and only $R\in\mathbb{R}^{r\times r}$ is trainable, reducing the number of parameters to $r^{2}$.
+We denote the full FT gradient as $g=\frac{\partial L}{\partial W}$, and the LoRA-XS gradient as $g^{R}_{\text{LoRA-XS}}=\frac{\partial L}{\partial R}$, where $L$ is the loss function.
 3.2 Motivation
-LoRA-XS (Bałazy et al., 2024), although having significantly fewer learnable parameters compared to LoRA, exhibits suboptimal performance.
-The difference in architecture causes constraints on the type of updates LoRA-XS can learn.
+LoRA-XS (Bałazy et al., 2024) has significantly fewer learnable parameters than LoRA but performs suboptimally.
+LoRA-XS’s architecture causes constraints on the type of updates it can learn.
 The subspace of learned updates is characterized in Lemma 1.
+This implies that while $\Delta W$ is constrained to be rank $\leq r$, it also needs to have column and row spaces defined by those of $B$ and $A$, respectively.
+In contrast, LoRA can learn any update $\Delta W$ as long as $\text{rank}(\Delta W)\leq r$.
+Thus, the low expressivity of LoRA-XS as compared to LoRA can account for the performance drop.
 Lemma 1.
-Let $\Delta W$ be an update learned by LoRA-XS .
-Then, the set of all $\Delta W$, say $\mathcal{W}_{LoRA-XS}$, can be given by $\{M\in\mathbb{R}^{m\times n}|\text{Col}(W)\subseteq\text{Col}(M)$  $\wedge\text{ Row}(M)\subseteq\text{Row}(A)\}$
+Let $\Delta W$ be an update learned with LoRA-XS.
+Then, the set of all possible $\Delta W$, say $\mathcal{W}_{LoRA-XS}$, is given as: $\{M\in\mathbb{R}^{m\times n}|\text{Col}(M)\subseteq\text{Col}(B)$ $\wedge\text{Row}(M)\subseteq\text{Row}(A)\},$ where $\text{Col}(M)$ and $\text{Row}(M)$ are column and row spaces of matrix $M$ respectively.
 Proof.
-Appendix A.2
-The above lemma implies that while $\Delta W$ is constrained to be $r$ ranked or lower, it also needs to have row and column spaces defined by those of $B$ and $A$ respectively.
-As against this, LoRA can learn any update $\Delta W$ as long as $\text{rank}(\Delta W)\leq r$.
-Thus the low expressivity of LoRA-XS as compared to LoRA can account for the performance drop.
-We identify three key limitations which arise due to this and otherwise:
-1) Inadequate Gradient Approximation: LoRA optimization is mathematically equivalent to full fine-tuning using a constrained low-rank gradient.
+See Appendix A.1. ∎
+We identify three key limitations, which arise due to this and otherwise:
+1) Inadequate Gradient Approximation: LoRA optimization is mathematically equivalent to full FT using a constrained low-rank gradient.
 The gradient of LoRA does not optimally approximate the full gradient, and needs to be tuned at each step.
 LoRA-Pro (Wang et al., 2024b) finds that this results in suboptimal performances, and provides a closed form solution to optimize the gradients.
-In LoRA-XS , the gradient updates are restricted to an even more constrained low-rank space since the matrices $A$ and $B$ are fixed.
+In LoRA-XS, the gradient updates are restricted to an even more constrained low-rank space since $A$ and $B$ are fixed.
 We posit that the limitation becomes particularly severe when the ideal updates lie outside the space spanned by fixed $A$ and $B$, and consequently has a larger impact on performance.
 2) Suboptimal Initialization: While initialization impacts all low-rank methods, it becomes critical in LoRA-XS where $A$ and $B$ are frozen.
-Unlike LoRA where poor initialization can be compensated through training, LoRA-XS must rely entirely on its initial subspace defined by $A$ and $B$.
+Unlike LoRA where poor initialization can be compensated through training, LoRA-XS relies entirely on its initial subspace defined by $A$ and $B$.
 Consider the zero initialization of the $B$ matrix, for example.
-While LoRA may experience some performance degradation in this case (Wang et al., 2024a) (Meng et al., 2024), the ideal low-rank update $\Delta W$ can still be reached through gradient descent.
-In fact, zero initialization for the $B$ matrix is a commonly used choice, including in the original LoRA implementation (Hu et al., 2021).
-However, in the case of LoRA-XS , this would result in no learning, as the product $BRA$ would remain zero.
-The current LoRA-XS method leverages the most significant subspaces spanned by the columns of pre-trained weights $W_{0}$, inspired by (Meng et al., 2024).
-This initialization is inadequate because it fails to capture the specific subspaces relevant to the fine-tuning task.
-This issue, known to affect LoRA ’s performance (Wang et al., 2024a), could have an even greater impact on LoRA-XS for the reasons discussed above.
-3) Hyperparameter Sensitivity: The scaling factor $\alpha$, present in almost every LoRA based fine-tuning method requires careful tuning to maintain stability during training.
-This hyperparameter acts as a bridge between the low-rank and full-rank spaces, compensating for the dimensional mismatch in gradients.
-Poor tuning of $\alpha$ can lead to unstable training or slow convergence (Kalajdzievski, 2023), adding complexity to the adaptation process and potentially limiting practical deployment.
-3.3 Approximation of the full fine-tuning gradient
-As mentioned above, LoRA optimization is mathematically equivalent to full fine-tuning using a constrained low-rank gradient.
-However the update generated using the gradients of LoRA does not result in the same update which the low-rank gradient would have generated.
+While LoRA may experience some performance degradation in this case (Wang et al., 2024a; Meng et al., 2024), the ideal low-rank update $\Delta W$ can still be reached through gradient descent.
+In fact, zero initialization for the $B$ matrix is commonly used, including in the original LoRA paper (Hu et al., 2021).
+However, in LoRA-XS, this results in no learning, as the product $BRA$ remains zero.
+LoRA-XS uses the most significant subspaces spanned by the columns of pre-trained weights for initialization, inspired by Meng et al.
+(2024).
+This initialization is not aligned well with FT because it fails to capture the specific subspaces relevant to the FT task.
+3) Hyperparameter Sensitivity: The scaling factor $s$, present in almost every LoRA based FT method requires tuning to maintain stability during training.
+This factor acts as a bridge between the low-rank and full-rank spaces, compensating for the dimensional mismatch in gradients.
+Poor tuning of $s$ can lead to unstable training or slow convergence, e.g., see rsLoRA (Kalajdzievski, 2023), adding complexity and potentially limiting practical deployment.
+3.3 Approximation of the full FT gradient
+As mentioned, LoRA optimization is mathematically equivalent to full FT using a constrained low-rank gradient.
+However, the update generated using the gradients of LoRA does not result in the same update which the low-rank gradient would have generated.
 The following holds true for LoRA-XS as well.
-To understand this, let us look at the change in weight $W$ and its relationship with changing of low-rank matrix $R$.
-The relationship can be simply given by $dW=\frac{\partial W}{\partial R}g^{R}$.
-This relationship implies that updating matrix $R$ with gradient $g^{R}$ is equivalent to updating $W$ with low rank equivalent gradient $\tilde{g}$ in full fine-tuning as described in Definition 1.
-Definition 1 (Equivalent Gradient).
-In the context of LoRA-XS optimization, we define the equivalent gradient as: $\tilde{g}=sBg^{R}A$ where $s$ is the scaling factor, and $g^{R}$ is the gradient with respect to matrix $R$.
+To understand this, let us look at the change in weight $W$ and its relationship with changing of low-rank matrix $R$, which can be simply given by $\mathrm{d}W=-sB(\mathrm{d}R)A.$ This implies that updating $R$ with gradient $g^{R}$ is equivalent to updating $W$ with low rank equivalent gradient $\tilde{g}$ in full FT as described in Definition 1.
+Definition 1.
+We define the equivalent gradient as: $\tilde{g}=sBg^{R}A$ where $g^{R}$ is the gradient of $L$ with respect to $R$.
 The equivalent gradient describes the virtual low-rank gradient of matrix $W$ in LoRA-XS optimization process, despite $W$ not being directly trainable.
-This gradient determines how updates to $R$ affect the overall weight matrix.
-To bridge the performance gap between LoRA-XS and full fine-tuning, our goal is to minimize the discrepancy between the equivalent gradient $\tilde{g}$ and the full gradient $g$.
+This gradient determines how updates to $R$ affect $W$.
+To bridge the performance gap between LoRA-XS and full FT, we aim to minimize the discrepancy between the equivalent gradient $\tilde{g}$ and the full gradient $g$.
 First, we establish the relationship between gradients in LoRA-XS optimization in Lemma 2.
 Lemma 2.
-During LoRA-XS optimization, the gradient of the loss with respect to matrix $R$ can be expressed in terms of the gradient with respect to the weight matrix $W$ as: $g^{R}_{LoRA-XS}=sB^{T}gA^{T}$
+The gradient of the loss with respect to matrix $R$ can be expressed in terms of the gradient with respect to the weight matrix $W$ as: $g^{R}_{LoRA-XS}=sB^{\top}gA^{\top}$
 Proof.
-Appendix A.2
-Leveraging this relationship, we can formulate our objective to minimize the distance between the equivalent gradient and the full gradient.
-We do not have access to the full fine-tuning gradient $g$ during LoRA-XS based fine-tuning.
+See Appendix A.2. ∎
+We can now formulate our objective to minimize the distance between the equivalent gradient and the full gradient.
+We do not have access to the full FT gradient $g$ during LoRA-XS based FT.
 Thus we need to find the ideal gradient with respect to $R$, given by $g^{R}$, and subsequently the optimal approximation $\tilde{g}$, in terms of the gradient which is available to us during training: $g^{R}_{LoRA-XS}$.
@@
 Theorem 3.
-[…]
+The optimal solution for the objective $\text{min}_{g^{R}}||\tilde{g}-g||^{2}_{F}$, such that $\tilde{g}=sBg^{R}A$, is: $g^{R}=\dfrac{1}{s^{2}}(B^{\top}B)^{-1}g^{R}_{LoRA-XS}(AA^{\top})^{-1}$ (4)
 Proof.
-Appendix A.3. ∎
-[…]
+See Appendix A.3. ∎
+The closed-form solution in Theorem 3 solves the optimization problem $\text{min}_{g^{R}}||\tilde{g}-g||^{2}_{F}$, but by itself doesn’t ensure the loss will decrease when updating $R$.
+Through Theorem 4, we prove that the change in loss is non-positive ($\Delta L\leq 0$).
+This property is fundamental to optimization as it guarantees consistent loss minimization throughout training.
 Theorem 4.
-Consider the update for matrix $R$ using the closed-form solution derived in Theorem 3: $R\leftarrow R-\eta g^{R}$ where $\eta\geq 0$ is the learning rate.
-This update guarantees a reduction in the loss function, similar to traditional gradient descent methods.
-Specifically, the change in loss $\mathrm{d}L$ is given by: $\mathrm{d}L=-\eta\langle g^{R}_{LoRA-XS},g^{R}\rangle\leq 0$
+Consider the update for matrix $R$ using the solution derived in Theorem 3: $R\leftarrow R-\eta g^{R}$ where $\eta>0$ is the (sufficiently small) learning rate.
+This update guarantees a reduction in the loss $\Delta L$, given by: $\Delta L=-\eta\langle g^{R}_{LoRA-XS},g^{R}\rangle_{F}+o(\eta)\leq 0.$
 Proof.
-Appendix A.4. ∎
+See Appendix A.4. ∎
 3.4 Initialization using update approximation
-Let us first consider the objective in fine-tuning, the primary goal is to update weights to better suit the target task and domain.
+In FT, the primary goal is to update weights to better suit the target task.
 The initial gradient steps are particularly informative, as they indicate the direction of desired adaptation.
-We propose leveraging this insight by using the first update step from full fine-tuning to initialize our low-rank matrices.
-This approach offers two key advantages: First, it ensures the low-rank space captures the most relevant subspace for the target task rather than relying on pre-trained weight properties.
-Second, since $A$ and $B$ matrices will remain fixed in LoRA-XS, initializing them to span the subspace of early adaptation increases the likelihood of capturing useful updates throughout training.
-This would also ensure that the final update is learnt in the correct subspace, of which we have no apriori information besides the first step of full fine-tuning.
-In essence, our method can be summarized as follows: Set an initialization which best approximates the first step of full fine-tuning.
-Formally, given a full fine-tuning update $\Delta W$, our initialization should satisfy:
+We leverage this insight by using the first update step from full FT for initialization.
+This approach offers two key advantages.
+First, it ensures the low-rank space captures the most relevant subspace for the target task rather than relying on pre-trained properties.
+Second, since $A$ and $B$ are fixed, initializing them to span the subspace of early adaptation increases the likelihood of capturing useful updates throughout training.
+This also ensures that the final update is learnt in the correct subspace, of which we have no apriori information besides the first full FT step.
+Our method is summarized as: set such initialization that best approximates the first step of full FT.
+Given a full FT update $\Delta W_{first-step}$, our initialization satisfies:
 $sB_{init}R_{init}A_{init}\approx\Delta W_{first-step}$ (5)
-The first step of full fine-tuning, for Adam-based optimizers such as AdamW, for a sample $x_{i}$ is:
+The first step of full FT, for Adam-based optimizers such as AdamW, for sample $x_{i}$ is:
 $\Delta W_{first-step}=-\eta\times\textbf{sign}(\nabla_{W}\mathcal{L}(W_{0},x_{i}))$ (6)
-Using a single sample may lead to noisy estimates. Instead, we compute a more stable initialization by averaging gradients over a subset of the training data:
-$\Delta W_{avg}=-\eta\textbf{sign}(\sum_{i=0}^{n\leq|\mathbb{X}|}\nabla_{W}\mathcal{L}(W_{0},x_{i}))\dots x_{i}\in\mathbb{X}$ (7)
-This averaged gradient better captures the general direction of adaptation required for the target task while being less sensitive to individual sample variations.
+However, the usage of a single sample may lead to noisy estimates. Instead, we compute a more stable initialization by averaging gradients over a subset of the training data:
+$\Delta W_{avg}=-\eta\textbf{sign}(\sum_{i=0}^{n\leq|\mathbb{X}|}\nabla_{W}\mathcal{L}(W_{0},x_{i})),\quad x_{i}\in\mathbb{X}$ (7)
+This better captures the general direction of adaptation required for the target task while being less sensitive to individual sample variations.
 We can then use truncated SVD to obtain a low-rank approximation of $\Delta W_{\text{avg}}$, and express it as $sBRA$.
-There exist infinite combinations of $B$ and $A$ which can obey the above relationship.
-For instance, we can initialize $B$ and $A$ as $US$ and $V^{T}$ and keep $R$ as $I/s$.
-This initialization is equivalent to the $B$ and $A$ initialization in LoRA-XS but by approximating the update rather than the pre-trained matrix.
+There exist infinite combinations of $B$ and $A$ which can obey this relationship.
+For instance, we can initialize $B$ and $A$ as $US$ and $V^{\top}$ and keep $R$ as $I/s$.
+This is equivalent to the $B$ and $A$ initialization in LoRA-XS but by approximating the update rather than the pre-trained matrix.
 We note that the above process can be computed for any optimizer, by approximating the corresponding first step.
@@
 3.5 Hyperparameter independence
-The hyperparameter $\alpha$ is used in LoRA and subsequent decomposition based models to tackle the issue of instability caused to improper scaling of the updates.
+The hyperparameter $\alpha$ is used in LoRA and other decomposition-based method to tackle the issue of instability caused to improper scaling of the updates.
 The gradient scaling is accounted for, by adding a hyperparameter to normalize the updates.
-The importance of scaling is well documented in methods like rank stabilization (Kalajdzievski, 2023).
-However, the full fine-tuning gradient $g$ needs no such tuning.
-We claim that approximating the full fine-tuning gradient removes the need for introducing a scaling factor.
+The importance of scaling is shown in methods like rank stabilization (Kalajdzievski, 2023).
+However, the full FT gradient $g$ needs no such tuning.
+We claim that approximating the full FT gradient removes the need for introducing a scaling factor, as shown in Theorem 5.
 Theorem 5.
@@
 Proof.
-Appendix A.5
+See Appendix A.5. ∎
 The hyperparameter independence of the equivalent gradient eliminates the need for manual gradient scaling.
-Updates to weight matrix $W$ depend solely on this gradient (modulo learning rate), making any additional scaling redundant.
-This can be understood by examining the relationship with full fine-tuning gradient $g$.
+Updates to $W$ depend solely on this gradient (modulo learning rate), making any additional scaling redundant.
+This can be understood by examining the relationship with the full FT gradient $g$.
 Since $g$ is naturally scaled for optimal weight updates, and our method approximates $g$ in a constrained subspace, the equivalent gradient inherits appropriate scaling automatically.
-Notably, this property is unique to our gradient approximation approach and does not hold for standard LoRA-XS training.
+This property is unique to our gradient approximation approach and does not hold for standard LoRA-XS.
 3.6 LoRA-SB: Update approximation initialization is a silver bullet
 The solutions discussed above independently address the gradient approximation and initialization problems, while also providing hyperparameter independence.
-Our proposed method, LoRA-SB, elegantly combines these solutions through a simple initialization strategy.
-Our initialization strategy is derived from approximating the first step of full fine-tuning:
-$U,S,V^{T}\leftarrow\textbf{SVD}(\Delta W_{avg})$ (8)
+Our proposed method, LoRA-SB, elegantly combines these solutions through a simple initialization strategy, derived from approximating the first step of full FT:
+$U,S,V^{\top}\leftarrow\textbf{SVD}(\Delta W_{avg})$ (8)
 $B_{init}\leftarrow U[1:r]$ (9)
@@
 $R_{init}\leftarrow\dfrac{1}{s}S[1:r,1:r]$ (11)
-By the Eckart-Young theorem (Eckart & Young, 1936; Mirsky, 1960), this gives the optimal rank-$r$ approximation of the initial full fine-tuning update.
+By the Eckart-Young theorem (Eckart & Young, 1936; Mirsky, 1960), this gives the optimal rank-$r$ approximation of the full FT update.
 where $U$, $S$, $V$ are obtained from truncated SVD of the averaged first update $\Delta W_{\text{avg}}$.
-This initialization leads to several key advantages that address the problems identified earlier.
+This initialization leads to several key advantages.
 Simplified Gradient Optimization.
-This initialization ensures $B_{\text{init}}$ and $A_{\text{init}}$ form orthonormal bases in $\mathbb{R}^{m}$ and $\mathbb{R}^{n}$ respectively, leading to $B^{T}B=AA^{T}=I$.
-With fixed $B$ and $A$ matrices being orthonormal, the optimal update step derived in Equation 3 simplifies to:
-$g^{R}=\dfrac{1}{s^{2}}(B^{T}B)^{-1}g^{R}_{LoRA-XS}(AA^{T})^{-1}=\frac{1}{s^{2}}g^{R}_{LoRA-XS}$ (12)
-This eliminates the need for complex matrix inversions during training.
-Optimum Update Approximation. Our initialization guarantees that the first update optimally approximates the full fine-tuning weight updates:
-$sB_{\text{init}}R_{\text{init}}A_{\text{init}}\approx\Delta W_{avg}$ (13)
-By the Eckart-Young theorem, this gives the optimal rank-$r$ approximation of the initial full fine-tuning update.
-Hyperparameter Independence. As shown in Theorem 5, when gradient approximation is applied with orthonormal $B$ and $A$, the hyperparameter $s$ can be set to 1, resulting in the following:
+Our initialization ensures $B_{\text{init}}$ and $A_{\text{init}}$ form orthonormal bases in $\mathbb{R}^{m}$ and $\mathbb{R}^{n}$ respectively, leading to $B^{\top}B=AA^{\top}=I$.
+With fixed $B$ and $A$ matrices being orthonormal, the need for complex matrix inversions during training is eliminated, , as the optimal update step, derived in Equation 3, simplifies to:
+$g^{R}=\dfrac{1}{s^{2}}(B^{\top}B)^{-1}g^{R}_{LoRA-XS}(AA^{\top})^{-1}=\frac{1}{s^{2}}g^{R}_{LoRA-XS}.$
+Optimal Update Approximation. Our initialization guarantees that the first update optimally approximates the full FT weight updates:
+$sB_{\text{init}}R_{\text{init}}A_{\text{init}}\approx\Delta W_{avg}.$
+By the Eckart-Young theorem, this gives the optimal rank-$r$ approximation of the initial full FT update.
+Hyperparameter Independence. As shown in Theorem 5, when gradient approximation is applied with orthonormal $B$ and $A$, the hyperparameter $s$ can be set to 1, resulting in:
 $\boxed{g^{R}=g^{R}_{\text{LoRA-XS}}}$ (12)
-This demonstrates that our initialization guarantees optimal gradient approximation at every step, without requiring any scaling factor!
+This demonstrates that our initialization guarantees optimal gradient approximation at every step, without requiring any scaling factor.
 Guaranteed Loss Reduction.
@@
 Thankfully, we don’t have to choose between the two.
-For SGD, we prove that setting $B_{init}$ and $A_{init}$ using Equations 8-11, results in the first update of LoRA-XS to best approximate the direction of the update of full fine-tuning.
+For SGD, we prove that setting $B_{init}$ and $A_{init}$ using Equations 8-11, results in the first update of LoRA-XS to best approximate the direction of the update of full FT (Theorem 6).
 Theorem 6.
-If $A_{init}$ and $B_{init}$ are initialized using LoRA-SB for the first SGD optimizer, then $\Delta(B_{init}R_{init}A_{init})\approx\Delta W$
+If $A_{init}$ and $B_{init}$ are initialized using LoRA-SB for the first step of SGD optimizer, then $\Delta(B_{init}R_{init}A_{init})\approx\Delta W$
 Proof.
-Appendix A.6
+See Appendix A.6. ∎
 LoRA-SB Advantages over LoRA.
-Notably, many of the properties described above are not achievable with standard LoRA methods.
+Many properties described above are not achievable with standard LoRA methods.
 Even if $B$ and $A$ are initialized as orthonormal in LoRA, subsequent updates do not preserve this property because $B$ and $A$ are trainable.
 This results in several challenges:
-  • Potential instability of $(B^{T}B)^{-1}$ and $(AA^{T})^{-1}$, as they are not guaranteed to remain non-singular during training.
+  • Potential instability of $(B^{\top}B)^{-1}$ and $(AA^{\top})^{-1}$, as they are not guaranteed to remain non-singular during training.
   • Inability to ensure consistent loss reduction due to potential rank deficiency—$B$ and $A$ may not remain full-rank throughout training.
   • Necessity to fine-tune the hyperparameter $\alpha$.
-  • Repeated re-computation of $B^{T}B$ and $AA^{T}$ is required at each optimizer step for accurate gradient approximation.
-LoRA-SB’s simple initialization strategy solves multiple problems simultaneously, offering a robust and efficient approach to low-rank adaptation.
+  • Repeated re-computation of $B^{\top}B$ and $AA^{\top}$ is required at each optimizer step for accurate gradient approximation.
 Algorithm.
 We present a PyTorch-like implementation of our method in Algorithm 1.
-To optimize GPU memory usage during initialization, we hook into the backward pass of PyTorch and compute the gradient for one layer at a time, immediately discarding the computed gradients (Lv et al., 2024; Wang et al., 2024a).
-This ensures that memory usage remains at $O(1)$, independent of the number of layers.
-This approach keeps memory consumption under control, ensuring it never exceeds the requirements of subsequent LoRA-SB fine-tuning.
-For very large batch sizes, one can also use gradient accumulation and quantization to further manage memory usage within the specified limits.
+To optimize GPU memory usage during initialization, we hook into the backward pass of PyTorch and compute the gradients layerwise, immediately discarding the computed gradients (Lv et al., 2024; Wang et al., 2024a).
+This ensures $O(1)$ memory usage, independent of the number of layers, keeping memory consumption well within manageable limits and ensuring it does not exceed the memory requirements of subsequent LoRA-SB FT.
+For large batch sizes, memory usage can be further optimized through gradient accumulation and quantization.
 Algorithm 1 LoRA-SB, PyTorch-like
@@
        # Convert to LoRA-SB model
-   $\text{{sb\_model}}\leftarrow$ lora_SB(model, B, R, A)
+[…]
        return sb_model
   # Load pre-trained model
-$\text{{model}}\leftarrow$ AutoModel(base_model)
+[…]
   # Initialize LoRA-SB with dataset D
-$\text{{sb\_model}}\leftarrow$ initSB(model, D)
+[…]
   # Standard training, only R trainable
-$\text{{trainer}}\leftarrow$ Trainer(sb_model, ...)
-$\text{{trainer}}.$train()
+[…]
+[…]
 4 Experiments
@@
 We fine-tune RoBERTa-large (Liu et al., 2019), Llama-3.2 3B (Dubey et al., 2024), Mistral-7B (Jiang et al., 2023), and Gemma-2 9B (Team et al., 2024), showcasing our method’s adaptability across a variety of tasks and model architectures.
-We implement all models using PyTorch (Paszke et al., 2019) and the popular HuggingFace Transformers library (Wolf et al., 2020).
+We use PyTorch (Paszke et al., 2019) and the HuggingFace Transformers library (Wolf et al., 2020) for our implementations.
 We run all experiments on a single NVIDIA A6000 GPU and report results as the average of three random seeds.
@@
 We provide a benchmark of this overhead in Section 5.
-These samples are randomly selected from the dataset’s training set in each run.
+The samples are randomly selected from the training set in each run.
+Detailed hyperparameter settings for all tasks are given in Appendix C.
+Baselines.
+We compare LoRA-SB against full FT, LoRA, LoRA-XS, rsLoRA, and PiSSA.
+rsLoRA introduces a rank-scaled stabilization factor ($\alpha/\sqrt{r}$) to enhance stability, while PiSSA updates only the principal components of the pre-trained weight $W$ and freezes the residuals.
 4.1 Arithmetic Reasoning
-Details.
-We fine-tune Mistral-7B (Jiang et al., 2023) and Gemma-2 9B (Team et al., 2024) on 50K samples from the MetaMathQA (Yu et al., 2024) dataset and evaluate them on the GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) benchmarks.
-Evaluation focuses solely on the final numeric answer.
-We apply LoRA modules to the key, value, query, attention output, and all fully connected weight matrices, training with ranks $r=\{32,64,96\}$.
-Detailed hyperparameter settings are provided in Appendix C.
-Results.
-We present the results for both models in Table 1.
-LoRA-SB significantly outperforms LoRA-XS across all parameter settings.
-Notably, LoRA-SB surpasses the performance of LoRA ($r=32$) while using 40x fewer trainable parameters for Mistral-7B and 90x fewer for Gemma-2 9B at ranks $r=96$ and $r=64$, respectively.
-We present training loss curves comparing LoRA-SB and LoRA-XS in Figure 2.
-Thanks to superior initialization, LoRA-SB starts with a lower initial loss compared to LoRA-XS.
-Additionally, due to optimal gradient approximation, LoRA-SB maintains a consistently better loss curve throughout training and converges to a superior final value.
-Table 1: Accuracy comparison of fine-tuning methods on Mistral-7B and Gemma-2 9B across the arithmetic reasoning benchmarks GSM8K and MATH, after training on MetaMathQA.
+Table 1: Accuracy comparison of FT methods on Mistral-7B and Gemma-2 9B across the arithmetic reasoning benchmarks GSM8K and MATH, after training on MetaMathQA.
 # Params denotes the number of trainable parameters.
@@
 LoRA $32$ $83.88$ M $61.94$ $15.98$
+rsLoRA $32$ $83.88$ M $62.15$ $16.24$
+PiSSA $32$ $83.88$ M $62.43$ $16.52$
 LoRA-XS $32$ $0.23$ M $54.28$ $13.36$
@@
 LoRA $32$ $108.04$ M $76.19$ $36.56$
+rsLoRA $32$ $108.04$ M $76.84$ $36.88$
+PiSSA $32$ $108.04$ M $77.12$ $37.04$
 LoRA-XS $32$ $0.30$ M $74.07$ $34.62$
@@
 LoRA-SB $96$ $2.71$ M $\mathbf{78.40}$ $\mathbf{37.70}$
+Details.
+We fine-tune Mistral-7B (Jiang et al., 2023) and Gemma-2 9B (Team et al., 2024) on 50K samples from the MetaMathQA (Yu et al., 2024) dataset and evaluate them on the GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) benchmarks.
+We apply LoRA modules to the key, value, query, attention output, and all fully connected weight matrices, training with ranks $r=\{32,64,96\}$.
+Results.
+We present results in Table 1.
+LoRA-SB significantly outperforms LoRA-XS across all settings.
+Notably, LoRA-SB outperforms LoRA-based methods ($r=32$) while using 40x fewer trainable parameters for Mistral-7B and 90x fewer for Gemma-2 9B at ranks $r=96$ and $r=64$, respectively.
+We present training loss curves comparing LoRA-SB and LoRA-XS in Figure 2.
+Thanks to superior initialization, LoRA-SB starts with a lower initial loss compared to LoRA-XS.
+Additionally, due to optimal gradient approximation, LoRA-SB maintains a consistently better loss curve throughout and converges to a superior final value.
 (a) Mistral-7B
@@
 Details.
-We fine-tune Llama-3.2 3B (Dubey et al., 2024) on CommonSense170K, a dataset that aggregates eight commonsense reasoning tasks (Hu et al., 2023).
+We fine-tune Llama-3.2 3B (Dubey et al., 2024) on CommonSense170K, a dataset with eight commonsense reasoning tasks (Hu et al., 2023).
 We evaluate the model’s performance on each dataset individually, which include BoolQ (Clark et al., 2019), SIQA (Sap et al., 2019), PIQA (Bisk et al., 2020), ARC-Challenge (Clark et al., 2018), ARC-Easy (Clark et al., 2018), OBQA (Mihaylov et al., 2018), WinoGrande (Sakaguchi et al., 2021), and HellaSwag (Zellers et al., 2019).
-Each example is framed as a multiple-choice question where the model generates the correct answer without explanations.
-We use the prompt template from (Hu et al., 2023).
 LoRA modules are applied to the key, value, query, attention output, and all fully connected weight matrices, training with ranks $r=\{32,64,96\}$.
-Hyperparameter details are provided in Appendix C.
 Results.
 We present the results in Table 2.
-LoRA-SB consistently outperforms LoRA-XS across all parameter settings.
-Additionally, LoRA-SB ($r=96$) exceeds the performance of LoRA ($r=32$) with 27x fewer trainable parameters.
-Table 2: Accuracy comparison of fine-tuning methods on Llama-3.2 3B across eight commonsense reasoning datasets.
+LoRA-SB consistently outperforms LoRA-XS across all settings.
+In addition, LoRA-SB ($r=96$) outperforms LoRA-based methods ($r=32$) with 27x fewer trainable parameters.
+Table 2: Accuracy comparison of FT methods on Llama-3.2 3B across eight commonsense reasoning datasets.
 # Params denotes the number of trainable parameters.
@@
 Llama-3.2 3B Full FT - $3.21$ B $70.43$ $85.64$ $80.45$ $91.92$ $85.02$ $88.52$ $75.29$ $81.88$ $82.39$
-LoRA $32$ $48.63$ M $70.03$ $\mathbf{85.20}$ $79.12$ $90.71$ $82.24$ $86.91$ $74.32$ $\mathbf{81.87}$ $81.44$
+LoRA $32$ $48.63$ M $70.03$ $85.20$ $79.12$ $90.71$ $82.24$ $86.91$ $74.32$ $\mathbf{81.87}$ $81.30$
+rsLoRA $32$ $48.63$ M $69.81$ $85.05$ $78.92$ $90.45$ $82.02$ $86.71$ $74.18$ $81.72$ $81.11$
+PiSSA $32$ $48.63$ M $70.12$ $\mathbf{85.42}$ $79.44$ $90.88$ $82.68$ $87.23$ $74.61$ $81.79$ $81.52$
 LoRA-XS $32$ $0.20$ M $65.01$ $82.87$ $76.17$ $87.32$ $80.12$ $84.78$ $70.31$ $75.71$ $77.79$
@@
 LoRA-SB $32$ $0.20$ M $66.33$ $84.06$ $78.91$ $89.04$ $81.37$ $86.62$ $72.44$ $76.97$ $79.47$
-LoRA-SB $64$ $0.80$ M $68.35$ $84.55$ $79.94$ $\mathbf{91.68}$ $83.03$ $87.84$ $74.83$ $80.12$ $81.29$
-LoRA-SB $96$ $1.81$ M $\mathbf{70.34}$ $84.76$ $\mathbf{80.19}$ $91.62$ $\mathbf{84.61}$ $\mathbf{87.92}$ $\mathbf{74.74}$ $81.20$ $\mathbf{81.92}$
+LoRA-SB $64$ $0.80$ M $68.35$ $84.55$ $79.94$ $\mathbf{91.68}$ $83.03$ $87.84$ $\mathbf{74.83}$ $80.12$ $81.29$
+LoRA-SB $96$ $1.81$ M $\mathbf{70.34}$ $84.76$ $\mathbf{80.19}$ $91.62$ $\mathbf{84.61}$ $\mathbf{87.92}$ $74.74$ $81.20$ $\mathbf{81.92}$
 4.3 Natural Language Understanding
 Details.
-We fine-tune RoBERTa-large (Liu et al., 2019) on GLUE, a sequence classification benchmark that contains several datasets, covering domains like sentiment analysis and natural language inference.
+We fine-tune RoBERTa-large (Liu et al., 2019) on GLUE, a popular language understanding classification benchmark that contains several datasets.
 The datasets we evaluate on are: CoLA (Warstadt et al., 2019), RTE, MRPC (Dolan & Brockett, 2005), SST-2 (Socher et al., 2013), QNLI (Rajpurkar et al., 2018), and STS-B (Cer et al., 2017).
-LoRA modules are applied only to the self-attention layers, following the configuration in the original LoRA paper (Hu et al., 2021), with ranks $r=\{8,16,24\}$.
-Detailed experimental settings are available in Appendix C.
+LoRA modules are applied only to the self-attention layers, with ranks $r=\{8,16,24\}$.
 Results.
 The results are shown in Table 3.
-LoRA-SB consistently outperforms LoRA-XS across all parameter configurations.
-Notably, LoRA-SB ($r=24$) outperforms LoRA ($r=8$) with 39x lesser trainable parameters.
-Table 3: Comparison of fine-tuning methods on RoBERTa-large across the GLUE benchmark datasets.
+LoRA-SB consistently outperforms LoRA-XS across all configurations.
+Additionally, LoRA-SB ($r=24$) outperforms LoRA-based methods ($r=8$) with 39x lesser trainable parameters.
+Table 3: Comparison of FT methods on RoBERTa-large across the GLUE benchmark datasets.
 # Params denotes the number of trainable parameters.
@@
 RoBERTa-large Full FT - $355.36$ M $68.44$ $83.42$ $90.21$ $91.76$ $93.92$ $96.21$ $87.33$
-LoRA 88 $2162.69$ K $68.02$ $82.98$ $90.05$ $91.43$ $93.42$ $95.98$ $86.98$
-LoRA-XS 88 $6.14$ K $61.07$ $75.23$ $86.21$ $89.29$ $92.44$ $94.72$ $83.16$
-LoRA-XS $16$ $24.57$ K $63.32$ $79.06$ $86.28$ $90.36$ $93.89$ $95.76$ $84.70$
+LoRA $8$ $2162.69$ K $68.02$ $82.98$ $90.05$ $91.43$ $93.42$ $95.98$ $86.98$
+rsLoRA $8$ $2162.69$ K $67.87$ $82.84$ $89.97$ $91.30$ $93.29$ $95.87$ $86.85$
+PiSSA $8$ $2162.69$ K $68.22$ $\mathbf{83.14}$ $90.10$ $91.59$ $93.55$ $96.03$ $87.10$
+LoRA-XS $8$ $6.14$ K $61.07$ $75.23$ $86.21$ $89.29$ $92.44$ $94.72$ $83.16$
+LoRA-XS $16$ $24.57$ K $63.32$ $79.06$ $86.28$ $90.36$ $93.69$ $95.76$ $84.70$
 LoRA-XS $24$ $55.20$ K $66.27$ $80.14$ $88.48$ $90.77$ $93.21$ $95.89$ $85.79$
-LoRA-SB 88 $6.14$ K $63.57$ $78.43$ $88.72$ $90.59$ $92.95$ $95.07$ $84.88$
-LoRA-SB $16$ $24.57$ K $64.36$ $82.31$ $89.71$ $91.24$ $93.89$ $95.87$ $86.23$
-LoRA-SB $24$ $55.20$ K $\mathbf{68.28}$ $\mathbf{83.03}$ $\mathbf{90.12}$ $\mathbf{91.65}$ $\mathbf{93.75}$ $\mathbf{96.11}$ $\mathbf{87.16}$
+LoRA-SB $8$ $6.14$ K $63.57$ $78.43$ $88.72$ $90.59$ $92.95$ $95.07$ $84.88$
+LoRA-SB $16$ $24.57$ K $64.36$ $82.31$ $89.71$ $91.24$ $\mathbf{93.89}$ $95.87$ $86.23$
+LoRA-SB $24$ $55.20$ K $\mathbf{68.28}$ $83.03$ $\mathbf{90.12}$ $\mathbf{91.65}$ $93.75$ $\mathbf{96.11}$ $\mathbf{87.16}$
 5 Analysis
 Optimal Initialization is Important!
-To isolate the impact of initialization on training, we perform truncated SVD on various matrices, including Kaiming initialization and $\Delta W_{avg}$ with varying levels of added Gaussian noise, as shown in Table 4.
+To isolate the impact of initialization, we take truncated SVD on various matrices, including Kaiming initialization (He et al., 2015) and $\Delta W_{avg}$ with varying levels of Gaussian noise, as shown in Table 4.
 By applying truncated SVD, we ensure optimal gradient approximation, leading to initialization matrices $B_{\text{init}}$ and $A_{\text{init}}$ that form orthonormal bases in $\mathbb{R}^{m}$ and $\mathbb{R}^{n}$, respectively.
@@
 Optimal Gradient Approximation is Important!
-We aim to examine the effect of optimal gradient approximation on training.
+We aim to examine the effect of optimal gradient approximation.
 Specifically, we want $B_{\text{init}}R_{\text{init}}A_{\text{init}}\approx\Delta W_{avg}$ without enforcing $B^{T}B=AA^{T}=I$.
-We achieve this through the following steps:
+We achieve this through:
 $U,S,V^{T}\leftarrow\textbf{SVD}(\Delta W_{avg})$ (14)
@@
 This construction ensures that $B_{\text{init}}R_{\text{init}}A_{\text{init}}\approx\Delta W_{avg}$, but only $AA^{T}=I$, while $B^{T}B\neq I$.
-This setup is suboptimal for gradient approximation since we do not explicity use the closed-form solution derived in Theorem 3.
+The setup is suboptimal for gradient approximation since we do not explicity use the closed-form solution derived in Theorem 3.
 We compare the resulting loss curves against LoRA-SB (which uses optimal gradient approximation) for Mistral-7B on MetaMathQA, as shown in Figure 3.
-Although both methods start similarly due to effective initialization, LoRA-SB converges to significantly better values, demonstrating the advantage of optimal gradient approximation.
-Furthermore, LoRA-SB achieves higher accuracies on the GSM8K and MATH benchmarks, with scores of $63.38$ and $17.44$ compared to $55.87$ and $12.74$, respectively.
+Although both start similarly due to effective initialization, LoRA-SB converges to significantly better values, demonstrating the advantage of optimal gradient approximation.
+Furthermore, LoRA-SB achieves higher accuracies on GSM8K and MATH, with scores of $63.38$ and $17.44$ compared to $55.87$ and $12.74$, respectively.
 Figure 3: Training loss for Mistral-7B on MetaMathQA, highlighting the impact of optimal gradient approximation.
-Training Time Overhead vs LoRA-XS
+Training Time Overhead vs LoRA-XS.
 As previously mentioned, we compute the update approximation using only $1/1000$ of the total training samples for each dataset.
 Table 5 presents the associated training time overhead for these computations, compared to LoRA-XS.
-The results show that the additional overhead is negligible, adding just 242–4 minutes compared to the total training time of 353–5 hours per epoch ($\approx 1.1\%$ to $1.3\%$).
+The results show that the additional overhead is negligible, adding just 24242–42 – 4 minutes compared to the total training time of 35353–53 – 5 hours per epoch ($\approx 1.1\%$ to $1.3\%$).
 Additionally, the update computation is performed only once, at the beginning of the first epoch, prior to training.
@@
 6 Conclusion
-In this work, we introduced LoRA-SB, a novel framework that bridges the gap between low-rank parameter-efficient fine-tuning and full fine-tuning.
-This is enabled by our carefully designed initialization strategy, which approximates the first step of full fine-tuning and ensures that the most relevant subspaces for task-specific adaptation are captured from the outset.
-Our theoretical analysis uncovered critical limitations in LoRA-XS: suboptimal gradient approximation and reliance on fixed initialization that fails to adapt to task-specific nuances.
-By addressing both issues through a unified approach, we demonstrate that LoRA-SB achieves optimal scaling and preserves critical update directions.
-Furthermore, our approach ensures hyperparameter independence by directly approximating the full fine-tuning gradient, thereby eliminating instability and convergence issues associated with scaling factors.
-Through extensive experiments on 4 models across 16 datasets covering mathematical reasoning, commonsense reasoning, and language understanding tasks, we demonstrate that our method exceeds the performance of LoRA while using upto 90x less parameters, and comprehensively outperforms LoRA-XS.
-Our work advances the state of PEFT while laying the groundwork for further innovations in low-rank adaptations for neural networks.
-Future work includes exploring adaptive layer-wise rank settings and evaluating the approach on other architectures, such as Vision Language Models (VLMs) and Vision Transformers (ViTs).
+In this work, we introduced LoRA-SB, which bridges the gap between low-rank PEFT and full FT.
+This is enabled by our initialization strategy, which approximates the first step of full FT and ensures that the most relevant subspaces for task-specific adaptation are captured.
+We achieve optimal gradient scaling and preserve update directions throughout training.
+Our approach ensures hyperparameter independence by approximating the full FT gradient, thereby eliminating instability issues associated with scaling factors.
+Through extensive experiments, we demonstrate that our method outperforms LoRA while using upto 90x less parameters, and comprehensively outperforms LoRA-XS.
+Our work advances PEFT while laying the groundwork for further innovations in low-rank adaptations for neural networks.
+Future work includes exploring adaptive layer-wise rank settings and integrating LoRA-SB with quantization.
+We also aim to evaluate its performance on other models, such as Vision Language Models (VLMs) and Vision Transformers (ViTs).
 7 Acknowledgements
@@
 Lemma.
-Let $\Delta W$ be an update learned by LoRA-XS .
-Then, the set of all $\Delta W$, say $\mathcal{W}_{LoRA-XS}$, can be given by $\mathcal{W}_{LoRA-XS}:=\{M\in\mathbb{R}^{m\times n}|\text{Col}(M)\subseteq\text{Col}(B)\wedge\text{ Row}(M)\subseteq\text{Row}(A)\}$
+Let $\Delta W$ be an update learned with LoRA-XS.
+Then, the set of all possible $\Delta W$, say $\mathcal{W}_{LoRA-XS}$, is given as: $\{M\in\mathbb{R}^{m\times n}\mid\text{Col}(M)\subseteq\text{Col}(B)\wedge\text{Row}(M)\subseteq\text{Row}(A)\},$ where for any $M\in\mathbb{R}^{m\times n}$ the column and row spaces are defined as $\text{Col}(M)=\{y\in\mathbb{R}^{m}\mid\exists x\in\mathbb{R}^{n}:\;y=Mx\}$ and $\text{Row}(M)=\{y\in\mathbb{R}^{n}\mid\exists x\in\mathbb{R}^{m}:\;y=M^{\top}x\}$.
 Proof.
-The column space of $\Delta W$ is given by:
-$\text{Col}(\Delta W)$ $=\{y\in\mathbb{R}^{m}\mid y=\Delta Wx,\,x\in\mathbb{R}^{n}\}$ (20)
-Now since $\Delta W=BRA$:
-$\text{Col}(\Delta W)=\{y\in\mathbb{R}^{m}\mid y=BRAx,\,x\in\mathbb{R}^{n}\}\implies\text{Col}(\Delta W)=\{y\in\mathbb{R}^{m}\mid y=Bz,\,z\in\text{Col}(RA)\}\dots\text{ for }z=RAx$ (21)
-Now:
-$\text{Col}(B)$ $=\{y\in\mathbb{R}^{m}\mid y=Bx,\,x\in\mathbb{R}^{r}\}$ (22)
-Now since $\text{Col}(RA)\subseteq\mathbb{R}^{r}$:
-$\text{Col}(\Delta W)$ $\subseteq\text{Col}(B)$ (23)
-A symmetric proof can now be given for row space of $\Delta W$. ∎
+Since $\Delta W=BRA$, we have
+$\text{Col}(\Delta W)=\{y\in\mathbb{R}^{m}\mid y=BRAx,\,x\in\mathbb{R}^{n}\}\implies\text{Col}(\Delta W)=\{y\in\mathbb{R}^{m}\mid y=Bz,\,z\in\text{Col}(RA)\}\subseteq\text{Col}(B).$
+That is, we proved that
+$\text{Col}(\Delta W)$ $\subseteq\text{Col}(B).$ (18)
+Following similar arguments, one can also show $\text{Row}(\Delta W)\subseteq\text{Row}(A)$. ∎
 A.2 Proof of Lemma 2
 Lemma.
-During LoRA-XS optimization, the gradient of the loss with respect to matrix $R$ can be expressed in terms of the gradient with respect to the weight matrix $W$ as: $g^{R}_{LoRA-XS}=sB^{T}gA^{T}$
+The gradient of the loss with respect to matrix $R$ can be expressed in terms of the gradient with respect to the weight matrix $W$ as: $g^{R}_{LoRA-XS}=sB^{\top}gA^{\top}$
 Proof.
 Let $L$ be the loss function. We have already defined $g$ and $g^{R}_{\text{LoRA-XS}}$ as:
-$g:=\frac{\partial L}{\partial W}\quad\&\quad g^{R}_{\text{LoRA-XS}}:=\frac{\partial L}{\partial R}$ (24)
-Now by chain rule:
-$\frac{\partial L}{\partial R}=\frac{\partial L}{\partial W}\frac{\partial W}{\partial R}\implies\frac{\partial L}{\partial R}=\frac{\partial L}{\partial W}\frac{\partial W}{\partial X}\frac{\partial X}{\partial R}\dots\text{ for }X=RA$ (25)
+$g:=\frac{\partial L}{\partial W}\quad\&\quad g^{R}_{\text{LoRA-XS}}:=\frac{\partial L}{\partial R}.$ (19)
+The chain rule gives
+$\frac{\partial L}{\partial R}=\frac{\partial L}{\partial W}\frac{\partial W}{\partial R}\implies\frac{\partial L}{\partial R}=\frac{\partial L}{\partial W}\frac{\partial W}{\partial X}\frac{\partial X}{\partial R}\quad\text{ for }X=RA$ (20)
 We know that for $W=sBX$:
-$\frac{\partial L}{\partial W}\frac{\partial W}{\partial X}=sB^{T}g\implies\frac{\partial L}{\partial R}=sB^{T}g\frac{\partial X}{\partial R}$ (26)
-Let $sB^{T}g=y$. We know that when $X=RA$:
-$y\frac{\partial X}{\partial R}=yA^{T}\implies\frac{\partial L}{\partial R}=yA^{T}=sB^{T}gA^{T}\$(27)
-$\text{Therefore, \quad}\boxed{g^{R}_{\text{LoRA-XS}}=sB^{T}gA^{T}}$ (28)
+$\frac{\partial L}{\partial W}\frac{\partial W}{\partial X}=sB^{\top}g\implies\frac{\partial L}{\partial R}=sB^{\top}g\frac{\partial X}{\partial R}$ (21)
+Let $sB^{\top}g=y$. We know that when $X=RA$:
+$y\frac{\partial X}{\partial R}=yA^{\top}\implies\frac{\partial L}{\partial R}=yA^{\top}=sB^{\top}gA^{\top}$ (22)
+[…]
 A.3 Proof of Theorem 3
 Theorem.
-[…]
+The optimal solution for the objective $\text{min}_{g^{R}}||\tilde{g}-g||^{2}_{F}$, such that $\tilde{g}=sBg^{R}A$, is: $g^{R}=\dfrac{1}{s^{2}}(B^{\top}B)^{-1}g^{R}_{LoRA-XS}(AA^{\top})^{-1}$ (24)
 Proof.
-Since we already defined the equivalent gradient $\tilde{g}:=sBg^{R}A$, the minimization problem can be formally denoted as:
+Since we already defined the equivalent gradient $\tilde{g}:=sBg^{R}A$, the minimization problem can be denoted as:
 $\operatorname*{arg\,min}_{g^{R}}F$ $=\|sBg^{R}A-g\|_{F}^{2}$ (25)
@@
 Using the same trick from before and substituting $g^{R}A=X$, we get:
-$2sB^{T}(sBg^{R}A-g)A^{T}=0\implies B^{T}(sBg^{R}A-g)A^{T}=0\implies B^{T}sBg^{R}AA^{T}=B^{T}gA^{T}$ (31)
+$2sB^{\top}(sBg^{R}A-g)A^{\top}=0\implies B^{\top}(sBg^{R}A-g)A^{\top}=0\implies B^{\top}sBg^{R}AA^{\top}=B^{\top}gA^{\top}$ (27)
 From Lemma 2, we get:
-$B^{T}gA^{T}=g^{R}_{\text{LoRA-XS}}/s\implies B^{T}sBg^{R}AA^{T}=g^{R}_{\text{LoRA-XS}}/s\implies B^{T}Bg^{R}AA^{T}=g^{R}_{\text{LoRA-XS}}/s^{2}$ (32)
-Now since $B$ and $A$ are full rank, multiplying both sides by $(B^{T}B)^{-1}$ and $(AA^{T})^{-1}$ on the left and right side respectively gives:
-$(B^{T}B)^{-1}(B^{T}Bg^{R}AA^{T})(AA^{T})^{-1}$ $=(B^{T}B)^{-1}g^{R}_{\text{LoRA-XS}}(AA^{T})^{-1}/s^{2}$ (33)
-$\text{Therefore,}\quad\boxed{g^{R}=\dfrac{1}{s^{2}}(B^{T}B)^{-1}g^{R}_{\text{LoRA-XS}}(AA^{T})^{-1}}$ (34)
+$B^{\top}gA^{\top}=g^{R}_{\text{LoRA-XS}}/s\implies B^{\top}sBg^{R}AA^{\top}=g^{R}_{\text{LoRA-XS}}/s\implies B^{\top}Bg^{R}AA^{\top}=g^{R}_{\text{LoRA-XS}}/s^{2}$ (28)
+Now since $B$ and $A$ are full rank, multiplying both sides by $(B^{\top}B)^{-1}$ and $(AA^{\top})^{-1}$ on the left and right side respectively gives:
+$(B^{\top}B)^{-1}(B^{\top}Bg^{R}AA^{\top})(AA^{\top})^{-1}$ $=(B^{\top}B)^{-1}g^{R}_{\text{LoRA-XS}}(AA^{\top})^{-1}/s^{2}$ (29)
+[…]
 A.4 Proof of Theorem 4
 Theorem.
-Consider the update for matrix $R$ using the closed-form solution derived in Theorem 3: $R\leftarrow R-\eta g^{R}$ where $\eta\geq 0$ is the learning rate.
-This update guarantees a reduction in the loss function, similar to traditional gradient descent methods.
-Specifically, the change in loss $\mathrm{d}L$ is given by: $\mathrm{d}L=-\eta{\langle g^{R}_{LoRA-XS},g^{R}\rangle}=-\eta{\langle g^{R}_{LoRA-XS},\frac{1}{s^{2}}(B^{T}B)^{-1}g^{R}_{LoRA-XS}(AA^{T})^{-1}\rangle}\leq 0$
+Consider the update for matrix $R$ using the solution derived in Theorem 3: $R\leftarrow R-\eta g^{R}$ where $\eta>0$ is the (sufficiently small) learning rate.
+This update guarantees a reduction in the loss $\Delta L$, given by: $\Delta L\coloneqq L(W_{0}+sB(R-\eta g^{R})A)-L(W_{0}+sBRA)=-\eta\langle g^{R}_{LoRA-XS},g^{R}\rangle_{F}+o(\eta)\leq 0$
 Proof.
-We establish that during optimization, the differential change in the loss function, $\mathrm{d}L$, can be expressed as:
-$\mathrm{d}L=-\eta\langle g^{R}_{\text{LoRA-XS}},\frac{1}{s^{2}}(B^{T}B)^{-1}g^{R}_{\text{LoRA-XS}}(AA^{T})^{-1}\rangle_{F}$ (35)
-Let us derive this expression step by step. First, the differential change in loss can be written as:
-$\mathrm{d}L=\langle\frac{\partial L}{\partial R},\mathrm{d}R\rangle_{F}$ (36)
-The update equation for $R$ follows:
-$\mathrm{d}R=-\eta g^{R}$ (37)
-Given that $\frac{\partial L}{\partial R}=g^{R}_{\text{LoRA-XS}}$, we obtain:
-$\mathrm{d}L$ $=-\eta\langle g^{R}_{\text{LoRA-XS}},g^{R}\rangle_{F}$ (38)
-$\mathrm{d}L$ $=-\eta\langle g^{R}_{\text{LoRA-XS}},\frac{1}{s^{2}}(B^{T}B)^{-1}g^{R}_{\text{LoRA-XS}}(AA^{T})^{-1}\rangle_{F}$
-To prove that $\mathrm{d}L\leq 0$, we demonstrate that:
-$\langle g^{R}_{\text{LoRA-XS}},\frac{1}{s^{2}}(B^{T}B)^{-1}g^{R}_{\text{LoRA-XS}}(AA^{T})^{-1}\rangle_{F}\geq 0$ (39)
-We first establish that $(B^{T}B)^{-1}$ is positive definite. For any non-zero vector $x$, since $B$ is full-rank:
-$\langle x,B^{T}Bx\rangle=\langle Bx,Bx\rangle=\|Bx\|^{2}>0$ (40)
-This proves $B^{T}B$, and consequently, $(B^{T}B)^{-1}$ are positive definite.
-Through Cholesky decomposition, we can express $(B^{T}B)^{-1}=VV^{T}$.
-Similarly, $(AA^{T})^{-1}$ is positive-definite and can be decomposed as $(AA^{T})^{-1}=UU^{T}$.
-Therefore:
-$\langle g^{R}_{\text{LoRA-XS}},\frac{1}{s^{2}}(B^{T}B)^{-1}g^{R}_{\text{LoRA-XS}}(AA^{T})^{-1}\rangle_{F}$ $=\frac{1}{s^{2}}\langle g^{R}_{\text{LoRA-XS}},VV^{T}g^{R}_{\text{LoRA-XS}}UU^{T}\rangle_{F}$ (41)
-$=\frac{1}{s^{2}}\langle V^{T}g^{R}_{\text{LoRA-XS}}U,V^{T}g^{R}_{\text{LoRA-XS}}U\rangle_{F}$
-$=\frac{1}{s^{2}}\|V^{T}g^{R}_{\text{LoRA-XS}}U\|_{F}^{2}\geq 0$
-Thus, we have proven:
-$\mathrm{d}L=-\eta\langle g^{R}_{\text{LoRA-XS}},\frac{1}{s^{2}}(B^{T}B)^{-1}g^{R}_{\text{LoRA-XS}}(AA^{T})^{-1}\rangle_{F}\leq 0$ (42)
-For our specific initialization where $(B^{T}B)=I$, $(AA^{T})=I$, and $s=1$, the result simplifies to:
-$\mathrm{d}L$ $=-\eta\langle g^{R}_{\text{LoRA-XS}},g^{R}_{\text{LoRA-XS}}\rangle_{F}\leq 0$ (43)
+Assuming that $L$ is differentiable, we use Taylor’s theorem and get
+$\Delta L$ $\coloneqq L(W_{0}+sB(R-\eta g^{R})A)-L(W_{0}+sBRA)$
+$=\left\langle\frac{\partial L}{\partial R},-\eta g^{R}\right\rangle_{F}+o(\eta)$
+$=-\frac{\eta}{s^{2}}\langle g^{R}_{\text{LoRA-XS}},(B^{\top}B)^{-1}g^{R}_{\text{LoRA-XS}}(AA^{\top})^{-1}\rangle_{F}+o(\eta),$ (31)
+where in the last step we also used the definition of $g^{R}_{\text{LoRA-XS}}$ and the result of Theorem 3. To prove $\Delta L\leq 0$ for small enough $\eta$, it is sufficient to show that
+$\langle g^{R}_{\text{LoRA-XS}},(B^{\top}B)^{-1}g^{R}_{\text{LoRA-XS}}(AA^{\top})^{-1}\rangle_{F}\geq 0.$ (32)
+Next, we note that matrices $B^{\top}B\in\mathbb{R}^{r\times r}$ and $AA^{\top}\in\mathbb{R}^{r\times r}$ are positive definite since they are positive semi-definite and matrices $B$ and $A$ are full-rank (i.e., with rank $r$) matrices, which means that $B^{\top}B$ and $AA^{\top}$ have non-zero eigenvalues.
+Therefore, $(B^{\top}B)^{-1}$ and $(AA^{\top})^{-1}$ are also positive definite, implying that there exist matrices $U$ and $V$ such that $(B^{\top}B)^{-1}=VV^{\top}$ and $(AA^{\top})^{-1}=UU^{\top}$ (e.g., one can find such matrices using Cholesky decomposition).
+Then, we have
+$\langle g^{R}_{\text{LoRA-XS}},(B^{\top}B)^{-1}g^{R}_{\text{LoRA-XS}}(AA^{\top})^{-1}\rangle_{F}$ $=\langle g^{R}_{\text{LoRA-XS}},VV^{\top}g^{R}_{\text{LoRA-XS}}UU^{\top}\rangle_{F}$
+$=\frac{1}{s^{2}}\langle V^{\top}g^{R}_{\text{LoRA-XS}}U,V^{\top}g^{R}_{\text{LoRA-XS}}U\rangle_{F}$
+$=\|V^{\top}g^{R}_{\text{LoRA-XS}}U\|_{F}^{2}\geq 0.$
+This concludes the proof. ∎
+For our specific initialization where $(B^{\top}B)=I$, $(AA^{\top})=I$, and $s=1$, the result simplifies to:
+$\Delta L$ $=-\eta\langle g^{R}_{\text{LoRA-XS}},g^{R}_{\text{LoRA-XS}}\rangle_{F}+o(\eta)\leq 0.$ (33)
 A.5 Proof of Theorem 5
 Theorem.
-The equivalent gradient $\tilde{g}$ is hyperparameter $s$ independent when $\tilde{g}=sBg^{R}A\dots\text{ but not when }\tilde{g}=sBg^{R}_{LoRA-XS}A$
+The equivalent gradient $\tilde{g}$ is hyperparameter $s$ independent when $\tilde{g}=sBg^{R}A\quad\text{ but not when }\quad\tilde{g}=sBg^{R}_{LoRA-XS}A$
 Proof.
 Let $g$ be the full fine-tuning gradient.
-We want to prove that $\tilde{g}$ does not depend on $s$, so we try to express it in terms of $g$ which does not depend on the LoRA-XS training process or reparameterization
+We want to prove that $\tilde{g}$ does not depend on $s$, so we try to express it in terms of $g$ which does not depend on the LoRA-XS training process or reparameterization.
 1) For $\tilde{g}=sBg^{R}A$:
-$g^{R}=\frac{1}{s^{2}}(B^{T}B)^{-1}g^{R}_{\text{LoRA-XS}}(AA^{T})^{-1}\implies\tilde{g}=\frac{s}{s^{2}}B(B^{T}B^{-1})g^{R}_{\text{LoRA-XS}}(AA^{T})^{-1}A$ (44)
-Now since $g^{R}_{\text{LoRA-XS}}=sB^{T}gA^{T}$:
-$\tilde{g}=\frac{1}{s}B(B^{T}B^{-1})sB^{T}gA^{T}(AA^{T})^{-1}A=B(B^{T}B^{-1})B^{T}gA^{T}(AA^{T})^{-1}A$ (45)
-which is $s$ independent!
+$g^{R}=\frac{1}{s^{2}}(B^{\top}B)^{-1}g^{R}_{\text{LoRA-XS}}(AA^{\top})^{-1}\implies\tilde{g}=\frac{s}{s^{2}}B(B^{\top}B^{-1})g^{R}_{\text{LoRA-XS}}(AA^{\top})^{-1}A$ (34)
+Now since $g^{R}_{\text{LoRA-XS}}=sB^{\top}gA^{\top}$:
+$\tilde{g}=\frac{1}{s}B(B^{\top}B^{-1})sB^{\top}gA^{\top}(AA^{\top})^{-1}A=B(B^{\top}B^{-1})B^{\top}gA^{\top}(AA^{\top})^{-1}A.$ (35)
+which is $s$-independent.
 2) For $\tilde{g}=sBg^{R}_{\text{LoRA-XS}}A$
-$g^{R}_{\text{LoRA-XS}}=sB^{T}gA^{T}\implies\tilde{g}=sB(sB^{T}gA^{T})A\implies\tilde{g}=s^{2}BB^{T}gA^{T}A$ (46)
-which is not $s$ independent! ∎
+$g^{R}_{\text{LoRA-XS}}=sB^{\top}gA^{\top}\implies\tilde{g}=sB(sB^{\top}gA^{\top})A\implies\tilde{g}=s^{2}BB^{\top}gA^{\top}A$ (36)
+which is not $s$-independent. ∎
 A.6 Proof of Theorem 6
 Theorem.
-If $A_{init}$ and $B_{init}$ are initialized using LoRA-SB for the first SGD optimizer, then $\Delta(B_{init}R_{init}A_{init})\approx\Delta W$
+If $A_{init}$ and $B_{init}$ are initialized using LoRA-SB for the first step of SGD optimizer, then $\Delta(B_{init}R_{init}A_{init})\approx\Delta W$
 Proof.
 Consider a gradient descent step with learning rate $\eta$ and updates for $R$:
-$\Delta R=-\eta\nabla_{R}\mathcal{L}(R)\implies B\Delta RA=-\eta B\nabla_{R}\mathcal{L}(R)A$ (47)
+$\Delta R=-\eta\nabla_{R}\mathcal{L}(R)\implies B\Delta RA=-\eta B\nabla_{R}\mathcal{L}(R)A.$ (37)
 To measure its approximation quality of update of the weights in full finetuning:
-$\Delta W$ $=-\eta\nabla_{W}\mathcal{L}(W_{0})$ (48)
+$\Delta W$ $=-\eta\nabla_{W}\mathcal{L}(W_{0}).$ (38)
 We use Frobenius norm of the difference between these two updates as a criterion:
-$\|B\Delta RA-\eta\mathcal{L}_{W}(W_{0})\|_{F}$ $=\eta\|B\nabla_{R}\mathcal{L}(R)A-\mathcal{L}_{W}(W_{0})\|_{F}$ (49)
+$\|B\Delta RA-\eta\nabla\mathcal{L}_{W}(W_{0})\|_{F}$ $=\eta\|B\nabla_{R}\mathcal{L}(R)A-\nabla\mathcal{L}_{W}(W_{0})\|_{F}.$ (39)
 We have shown before that:
-$\nabla_{R}\mathcal{L}$ $=B^{T}\nabla_{W}\mathcal{L}A^{T}$ (50)
-The problem becomes:
-$\min_{A_{\text{init}},B_{\text{init}}}\|B^{T}(B^{T}\nabla_{W}\mathcal{L}A^{T})A-\nabla_{W}\mathcal{L}\|_{F}\dots\text{ where }\nabla_{W}\mathcal{L}=USV^{T}$ (51)
+$\nabla_{R}\mathcal{L}$ $=B^{\top}\nabla_{W}\mathcal{L}A^{\top}.$ (40)
+The problem now becomes:
+$\min_{A_{\text{init}},B_{\text{init}}}\|B^{\top}(B^{\top}\nabla_{W}\mathcal{L}A^{\top})A-\nabla_{W}\mathcal{L}\|_{F}\quad\text{ where }\nabla_{W}\mathcal{L}=USV^{\top}.$ (41)
 Using our initialization, we get:
-$\|BB^{T}\nabla_{W}\mathcal{L}A^{T}A-\nabla_{W}\mathcal{L}\|_{F}$ $=\|U_{IR}U_{IR}^{T}UsV^{T}V_{IR}V_{IR}^{T}-UsV^{T}\|_{F}$ (52)
-Now:
-$U_{IR}U_{IR}^{T}UsV^{T}V_{IR}V_{IR}^{T}$ $=\sum_{i=1}^{r}\sigma_{i}u_{i}v_{i}^{T}$ (53)
-Now rank of $W^{\prime}$ where:
-$W^{\prime}$ $=U_{IR}U_{IR}^{T}UsV^{T}V_{IR}V_{IR}^{T}$ (54)
-is $\leq r$ since rank of $B_{\text{init}}$ and $A_{\text{init}}$ is $r$
-By Eckart-Young Theorem, the optimal low-rank optimization w.r.t Frobenius norm is:
-$W^{\prime*}$ $=\underset{\text{rank}(W^{\prime})=r}{\text{arg min}}\|W^{\prime}-\nabla_{W}\mathcal{L}\|_{F}=\sum_{i=1}^{r}\sigma_{i}u_{i}v_{i}^{T}$ (55)
-Since we get an identical expression, our solution is optimal. ∎
-Appendix B Inference Costs
+$\|BB^{\top}\nabla_{W}\mathcal{L}A^{\top}A-\nabla_{W}\mathcal{L}\|_{F}$ $=\|U_{IR}U_{IR}^{\top}USV^{\top}V_{IR}V_{IR}^{\top}-USV^{\top}\|_{F}.$ (42)
+Moreover, we also have
+$U_{IR}U_{IR}^{\top}USV^{\top}V_{IR}V_{IR}^{\top}$ $=\sum_{i=1}^{r}\sigma_{i}u_{i}v_{i}^{\top}.$ (43)
+The rank of $W^{\prime}$ such that
+$W^{\prime}$ $=U_{IR}U_{IR}^{\top}USV^{\top}V_{IR}V_{IR}^{\top}$ (44)
+is $\leq r$, since the corresponding ranks of $B_{\text{init}}$ and $A_{\text{init}}$ is $r$. Using the Eckart-Young Theorem, we find the optimal low-rank solution as:
+$W^{\prime*}$ $=\underset{\text{rank}(W^{\prime})=r}{\text{arg min}}\|W^{\prime}-\nabla_{W}\mathcal{L}\|_{F}=\sum_{i=1}^{r}\sigma_{i}u_{i}v_{i}^{\top}.$ (45)
+Since we also get an identical expression, our solution is optimal. ∎
+Appendix B Inference Overhead vs LoRA
 As discussed in Section 5, LoRA-SB introduces minimal inference cost overhead due to the insertion of the $r\times r$ matrix $R$ between $B$ and $A$, as well as the requirement for higher ranks to match LoRA’s performance.
@@
 Model Method Rank MACs FLOPs
-RoBERTa-large LoRA 88 $77.86$ G $155.79$ G
+RoBERTa-large LoRA $8$ $77.86$ G $155.79$ G
 LoRA-SB $16$ $78.42$ G $156.91$ G
@@
 In line with LoRA-XS settings, we fixed $\alpha$ at $16$ for their baseline.
+For all tasks, we followed the baseline configurations provided in the PiSSA (Meng et al., 2024) and rsLoRA (Kalajdzievski, 2023) papers for our comparisons.
+Table 7: Hyperparameter settings for training Mistral-7B and Gemma-2 9B on MetaMathQA, and Llama-3.2 3B on Commonsense170K.
 Mistral-7B / Gemma-2 9B Llama-3.2 3B
 Optimizer AdamW AdamW
-Batch size 11 66
+Batch size $1$ $6$
 Max. Seq. Len $512$ $256$
 Grad Acc. Steps $32$ $24$
-Epochs 11 22
-Dropout 0 $0.05$
+Epochs $1$ $2$
+Dropout 00 $0.05$
 Learning Rate $1\times 10^{-4}$ $2\times 10^{-3}$
@@
 Warmup Ratio $0.02$ $0.02$
-Table 7: Hyperparameter settings for training Mistral-7B and Gemma-2 9B on MetaMathQA, and Llama-3.2 3B on Commonsense170K.
+Table 8: Hyperparameter settings for RoBERTa-large on GLUE.
 CoLA RTE MRPC SST-2 QNLI STS-B
@@
 Epochs $30$ $30$ $30$ $15$ $15$ $30$
-Dropout 0
+Dropout 00
 Learning Rate $1\times 10^{-3}$
@@
 Warmup Ratio $0.06$
-Table 8: hyperparameter settings for RoBERTa-large on GLUE.
 Appendix D Dataset Details
@@
 We assess this dataset using two benchmarks: GSM8K (Cobbe et al., 2021), which consists of grade-school math problems requiring multi-step reasoning, and MATH (Hendrycks et al., 2021), which presents difficult, competition-level math problems.
-CommonSense170K is a comprehensive dataset that consolidates eight commonsense reasoning datasets (Hu et al., 2023), as described below:
+Evaluation focuses solely on the final numeric answer.
+CommonSense170K is a comprehensive dataset that consolidates eight commonsense reasoning datasets (Hu et al., 2023).
+Each example is framed as a multiple-choice question where the model generates the correct answer without explanations.
+We use the prompt template from (Hu et al., 2023).
+The individual datasets used are described below:
   1. 1.

### v4

@@
 Abstract
-Low-rank adapters have become standard for efficiently fine-tuning large language models (LLMs), but they often fall short of achieving the performance of full fine-tuning.
+Low-rank adapters have become standard for efficiently fine-tuning large language models, but they often fall short of achieving the performance of full fine-tuning.
 We propose a method, LoRA Silver Bullet or LoRA-SB, that approximates full fine-tuning within low-rank subspaces using a carefully designed initialization strategy.
-We theoretically demonstrate that the architecture of LoRA-XS—which inserts a learnable $r\times r$ matrix between $B$ and $A$ while keeping other matrices fixed—provides the precise conditions needed for this approximation.
-We leverage its constrained update space to achieve optimal scaling for high-rank gradient updates while removing the need for hyperparameter tuning.
+We theoretically demonstrate that the architecture of LoRA-XS, which inserts a learnable $r\times r$ matrix between $B$ and $A$ while keeping other matrices fixed, provides the precise conditions needed for this approximation.
+We leverage its constrained update space to achieve optimal scaling for high-rank gradient updates while removing the need for scaling factor tuning.
 We prove that our initialization offers an optimal low-rank approximation of the initial gradient and preserves update directions throughout training.
-Extensive experiments across mathematical reasoning, commonsense reasoning, and language understanding tasks demonstrate that our approach exceeds the performance of standard LoRA while using 27-90 times fewer learnable parameters, and comprehensively outperforms LoRA-XS.
-Our findings establish that it is possible to simulate full fine-tuning in low-rank subspaces, and achieve significant efficiency gains without sacrificing performance.
-Our code is publicly available at https://github.com/RaghavSinghal10/lora-sb.
-Machine Learning, ICML
+Extensive experiments across mathematical reasoning, commonsense reasoning, and language understanding tasks demonstrate that our approach exceeds the performance of LoRA (and baselines) while using 27-90 times fewer learnable parameters, and comprehensively outperforms LoRA-XS.
+Our findings establish that it is possible to simulate full fine-tuning in low-rank subspaces, and achieve significant parameter efficiency gains without sacrificing performance.
+Our code is publicly available at: https://github.com/CERT-Lab/lora-sb.
 1 Introduction
-Pre-trained language models have become central to natural language processing, achieving state-of-the-art performance across diverse tasks (Radford et al., 2021; Kirillov et al., 2023; Achiam et al., 2023).
-While these models excel at general-purpose capabilities (Bubeck et al., 2023; Hao et al., 2022), adapting them to specific downstream tasks often requires fine-tuning (FT).
-Although in-context learning (Brown et al., 2020; Radford et al., 2019) has gained popularity for its simplicity, it falls short in both performance and efficiency compared to FT (Liu et al., 2022).
+Pre-trained language models have become central to natural language processing, achieving state-of-the-art performance across diverse tasks (35, 21, 1).
+While these models excel at general-purpose capabilities (4, 14), adapting them to specific downstream tasks often requires fine-tuning (FT).
 At the same time, full FT, while highly effective, is computationally expensive and impractical at scale.
 Parameter-efficient fine-tuning (PEFT) has become vital for adapting large language models (LLMs) under computational constraints.
-Low-rank methods like LoRA (Hu et al., 2021) address this by reducing learnable parameters via low-rank updates, sparking advancements in optimization, initialization, structured matrices, and adaptive rank selection (Zhang et al., 2023; Wang et al., 2024b, a).
-However, these methods face trade-offs: either retain many parameters to match full FT or sacrifice performance for extreme efficiency (Hu et al., 2021; Ding et al., 2023; Wang et al., 2024b).
+Low-rank methods like LoRA (17) address this by reducing learnable parameters via low-rank updates, sparking advancements in optimization, initialization, structured matrices, and adaptive rank selection (52, 46, 45).
+However, these methods face trade-offs: either retain many parameters to match full FT or sacrifice performance for extreme efficiency (17, 10, 46).
 This raises a critical question: Can we design low-rank methods that achieve full FT-level performance while drastically reducing parameter counts?
 Figure 1: LoRA-SB.
-LoRA-XS (Bałazy et al., 2024) reduces parameters compared to LoRA (Hu et al., 2021) by inserting a learnable $r\times r$ matrix $R$ between $B$ and $A$, while keeping other matrices fixed, leading to $W=W_{0}+sBRA$.
+LoRA-XS (2) reduces parameters compared to LoRA (17) by inserting a learnable $r\times r$ matrix $R$ between $B$ and $A$, while keeping other matrices fixed, leading to $W=W_{0}+sBRA$.
 Our method, LoRA-SB, uses the same architecture.
@@
 We initialize $B$, $R$, and $A$ such that the equivalent gradient $\tilde{g}_{SB}$ provably best approximates the full FT gradient $g$ in low rank subspaces at each step.
-In essence, we simulate the entire full FT process optimally within low-rank subspaces by utilizing only the first gradient $g_{1}$ (shown in green) from full FT.
+In essence, we simulate the entire full FT process optimally within low-rank subspaces by utilizing only the first full FT gradient $g_{1}$.
 Low-rank decomposition methods operate on a fundamental premise: FT requires learning only a low-rank update to the pre-trained weights.
-Some theoretical work extends this hypothesis, suggesting that methods like LoRA can learn any low-rank approximation of the full FT gradient.
 However, the gradients computed by these methods do not inherently possess this property.
-For instance, LoRA’s gradients need explicit optimization at each step to better approximate the full FT gradient (Wang et al., 2024b).
-Additionally, initialization has emerged as a critical factor in low-rank adaptation, as highlighted by recent works like Pissa-LoRA (Meng et al., 2024) and LoRA-GA (Wang et al., 2024a).
-We analyze these limitations in the context of the architecture of LoRA-XS (Bałazy et al., 2024)—which inserts a learnable $r\times r$ matrix between $B$ and $A$ while keeping other matrices fixed—and demonstrate that these challenges are even more pronounced.
+For instance, LoRA’s gradients need explicit optimization at each step to better approximate the full FT gradient (46).
+Additionally, initialization has emerged as a critical factor in low-rank adaptation, as highlighted by recent works like PiSSA-LoRA (30) and LoRA-GA (45).
+We analyze these limitations in the context of the architecture of LoRA-XS (2), which inserts a learnable $r\times r$ matrix between $B$ and $A$ while keeping other matrices fixed, and demonstrate that these challenges are even more pronounced.
 While exploring solutions inspired by LoRA-based methods, we discover a remarkable property unique to LoRA-XS: through careful initialization of $A$ and $B$, we can simulate the full FT optimization in low rank subspaces through entire training, as shown in Figure 1.
-Our initialization provides optimal scaling for approximating high-rank full FT gradients and eliminates need for scaling the hyperparameter $\alpha$.
+Our initialization provides optimal scaling for approximating high-rank full FT gradients and eliminates need for tuning the hyperparameter $\alpha$.
+The peak memory usage of LoRA-SB never exceeds that of LoRA or other baselines, and its training-time overhead relative to LoRA is negligible ($\approx 1.1\%-1.3\%$).
 Our key contributions are:
-  • We theoretically formalize the limitations of LoRA-XS, showing how its constrained update space leads to suboptimal gradient approximation, initialization sensitivity, and hyperparameter dependence.
-  • We propose an initialization strategy derived from approximating the first step of full FT, proving it provides optimal low-rank approximation of the initial gradient and preserves update directions throughout training.
-  • We prove that our initialization makes gradient optimization hyperparameter-independent and guarantees convergence by maintaining orthonormal bases, eliminating the need for any tuning of the scaling factor.
-  • Through extensive experiments on $4$ models across $16$ datasets covering mathematical reasoning, commonsense reasoning, and language understanding tasks, we demonstrate that our method surpasses LoRA’s performance while using 27-90x less learnable parameters, and comprehensively outperforms LoRA-XS.
-2 Related Work
-Parameter-Efficient Fine-Tuning (PEFT).
-PEFT methods have become essential for adapting large pre-trained models under computational constraints.
-Early techniques like AdapterFusion (Pfeiffer et al., 2021) and Prefix-Tuning (Li & Liang, 2021) enabled task-specific adaptation with minimal parameter updates.
-Advances like soft prompts (Lester et al., 2021) further reduced trainable parameter counts while maintaining strong performance.
-Recent approaches have explored operating directly on model representations (Wu et al., 2024).
-Low-Rank Decomposition Methods.
-LoRA (Hu et al., 2021) demonstrated that weight updates during FT could be efficiently approximated using low-rank matrices, drastically reducing parameter counts.
-Building on this insight, variants such as QLoRA (Dettmers et al., 2023) and AdaLoRA (Zhang et al., 2023) extended the paradigm through quantization and adaptive allocation strategies.
-The applicability of low-rank techniques has also been explored in pretraining with GaLore (Zhao et al., 2024) and ReLoRA (Lialin et al., 2023), highlighting the versatility of low-rank adaptation methods.
-LoRA-based methods have also been applied in other domains, such as efficient federated FT (Sun et al., 2024; Singhal et al., 2025).
-Enhancing LoRA Performance.
-Recent efforts have focused on optimizing LoRA’s performance.
-PiSSA (Meng et al., 2024) demonstrated improvements by initializing matrices with principal components of pre-trained weights.
-LoRA-Pro (Wang et al., 2024b) and LoRA-GA (Wang et al., 2024a) improved gradient approximation, aligning low-rank updates more closely with full FT.
-Methods like DoRA (Liu et al., 2024) and rsLoRA (Kalajdzievski, 2023) introduced decomposition-based and scaling stabilization techniques to enhance learning stability and expand LoRA’s utility.
-Improving Efficiency in LoRA Variants.
-Efficiency-focused innovations have pushed LoRA toward more parameter savings.
-LoRA-XS (Bałazy et al., 2024) achieves this by inserting a small trainable weight matrix into frozen low-rank matrices.
-VeRA (Kopiczko et al., 2024) shares low-rank matrices across layers, relying on scaling vectors for task-specific adaptation.
-Tied-LoRA (Renduchintala et al., 2024) leverages weight tying to reduce parameter usage at higher ranks, while HydraLoRA (Tian et al., 2024) introduces an asymmetric architecture for improvement.
+  • We formalize the limitations of LoRA-XS, showing how its constrained update space leads to suboptimal gradient approximation, initialization sensitivity, and scaling dependence.
+  • We propose an initialization strategy derived from using the first step of full FT, which provides an optimal approximation of the initial gradient and preserves update directions throughout.
+  • We prove our initialization makes gradient optimization scaling-independent and guarantees convergence by maintaining orthonormal bases, eliminating need for tuning the scaling factor $\alpha$.
+  • Through extensive experiments on $4$ models across $16$ datasets covering mathematical reasoning, commonsense reasoning, and language understanding, we demonstrate that LoRA-SB surpasses LoRA while using 27-90x less learnable parameters, and comprehensively outperforms LoRA-XS.
 2 Methodology
@@
 where $B$ and $A$ are fixed, and only $R\in\mathbb{R}^{r\times r}$ is trainable, reducing the number of parameters to $r^{2}$.
-We denote the full FT gradient as $g=\frac{\partial L}{\partial W}$, and the LoRA-XS gradient as $g^{R}_{\text{LoRA-XS}}=\frac{\partial L}{\partial R}$, where $L$ is the loss function.
+We denote the full FT gradient: $g=\frac{\partial L}{\partial W}$; LoRA-XS gradient: $g^{R}_{\text{LoRA-XS}}=\frac{\partial L}{\partial R}$; $L$ is the loss function.
 2.2 Motivation
-LoRA-XS (Bałazy et al., 2024) has significantly fewer learnable parameters than LoRA but performs suboptimally.
+LoRA-XS (2) has significantly fewer learnable parameters than LoRA but performs suboptimally.
 LoRA-XS’s architecture causes constraints on the type of updates it can learn.
@@
 Let $\Delta W$ be an update learned with LoRA-XS.
-Then, the set of all possible $\Delta W$, say $\mathcal{W}_{LoRA-XS}$, is given as: $\{M\in\mathbb{R}^{m\times n}|\text{Col}(M)\subseteq\text{Col}(B)$ $\wedge\text{Row}(M)\subseteq\text{Row}(A)\},$ where $\text{Col}(M)$ and $\text{Row}(M)$ are column and row spaces of matrix $M$ respectively.
+Then, the set of all possible $\Delta W$, say $\mathcal{W}_{LoRA-XS}$, is given as: $\mathcal{W}_{LoRA-XS}=\{M\in\mathbb{R}^{m\times n}|\text{Col}(M)\subseteq\text{Col}(B)\wedge\text{Row}(M)\subseteq\text{Row}(A)\},$ where $\text{Col}(M)$ and $\text{Row}(M)$ are column and row spaces of matrix $M$ respectively.
 Proof.
-See Appendix A.1. ∎
+See Appendix B.1. ∎
 We identify three key limitations, which arise due to this and otherwise:
@@
 The gradient of LoRA does not optimally approximate the full gradient, and needs to be tuned at each step.
-LoRA-Pro (Wang et al., 2024b) finds that this results in suboptimal performances, and provides a closed form solution to optimize the gradients.
+LoRA-Pro (46) finds that this results in suboptimal performances, and provides a closed form solution to optimize the gradients.
 In LoRA-XS, the gradient updates are restricted to an even more constrained low-rank space since $A$ and $B$ are fixed.
@@
 Consider the zero initialization of the $B$ matrix, for example.
-While LoRA may experience some performance degradation in this case (Wang et al., 2024a; Meng et al., 2024), the ideal low-rank update $\Delta W$ can still be reached through gradient descent.
-In fact, zero initialization for the $B$ matrix is commonly used, including in the original LoRA paper (Hu et al., 2021).
+While LoRA may experience some performance degradation in this case (45, 30), the ideal low-rank update $\Delta W$ can still be reached through gradient descent.
+In fact, zero initialization for the $B$ matrix is commonly used, including in the original LoRA paper (17).
 However, in LoRA-XS, this results in no learning, as the product $BRA$ remains zero.
-LoRA-XS uses the most significant subspaces spanned by the columns of pre-trained weights for initialization, inspired by Meng et al.
-(2024).
+LoRA-XS uses the most significant subspaces spanned by the columns of pre-trained weights for initialization, inspired by PiSSA (30).
 This initialization is not aligned well with FT because it fails to capture the specific subspaces relevant to the FT task.
-3) Hyperparameter Sensitivity: The scaling factor $s$, present in almost every LoRA based FT method requires tuning to maintain stability during training.
+3) Scaling Factor Sensitivity: The scaling factor $s$, present in almost every LoRA based method, requires tuning to maintain stability during training.
 This factor acts as a bridge between the low-rank and full-rank spaces, compensating for the dimensional mismatch in gradients.
-Poor tuning of $s$ can lead to unstable training or slow convergence, e.g., see rsLoRA (Kalajdzievski, 2023), adding complexity and potentially limiting practical deployment.
+Poor tuning of $s$ can lead to unstable training or slow convergence (rsLoRA (20)), adding complexity and potentially limiting practical deployment.
 2.3 Approximation of the full FT gradient
-As mentioned, LoRA optimization is mathematically equivalent to full FT using a constrained low-rank gradient.
+As mentioned, LoRA optimization is equivalent to full FT using a constrained low-rank gradient.
 However, the update generated using the gradients of LoRA does not result in the same update which the low-rank gradient would have generated.
 The following holds true for LoRA-XS as well.
-To understand this, let us look at the change in weight $W$ and its relationship with changing of low-rank matrix $R$, which can be simply given by $\mathrm{d}W=-sB(\mathrm{d}R)A.$ This implies that updating $R$ with gradient $g^{R}$ is equivalent to updating $W$ with low rank equivalent gradient $\tilde{g}$ in full FT as described in Definition 1.
+To understand this, let us look at the change in weight $W$ and its relationship with changing of low-rank matrix $R$, which can be simply given by $\mathrm{d}W=-sB(\mathrm{d}R)A.$ This implies that updating $R$ with gradient $g^{R}$ is equivalent to updating $W$ with low rank equivalent gradient $\tilde{g}$ in full FT (Definition 1).
 Definition 1.
-We define the equivalent gradient as: $\tilde{g}=sBg^{R}A$ where $g^{R}$ is the gradient of $L$ with respect to $R$.
+We define the equivalent gradient in LoRA-XS as: $\tilde{g}=sBg^{R}A$, where $g^{R}$ is the gradient of $L$ with respect to $R$.
 The equivalent gradient describes the virtual low-rank gradient of matrix $W$ in LoRA-XS optimization process, despite $W$ not being directly trainable.
@@
 Lemma 2.
-The gradient of the loss with respect to matrix $R$ can be expressed in terms of the gradient with respect to the weight matrix $W$ as: $g^{R}_{LoRA-XS}=sB^{\top}gA^{\top}$
+The gradient of the loss with respect to matrix $R$ can be expressed in terms of the gradient with respect to the weight matrix $W$ as: $g^{R}_{LoRA-XS}=sB^{\top}gA^{\top}$.
 Proof.
-See Appendix A.2. ∎
-We can now formulate our objective to minimize the distance between the equivalent gradient and the full gradient.
+See Appendix B.2. ∎
+We now formulate our objective to minimize the distance between the equivalent gradient and the full gradient.
 We do not have access to the full FT gradient $g$ during LoRA-XS based FT.
@@
 Theorem 3.
-The optimal solution for the objective $\text{min}_{g^{R}}||\tilde{g}-g||^{2}_{F}$, such that $\tilde{g}=sBg^{R}A$, is: $g^{R}=\dfrac{1}{s^{2}}(B^{\top}B)^{-1}g^{R}_{LoRA-XS}(AA^{\top})^{-1}$ (4)
+For full-rank AA and $B$ matrices, the optimal solution for the objective
+[…]
 Proof.
-See Appendix A.3. ∎
-The closed-form solution in Theorem 3 solves the optimization problem $\text{min}_{g^{R}}||\tilde{g}-g||^{2}_{F}$, but by itself doesn’t ensure the loss will decrease when updating $R$.
-Through Theorem 4, we prove that the change in loss is non-positive ($\Delta L\leq 0$).
-This property is fundamental to optimization as it guarantees consistent loss minimization throughout training.
+See Appendix B.3. ∎
+[…]
 Theorem 4.
-Consider the update for matrix $R$ using the solution derived in Theorem 3: $R\leftarrow R-\eta g^{R}$ where $\eta>0$ is the (sufficiently small) learning rate.
+Consider the update for matrix $R$ using the solution derived in Theorem 3:
+$R\leftarrow R-\eta g^{R}$, where $\eta>0$ is the (sufficiently small) learning rate.
 This update guarantees a reduction in the loss $\Delta L$, given by: $\Delta L=-\eta\langle g^{R}_{LoRA-XS},g^{R}\rangle_{F}+o(\eta)\leq 0.$
 Proof.
-See Appendix A.4. ∎
+See Appendix B.4. ∎
 2.4 Initialization using update approximation
@@
 $\Delta W_{avg}=-\eta\textbf{sign}(\sum_{i=0}^{n\leq|\mathbb{X}|}\nabla_{W}\mathcal{L}(W_{0},x_{i})),\quad x_{i}\in\mathbb{X}$ (6)
-This better captures the general direction of adaptation required for the target task while being less sensitive to individual sample variations.
-We can then use truncated SVD to obtain a low-rank approximation of $\Delta W_{\text{avg}}$, and express it as $sBRA$.
+Since AdamW is used as the optimizer for both full FT and LoRA-SB training, we approximate its first update step using the sign of the summed gradients rather than their raw values (see Appendix C for details).
+This better captures the direction of adaptation required for the target task while being less sensitive to individual sample variations.
+We then use truncated SVD to obtain a low-rank approximation of $\Delta W_{\text{avg}}$, and express it as $sBRA$.
 There exist infinite combinations of $B$ and $A$ which can obey this relationship.
@@
 This is equivalent to the $B$ and $A$ initialization in LoRA-XS but by approximating the update rather than the pre-trained matrix.
-We note that the above process can be computed for any optimizer, by approximating the corresponding first step.
+The above process can be computed for any optimizer, by approximating the corresponding first step.
 We compute this specifically for AdamW since we use it.
-3.5 Hyperparameter independence
-The hyperparameter $\alpha$ is used in LoRA and other decomposition-based method to tackle the issue of instability caused to improper scaling of the updates.
+2.5 Scaling Factor independence
+The hyperparameter $\alpha$ is used in LoRA and other decomposition-based methods to tackle instability caused to improper scaling of the updates.
 The gradient scaling is accounted for, by adding a hyperparameter to normalize the updates.
-The importance of scaling is shown in methods like rank stabilization (Kalajdzievski, 2023).
+The importance of scaling is shown in methods like rank stabilization (20).
 However, the full FT gradient $g$ needs no such tuning.
@@
 Theorem 5.
-The equivalent gradient $\tilde{g}$ is hyperparameter $s$ independent for $\tilde{g}=sBg^{R}A$, but not for $\tilde{g}=sBg^{R}_{LoRA-XS}A$
+The equivalent gradient $\tilde{g}$ is hyperparameter $s$ independent for $\tilde{g}=sBg^{R}A$, but not for $\tilde{g}=sBg^{R}_{LoRA-XS}A$.
 Proof.
-See Appendix A.5. ∎
-The hyperparameter independence of the equivalent gradient eliminates the need for manual gradient scaling.
+See Appendix B.5. ∎
+The scaling factor independence of the equivalent gradient eliminates the need for manual gradient scaling.
 Updates to $W$ depend solely on this gradient (modulo learning rate), making any additional scaling redundant.
@@
 2.6 LoRA-SB: Update approximation initialization is a silver bullet
-The solutions discussed above independently address the gradient approximation and initialization problems, while also providing hyperparameter independence.
-Our proposed method, LoRA-SB, elegantly combines these solutions through a simple initialization strategy, derived from approximating the first step of full FT:
+The solutions discussed independently address the gradient approximation and initialization problems, while also providing scaling factor independence.
+LoRA-SB, elegantly combines these solutions through a simple initialization strategy, derived from approximating the first full FT step:
 $U,S,V^{\top}\leftarrow\textbf{SVD}(\Delta W_{avg})$ (7)
-$B_{init}\leftarrow U[1:r]$ (9)
-$A_{init}\leftarrow V[1:r]$ (10)
-$R_{init}\leftarrow\dfrac{1}{s}S[1:r,1:r]$ (11)
-By the Eckart-Young theorem (Eckart & Young, 1936; Mirsky, 1960), this gives the optimal rank-$r$ approximation of the full FT update.
+$B_{init}\leftarrow U[1:r],A_{init}\leftarrow V[1:r],R_{init}\leftarrow\dfrac{1}{s}S[1:r,1:r]$ (8)
+By the Eckart-Young theorem (13, 32), this gives the optimal rank-$r$ approximation of the full FT update.
 where $U$, $S$, $V$ are obtained from truncated SVD of the averaged first update $\Delta W_{\text{avg}}$.
@@
 With fixed $B$ and $A$ matrices being orthonormal, the need for complex matrix inversions during training is eliminated, , as the optimal update step, derived in Equation 3, simplifies to:
-$g^{R}=\dfrac{1}{s^{2}}(B^{\top}B)^{-1}g^{R}_{LoRA-XS}(AA^{\top})^{-1}=\frac{1}{s^{2}}g^{R}_{LoRA-XS}.$
-Optimal Update Approximation. Our initialization guarantees that the first update optimally approximates the full FT weight updates:
-$sB_{\text{init}}R_{\text{init}}A_{\text{init}}\approx\Delta W_{avg}.$
-By the Eckart-Young theorem, this gives the optimal rank-$r$ approximation of the initial full FT update.
-Hyperparameter Independence. As shown in Theorem 5, when gradient approximation is applied with orthonormal $B$ and $A$, the hyperparameter $s$ can be set to 1, resulting in:
+$g^{R}=\dfrac{1}{s^{2}}(B^{\top}B)^{-1}g^{R}_{LoRA-XS}(AA^{\top})^{-1}=\frac{1}{s^{2}}g^{R}_{LoRA-XS}$
+Optimal Update Approximation.
+Our initialization guarantees that the first update optimally approximates the full FT weight updates: $sB_{\text{init}}R_{\text{init}}A_{\text{init}}\approx\Delta W_{avg}.$ By the Eckart-Young theorem, this is the optimal rank-$r$ approximation of the initial full FT update.
+Scaling Factor Independence.
+As shown in Theorem 5, when gradient approximation is applied with orthonormal $B$ and $A$, the hyperparameter $s$ can be set to 1, resulting in guaranteed optimal gradient approximation at every step, without requiring any scaling factor:
 $\boxed{g^{R}=g^{R}_{\text{LoRA-XS}}}$ (9)
-This demonstrates that our initialization guarantees optimal gradient approximation at every step, without requiring any scaling factor.
 Guaranteed Loss Reduction.
-Since $B$ is a tall orthonormal matrix and $A$ is a wide orthonormal matrix, they remain full rank throughout training.
-This ensures that $dL$ remains negative 4, guaranteeing stable optimization and convergence.
-Another heuristic which might lead to a good initialization is setting the weights $B$ and $A$, such that they match the first update also approximately matches the direction of $\Delta W$.
+Since $B$ is a tall orthonormal and $A$ a wide orthonormal matrix, they remain full rank throughout training.
+This ensures that $dL$ remains negative (Theorem 4), guaranteeing stable optimization and convergence.
 $\Delta(sB_{init}R_{init}A_{init})\approx\gamma\Delta W$ (10)
+Another heuristic which might lead to a good initialization is setting $B$ and $A$, such that the first update also approximately matches the $\Delta W$ direction (Equation 10).
 Thankfully, we don’t have to choose between the two.
-For SGD, we prove that setting $B_{init}$ and $A_{init}$ using Equations 8-11, results in the first update of LoRA-XS to best approximate the direction of the update of full FT (Theorem 6).
+For SGD, we prove that setting $B_{init}$ and $A_{init}$ using Equations 7-8, results in the first update of LoRA-XS to best approximate the direction of the full FT update (Theorem 6).
 Theorem 6.
-If $A_{init}$ and $B_{init}$ are initialized using LoRA-SB for the first step of SGD optimizer, then $\Delta(B_{init}R_{init}A_{init})\approx\Delta W$
+If $A_{init}$ and $B_{init}$ are initialized using LoRA-SB for the first step of SGD optimizer, then the update given by LoRA-SB, $\Delta(B_{init}R_{init}A_{init})$ , is the best low-rank approximation of full fine-tuning update, $\Delta W$.
 Proof.
-See Appendix A.6. ∎
+See Appendix B.6. ∎
+While Theorem 6 is stated for SGD, the result extends to other SGD-based optimizers such as AdamW.
+In practice, we use AdamW and approximate the first update by taking the sign of the averaged gradients, consistent with AdamW’s first-step behavior.
+This produces an initialization whose SVD still yields the optimal rank-$r$ approximation of the simulated full FT update.
+Initialization Memory.
+To optimize GPU memory during initialization, we hook into the backward pass and compute the gradients layerwise, immediately discarding the computed gradients (29, 45).
+This ensures $O(1)$ memory usage, independent of the number of layers, keeping GPU memory well within limits.
+This guarantees that the memory required for LoRA-SB initialization never exceeds the memory needed for subsequent LoRA-SB fine-tuning, and that the peak memory usage of the entire LoRA-SB algorithm never exceeds that of standard LoRA and other baselines.
 LoRA-SB Advantages over LoRA.
@@
 Even if $B$ and $A$ are initialized as orthonormal in LoRA, subsequent updates do not preserve this property because $B$ and $A$ are trainable.
-This results in several challenges:
-  • Potential instability of $(B^{\top}B)^{-1}$ and $(AA^{\top})^{-1}$, as they are not guaranteed to remain non-singular during training.
-  • Inability to ensure consistent loss reduction due to potential rank deficiency—$B$ and $A$ may not remain full-rank throughout training.
-  • Necessity to fine-tune the hyperparameter $\alpha$.
+This results in several challenges in using LoRA (even with optimal gradient approximation) compared to LoRA-SB:
+  • Potential instability of $(B^{\top}B)^{-1}$ and $(AA^{\top})^{-1}$, not guaranteed to remain non-singular throughout.
+  • Inability to ensure consistent loss reduction due to potential rank deficiency, $B$ and $A$ may not remain full-rank throughout training.
+  • Necessity to fine-tune the scaling factor hyperparameter $\alpha$.
   • Repeated re-computation of $B^{\top}B$ and $AA^{\top}$ is required at each optimizer step for accurate gradient approximation.
-Algorithm.
-We present a PyTorch-like implementation of our method in Algorithm 1.
-To optimize GPU memory usage during initialization, we hook into the backward pass of PyTorch and compute the gradients layerwise, immediately discarding the computed gradients (Lv et al., 2024; Wang et al., 2024a).
-This ensures $O(1)$ memory usage, independent of the number of layers, keeping memory consumption well within manageable limits and ensuring it does not exceed the memory requirements of subsequent LoRA-SB FT.
-For large batch sizes, memory usage can be further optimized through gradient accumulation and quantization.
-Algorithm 1 LoRA-SB, PyTorch-like
-  def initSB(model, D)
-       # Estimate gradient with n samples
-       $\mathbf{\Delta W_{avg}}\leftarrow$ est_grad(model, D, n)
-       # Initialize B, R, A
-       $\mathbf{B,R,A}\leftarrow$ trunc_SVD($\mathbf{\Delta W_{avg}}$)
-       # Convert to LoRA-SB model
-[…]
-       return sb_model
-  # Load pre-trained model
-[…]
-  # Initialize LoRA-SB with dataset D
-[…]
-  # Standard training, only R trainable
-[…]
-[…]
 3 Experiments
-We evaluate our method over $16$ different datasets on three widely-used NLP benchmarks, using models ranging from the 355 M-parameter RoBERTa-large model to the 9 B-parameter Gemma-2 model.
-Our setup spans both masked and autoregressive architectures, allowing us to comprehensively assess the effectiveness of our approach.
-We fine-tune RoBERTa-large (Liu et al., 2019), Llama-3.2 3B (Dubey et al., 2024), Mistral-7B (Jiang et al., 2023), and Gemma-2 9B (Team et al., 2024), showcasing our method’s adaptability across a variety of tasks and model architectures.
-We use PyTorch (Paszke et al., 2019) and the HuggingFace Transformers library (Wolf et al., 2020) for our implementations.
-We run all experiments on a single NVIDIA A6000 GPU and report results as the average of three random seeds.
-To save memory, we initialize base models in torch.bfloat16 precision.
-Appendix D provides detailed information on the datasets used.
-We compute the update approximation using only $\mathbf{1/1000}$ of each dataset’s total number of samples.
-This ensures that the additional training time overhead is minimal and has a negligible effect on overall efficiency.
-We provide a benchmark of this overhead in Section 5.
-The samples are randomly selected from the training set in each run.
-Detailed hyperparameter settings for all tasks are given in Appendix C.
-Baselines.
-We compare LoRA-SB against full FT, LoRA, LoRA-XS, rsLoRA, and PiSSA.
-rsLoRA introduces a rank-scaled stabilization factor ($\alpha/\sqrt{r}$) to enhance stability, while PiSSA updates only the principal components of the pre-trained weight $W$ and freezes the residuals.
-4.1 Arithmetic Reasoning
-Table 1: Accuracy comparison of FT methods on Mistral-7B and Gemma-2 9B across the arithmetic reasoning benchmarks GSM8K and MATH, after training on MetaMathQA.
-# Params denotes the number of trainable parameters.
-The best results among PEFT methods are highlighted in bold.
-Model Method Rank # Params Accuracy ($\uparrow$)
-GSM8K MATH
-Mistral-7B Full FT - $7.24$ B $63.87$ $17.65$
-LoRA $32$ $83.88$ M $61.94$ $15.98$
-rsLoRA $32$ $83.88$ M $62.15$ $16.24$
-PiSSA $32$ $83.88$ M $62.43$ $16.52$
-LoRA-XS $32$ $0.23$ M $54.28$ $13.36$
-LoRA-XS $64$ $0.92$ M $57.08$ $15.62$
-LoRA-XS $96$ $2.06$ M $58.53$ $16.42$
-LoRA-SB $32$ $0.23$ M $58.91$ $15.28$
-LoRA-SB $64$ $0.92$ M $60.73$ $16.28$
-LoRA-SB $96$ $2.06$ M $\mathbf{63.38}$ $\mathbf{17.44}$
-Gemma-2 9B Full FT - $9.24$ B $79.23$ $38.02$
-LoRA $32$ $108.04$ M $76.19$ $36.56$
-rsLoRA $32$ $108.04$ M $76.84$ $36.88$
-PiSSA $32$ $108.04$ M $77.12$ $37.04$
-LoRA-XS $32$ $0.30$ M $74.07$ $34.62$
-LoRA-XS $64$ $1.20$ M $75.02$ $36.46$
-LoRA-XS $96$ $2.71$ M $75.21$ $36.98$
-LoRA-SB $32$ $0.30$ M $75.44$ $36.66$
-LoRA-SB $64$ $1.20$ M $76.65$ $37.14$
-LoRA-SB $96$ $2.71$ M $\mathbf{78.40}$ $\mathbf{37.70}$
-Details.
-We fine-tune Mistral-7B (Jiang et al., 2023) and Gemma-2 9B (Team et al., 2024) on 50K samples from the MetaMathQA (Yu et al., 2024) dataset and evaluate them on the GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) benchmarks.
-We apply LoRA modules to the key, value, query, attention output, and all fully connected weight matrices, training with ranks $r=\{32,64,96\}$.
-Results.
-We present results in Table 1.
-LoRA-SB significantly outperforms LoRA-XS across all settings.
-Notably, LoRA-SB outperforms LoRA-based methods ($r=32$) while using 40x fewer trainable parameters for Mistral-7B and 90x fewer for Gemma-2 9B at ranks $r=96$ and $r=64$, respectively.
-We present training loss curves comparing LoRA-SB and LoRA-XS in Figure 2.
-Thanks to superior initialization, LoRA-SB starts with a lower initial loss compared to LoRA-XS.
-Additionally, due to optimal gradient approximation, LoRA-SB maintains a consistently better loss curve throughout and converges to a superior final value.
-(a) Mistral-7B
-(b) Gemma-2 9B
-Figure 2: Training loss curves for Mistral-7B and Gemma-2 9B on MetaMathQA, comparing LoRA-SB and LoRA-XS.
-4.2 Commonsense Reasoning
-Details.
-We fine-tune Llama-3.2 3B (Dubey et al., 2024) on CommonSense170K, a dataset with eight commonsense reasoning tasks (Hu et al., 2023).
-We evaluate the model’s performance on each dataset individually, which include BoolQ (Clark et al., 2019), SIQA (Sap et al., 2019), PIQA (Bisk et al., 2020), ARC-Challenge (Clark et al., 2018), ARC-Easy (Clark et al., 2018), OBQA (Mihaylov et al., 2018), WinoGrande (Sakaguchi et al., 2021), and HellaSwag (Zellers et al., 2019).
-LoRA modules are applied to the key, value, query, attention output, and all fully connected weight matrices, training with ranks $r=\{32,64,96\}$.
-Results.
-We present the results in Table 2.
-LoRA-SB consistently outperforms LoRA-XS across all settings.
-In addition, LoRA-SB ($r=96$) outperforms LoRA-based methods ($r=32$) with 27x fewer trainable parameters.
-Table 2: Accuracy comparison of FT methods on Llama-3.2 3B across eight commonsense reasoning datasets.
-# Params denotes the number of trainable parameters.
-The best results among PEFT methods are highlighted in bold.
-Model Method Rank # Params Accuracy ($\uparrow$)
+We evaluate over $16$ different datasets on $3$ widely-used benchmarks, using models ranging from the 355 M RoBERTa-large model to the 9 B Gemma-2 model.
+Our setup spans both masked and autoregressive architectures, allowing us to comprehensively assess the effectiveness of LoRA-SB.
+Specifically, we fine-tune RoBERTa-large (27), Llama-3.2 3B (12), Mistral-7B (19), and Gemma-2 9B (43).
+We compute the update approximation using only $\mathbf{1/1000}$ ($0.1\%$) of each dataset’s total size.
+This ensures that the training time overhead is minimal and has a negligible effect on efficiency.
+Detailed hyperparameter and dataset details are given in Appendix H and I, respectively.
+Baselines. We compare LoRA-SB against full FT, LoRA (17), LoRA-XS (2), and several popular variants of LoRA - rsLoRA (20), PiSSA (30), DoRA (26), and LoRA-Pro (46).
+Table 1: Comparison of FT methods on Mistral-7B and Gemma-2 9B across arithmetic benchmarks. # Params denotes the number of trainable parameters. Best results among PEFT methods are in bold.
+Method Rank Mistral-7B Gemma-2 9B
+# Params GSM8K ($\uparrow$) MATH ($\uparrow$) # Params GSM8K ($\uparrow$) MATH ($\uparrow$)
+Full FT - $7.24$ B $63.87$ $17.65$ $9.24$ B $79.23$ $38.02$
+LoRA 32 $83.88$ M $61.94$ $15.98$ $108.04$ M $76.19$ $36.56$
+rsLoRA 32 $83.88$ M $62.15$ $16.24$ $108.04$ M $76.84$ $36.88$
+PiSSA 32 $83.88$ M $62.43$ $16.52$ $108.04$ M $77.12$ $37.04$
+DoRA 32 $85.26$ M $62.65$ $16.64$ $109.88$ M $77.58$ $37.04$
+LoRA-Pro 32 $83.88$ M $63.07$ $17.32$ $108.04$ M $78.26$ $37.53$
+LoRA-XS 32 $0.23$ M $54.28$ $13.36$ $0.30$ M $74.07$ $34.62$
+LoRA-XS 64 $0.92$ M $57.08$ $15.62$ $1.20$ M $75.02$ $36.46$
+LoRA-XS 96 $2.06$ M $58.53$ $16.42$ $2.71$ M $75.21$ $36.98$
+LoRA-SB 32 $0.23$ M $58.91$ $15.28$ $0.30$ M $75.44$ $36.66$
+LoRA-SB 64 $0.92$ M $60.73$ $16.28$ $1.20$ M $76.65$ $37.14$
+LoRA-SB 96 $2.06$ M $\mathbf{63.38}$ $\mathbf{17.44}$ $2.71$ M $\mathbf{78.40}$ $\mathbf{37.70}$
+Table 2: Comparison of FT methods on Llama-3.2 3B across eight commonsense reasoning datasets. # Params denotes the number of trainable parameters. Best results among PEFT methods are in bold.
+Method Rank # Params Accuracy ($\uparrow$)
 BoolQ PIQA SIQA HellaS. WinoG. ARC-e ARC-c OBQA Avg.
-Llama-3.2 3B Full FT - $3.21$ B $70.43$ $85.64$ $80.45$ $91.92$ $85.02$ $88.52$ $75.29$ $81.88$ $82.39$
+Full FT - $3.21$ B $70.43$ $85.64$ $80.45$ $91.92$ $85.02$ $88.52$ $75.29$ $81.88$ $82.39$
 LoRA $32$ $48.63$ M $70.03$ $85.20$ $79.12$ $90.71$ $82.24$ $86.91$ $74.32$ $\mathbf{81.87}$ $81.30$
-rsLoRA $32$ $48.63$ M $69.81$ $85.05$ $78.92$ $90.45$ $82.02$ $86.71$ $74.18$ $81.72$ $81.11$
-PiSSA $32$ $48.63$ M $70.12$ $\mathbf{85.42}$ $79.44$ $90.88$ $82.68$ $87.23$ $74.61$ $81.79$ $81.52$
+rsLoRA $32$ $48.63$ M $69.81$ $85.63$ $78.92$ $90.45$ $82.02$ $86.71$ $74.18$ $81.72$ $81.11$
+PiSSA $32$ $48.63$ M $70.12$ $85.42$ $79.44$ $90.88$ $82.68$ $87.23$ $74.61$ $81.79$ $81.52$
+DoRA $32$ $49.40$ M $70.43$ $85.63$ $79.68$ $90.76$ $82.90$ $87.61$ $74.87$ $82.04$ $81.74$
+LoRA-Pro $32$ $48.63$ M $\mathbf{71.28}$ $\mathbf{85.81}$ $79.35$ $90.90$ $83.42$ $87.24$ $\mathbf{75.32}$ $81.74$ $81.88$
 LoRA-XS $32$ $0.20$ M $65.01$ $82.87$ $76.17$ $87.32$ $80.12$ $84.78$ $70.31$ $75.71$ $77.79$
@@
 LoRA-SB $32$ $0.20$ M $66.33$ $84.06$ $78.91$ $89.04$ $81.37$ $86.62$ $72.44$ $76.97$ $79.47$
-LoRA-SB $64$ $0.80$ M $68.35$ $84.55$ $79.94$ $\mathbf{91.68}$ $83.03$ $87.84$ $\mathbf{74.83}$ $80.12$ $81.29$
-LoRA-SB $96$ $1.81$ M $\mathbf{70.34}$ $84.76$ $\mathbf{80.19}$ $91.62$ $\mathbf{84.61}$ $\mathbf{87.92}$ $74.74$ $81.20$ $\mathbf{81.92}$
-4.3 Natural Language Understanding
-Details.
-We fine-tune RoBERTa-large (Liu et al., 2019) on GLUE, a popular language understanding classification benchmark that contains several datasets.
-The datasets we evaluate on are: CoLA (Warstadt et al., 2019), RTE, MRPC (Dolan & Brockett, 2005), SST-2 (Socher et al., 2013), QNLI (Rajpurkar et al., 2018), and STS-B (Cer et al., 2017).
-LoRA modules are applied only to the self-attention layers, with ranks $r=\{8,16,24\}$.
-Results.
-The results are shown in Table 3.
-LoRA-SB consistently outperforms LoRA-XS across all configurations.
-Additionally, LoRA-SB ($r=24$) outperforms LoRA-based methods ($r=8$) with 39x lesser trainable parameters.
-Table 3: Comparison of FT methods on RoBERTa-large across the GLUE benchmark datasets.
+LoRA-SB $64$ $0.80$ M $68.35$ $84.55$ $79.94$ $\mathbf{91.68}$ $83.03$ $87.84$ $74.83$ $80.12$ $81.29$
+LoRA-SB $96$ $1.81$ M $70.34$ $84.76$ $\mathbf{80.19}$ $91.62$ $\mathbf{84.61}$ $\mathbf{87.92}$ $74.74$ $81.20$ $\mathbf{81.92}$
+Table 3: Comparison of FT methods on RoBERTa-large across GLUE datasets.
 # Params denotes the number of trainable parameters.
-The best results among PEFT methods are highlighted in bold.
+Best results among PEFT methods are in bold.
 We use Pearson correlation for STS-B, Matthew’s correlation for CoLA, and accuracy for others.
-Higher is better for each metric.
-Model Method Rank # Params CoLA RTE MRPC STS-B QNLI SST-2 All
+Method Rank # Params CoLA RTE MRPC STS-B QNLI SST-2 All
 Mcc $\uparrow$ Acc $\uparrow$ Acc $\uparrow$ Corr $\uparrow$ Acc $\uparrow$ Acc $\uparrow$ Avg. $\uparrow$
-RoBERTa-large Full FT - $355.36$ M $68.44$ $83.42$ $90.21$ $91.76$ $93.92$ $96.21$ $87.33$
-LoRA $8$ $2162.69$ K $68.02$ $82.98$ $90.05$ $91.43$ $93.42$ $95.98$ $86.98$
-rsLoRA $8$ $2162.69$ K $67.87$ $82.84$ $89.97$ $91.30$ $93.29$ $95.87$ $86.85$
-PiSSA $8$ $2162.69$ K $68.22$ $\mathbf{83.14}$ $90.10$ $91.59$ $93.55$ $96.03$ $87.10$
-LoRA-XS $8$ $6.14$ K $61.07$ $75.23$ $86.21$ $89.29$ $92.44$ $94.72$ $83.16$
+Full FT - $355.36$ M $68.44$ $83.42$ $90.21$ $91.76$ $93.92$ $96.21$ $87.33$
+LoRA 88 $2162.69$ K $68.02$ $82.98$ $90.05$ $91.43$ $93.42$ $95.98$ $86.98$
+rsLoRA 88 $2162.69$ K $67.87$ $82.84$ $89.97$ $91.30$ $93.29$ $95.87$ $86.85$
+PiSSA 88 $2162.69$ K $68.22$ $83.14$ $90.10$ $91.59$ $93.55$ $96.03$ $87.10$
+DoRA 88 $2260.99$ K $68.05$ $83.04$ $89.93$ $91.34$ $93.11$ $95.82$ $86.88$
+LoRA-Pro 88 $2162.69$ K $67.98$ $\mathbf{83.40}$ $\mathbf{90.49}$ $91.38$ $93.37$ $95.98$ $87.10$
+LoRA-XS 88 $6.14$ K $61.07$ $75.23$ $86.21$ $89.29$ $92.44$ $94.72$ $83.16$
 LoRA-XS $16$ $24.57$ K $63.32$ $79.06$ $86.28$ $90.36$ $93.69$ $95.76$ $84.70$
 LoRA-XS $24$ $55.20$ K $66.27$ $80.14$ $88.48$ $90.77$ $93.21$ $95.89$ $85.79$
-LoRA-SB $8$ $6.14$ K $63.57$ $78.43$ $88.72$ $90.59$ $92.95$ $95.07$ $84.88$
+LoRA-SB 88 $6.14$ K $63.57$ $78.43$ $88.72$ $90.59$ $92.95$ $95.07$ $84.88$
 LoRA-SB $16$ $24.57$ K $64.36$ $82.31$ $89.71$ $91.24$ $\mathbf{93.89}$ $95.87$ $86.23$
-LoRA-SB $24$ $55.20$ K $\mathbf{68.28}$ $83.03$ $\mathbf{90.12}$ $\mathbf{91.65}$ $93.75$ $\mathbf{96.11}$ $\mathbf{87.16}$
+LoRA-SB $24$ $55.20$ K $\mathbf{68.28}$ $83.03$ $90.12$ $\mathbf{91.65}$ $93.75$ $\mathbf{96.11}$ $\mathbf{87.16}$
+3.1 Arithmetic Reasoning
+We fine-tune Mistral-7B (19) and Gemma-2 9B (43) on 50K samples from MetaMathQA (50) and evaluate on GSM8K (8) and MATH (16).
+We apply LoRA modules to the key, value, query, attention output, and all fully connected weight matrices, training with ranks $r=\{32,64,96\}$.
+We present results in Table 1.
+LoRA-SB significantly outperforms LoRA-XS across all settings.
+LoRA-SB outperforms LoRA-based methods ($r=32$) while using 40x fewer trainable parameters for Mistral-7B and 90x fewer for Gemma-2 9B at ranks $r=96$ and $r=64$, respectively.
+We present training loss curves comparing LoRA-SB and LoRA-XS in Figure 2.
+Thanks to superior initialization, LoRA-SB starts with a lower initial loss compared to LoRA-XS.
+Further, due to optimal gradient approximation, LoRA-SB maintains a consistently better loss throughout and converges to a superior final value.
+(a) Mistral-7B
+(b) Gemma-2 9B
+Figure 2: Training loss curves for Mistral-7B and Gemma-2 9B, comparing LoRA-SB and LoRA-XS.
+3.2 Commonsense Reasoning
+We fine-tune Llama-3.2 3B (12) on CommonSense170K, a dataset with eight commonsense reasoning tasks (18).
+LoRA modules are applied to the key, value, query, attention output, and all fully connected weight matrices, training with ranks $r=\{32,64,96\}$.
+We present the results in Table 2.
+LoRA-SB consistently outperforms LoRA-XS across all settings.
+In addition, LoRA-SB ($r=96$) outperforms LoRA-based methods ($r=32$) with 27x fewer trainable parameters.
+3.3 Natural Language Understanding
+We fine-tune RoBERTa-large (27) on GLUE, a popular language understanding benchmark.
+LoRA modules are applied only to the self-attention layers, with ranks $r=\{8,16,24\}$.
+Results are shown in Table 3.
+LoRA-SB consistently outperforms LoRA-XS across all settings.
+Additionally, LoRA-SB ($r=24$) outperforms LoRA-based methods ($r=8$) with 39x lesser trainable parameters.
 4 Analysis
 Optimal Initialization is Important!
-To isolate the impact of initialization, we take truncated SVD on various matrices, including Kaiming initialization (He et al., 2015) and $\Delta W_{avg}$ with varying levels of Gaussian noise, as shown in Table 4.
+To isolate the impact of initialization, we take truncated SVD on various matrices, including Kaiming initialization (15) and $\Delta W_{avg}$ with varying levels of Gaussian noise, as shown in Table 4.
 By applying truncated SVD, we ensure optimal gradient approximation, leading to initialization matrices $B_{\text{init}}$ and $A_{\text{init}}$ that form orthonormal bases in $\mathbb{R}^{m}$ and $\mathbb{R}^{n}$, respectively.
 This results in $B^{T}B=AA^{T}=I$, allowing us to isolate the effect of initialization.
-The results clearly demonstrate the significance of initialization—our approach consistently outperforms other variants.
+The results clearly demonstrate the significance of initialization, our approach consistently outperforms other variants.
 Table 4: Comparison of initialization strategies using Mistral-7B on GSM8K and MATH. All methods ensure optimal gradient approximation, with differences arising solely from the initialization.
@@
 LoRA-SB; trunc_SVD ($\Delta W_{avg}$) $\mathbf{63.38}$ $\mathbf{17.44}$
+Why Do We Use 0.1% of the Dataset Size for Initialization?
+We selected the $0.1\%$ initialization dataset-size heuristic based on experiments that suggested it provides a good tradeoff between quality and efficiency.
+Specifically, we conducted ablations varying the number of samples used for initialization when fine-tuning Mistral-7B and Gemma-2 9B on 50k samples from MetaMathQA.
+The results (Table 5) show that once the sample count exceeds a modest threshold (25 samples or $0.05\%$), performance quickly plateaus, indicating that the learned subspace is already sufficiently representative.
+Using $0.1\%$ of the training data (50 samples) consistently exceeds this threshold across tasks and models, while incurring negligible training time overhead.
+Table 5: Performance effect of number of samples used for initialization.
+# Samples Mistral-7B Gemma-2 9B
+GSM8K ($\uparrow$) MATH ($\uparrow$) GSM8K ($\uparrow$) MATH ($\uparrow$)
+11 $62.13$ $15.55$ $76.03$ $35.77$
+55 $62.78$ $16.86$ $77.49$ $37.24$
+$25$ $63.28$ $17.30$ $78.18$ $37.70$
+$50$ $63.38$ $17.44$ $78.40$ $37.70$
+$100$ $63.34$ $17.25$ $78.22$ $37.45$
+$200$ $63.45$ $17.36$ $78.43$ $37.87$
+$500$ $63.40$ $17.52$ $78.54$ $37.63$
 Optimal Gradient Approximation is Important!
@@
 $U,S,V^{T}\leftarrow\textbf{SVD}(\Delta W_{avg})$ (11)
-$B_{\text{init}}\leftarrow U[1:r]S[1:r,1:r]$ (15)
-$A_{\text{init}}\leftarrow V[1:r]$ (16)
-$R_{\text{init}}\leftarrow I$ (17)
-This construction ensures that $B_{\text{init}}R_{\text{init}}A_{\text{init}}\approx\Delta W_{avg}$, but only $AA^{T}=I$, while $B^{T}B\neq I$.
+$B_{\text{init}}\leftarrow U[1:r]S[1:r,1:r],A_{\text{init}}\leftarrow V[1:r],R_{\text{init}}\leftarrow I$ (12)
+This ensures that $B_{\text{init}}R_{\text{init}}A_{\text{init}}\approx\Delta W_{avg}$, but only $AA^{T}=I$, while $B^{T}B\neq I$.
 The setup is suboptimal for gradient approximation since we do not explicity use the closed-form solution derived in Theorem 3.
-We compare the resulting loss curves against LoRA-SB (which uses optimal gradient approximation) for Mistral-7B on MetaMathQA, as shown in Figure 3.
+We compare the resulting loss curves against LoRA-SB (which uses optimal gradient approximation) for Mistral-7B, as shown in Figure 3 in Appendix E.
 Although both start similarly due to effective initialization, LoRA-SB converges to significantly better values, demonstrating the advantage of optimal gradient approximation.
 Furthermore, LoRA-SB achieves higher accuracies on GSM8K and MATH, with scores of $63.38$ and $17.44$ compared to $55.87$ and $12.74$, respectively.
-Figure 3: Training loss for Mistral-7B on MetaMathQA, highlighting the impact of optimal gradient approximation.
-Training Time Overhead vs LoRA-XS.
-As previously mentioned, we compute the update approximation using only $1/1000$ of the total training samples for each dataset.
-Table 5 presents the associated training time overhead for these computations, compared to LoRA-XS.
-The results show that the additional overhead is negligible, adding just 24242–42 – 4 minutes compared to the total training time of 35353–53 – 5 hours per epoch ($\approx 1.1\%$ to $1.3\%$).
-Additionally, the update computation is performed only once, at the beginning of the first epoch, prior to training.
-Table 5: Training time overhead due to the initialization for various models on their respective tasks.
-Model Overhead Training Time/Epoch
-Mistral 7B 0:02:01 3:03:57
-Gemma-2 9B 0:03:46 4:13:24
-Llama-3.2 3B 0:03:54 4:54:31
-Inference Overhead vs LoRA
-LoRA-SB introduces a minimal inference cost overhead due to the insertion of the $r\times r$ matrix $R$ between $B$ and $A$, and the need for higher ranks to achieve comparable performance to LoRA.
-We benchmark the inference-time FLOPs and MACs across various models and find that the overhead is negligible.
-A detailed analysis can be found in Table 6 of Appendix B.
+Training Time and Inference.
+We provide detailed benchmarks of training time and inference performance in Appendix F and G, respectively.
+As shown, the initialization step in LoRA-SB introduces only a negligible training-time overhead compared to LoRA ($\approx 1.1\%-1.3\%$).
 5 Conclusion
@@
 We achieve optimal gradient scaling and preserve update directions throughout training.
-Our approach ensures hyperparameter independence by approximating the full FT gradient, thereby eliminating instability issues associated with scaling factors.
-Through extensive experiments, we demonstrate that our method outperforms LoRA while using upto 90x less parameters, and comprehensively outperforms LoRA-XS.
-Our work advances PEFT while laying the groundwork for further innovations in low-rank adaptations for neural networks.
-Future work includes exploring adaptive layer-wise rank settings and integrating LoRA-SB with quantization.
-We also aim to evaluate its performance on other models, such as Vision Language Models (VLMs) and Vision Transformers (ViTs).
+Our approach ensures scaling factor independence by approximating the full FT gradient, thereby eliminating potential instability issues.
+Through extensive experiments, we demonstrate that our method outperforms LoRA (and baselines) using upto 90x less parameters, and comprehensively outperforms LoRA-XS.
 6 Acknowledgements
 This research was supported by funding from Mohamed bin Zayed University of Artificial Intelligence (MBZUAI) and ADIA Lab.
-Appendix A Mathematical Proofs
+Appendix
+Appendix A Related Work
+Parameter-Efficient Fine-Tuning (PEFT).
+PEFT methods have become essential for adapting large pre-trained models under computational constraints.
+Early techniques like AdapterFusion (34) and Prefix-Tuning (24) enabled task-specific adaptation with minimal parameter updates.
+Advances like soft prompts (23) further reduced trainable parameter counts while maintaining strong performance.
+Recent approaches have explored operating directly on model representations (49).
+Low-Rank Decomposition Methods.
+LoRA (17) demonstrated that weight updates during FT could be efficiently approximated using low-rank matrices, drastically reducing parameter counts.
+Building on this insight, variants such as QLoRA (9) and AdaLoRA (52) extended the paradigm through quantization and adaptive allocation strategies.
+The applicability of low-rank techniques has also been explored in pretraining with GaLore (53) and ReLoRA (25), highlighting the versatility of low-rank adaptation methods.
+LoRA-based methods have also been applied in other domains, such as efficient federated FT (42, 40).
+Enhancing LoRA Performance.
+Recent efforts have focused on optimizing LoRA’s performance.
+PiSSA (30) demonstrated improvements by initializing matrices with principal components of pre-trained weights.
+LoRA-Pro (46) and LoRA-GA (45) improved gradient approximation, aligning low-rank updates more closely with full FT.
+Methods like DoRA (26) and rsLoRA (20) introduced decomposition-based and scaling stabilization techniques to enhance learning stability and expand LoRA’s utility.
+Improving Efficiency in LoRA Variants.
+Efficiency-focused innovations have pushed LoRA toward more parameter savings.
+LoRA-XS (2) achieves this by inserting a small trainable weight matrix into frozen low-rank matrices.
+VeRA (22) shares low-rank matrices across layers, relying on scaling vectors for task-specific adaptation.
+Tied-LoRA (37) leverages weight tying to reduce parameter usage at higher ranks, while HydraLoRA (44) introduces an asymmetric architecture for improvement.
+Appendix B Proofs
 In all the proofs below, we will use the notations defined in Section 2.
-A.1 Proof of Lemma 1
+B.1 Proof of Lemma 1
 Lemma.
 Let $\Delta W$ be an update learned with LoRA-XS.
-Then, the set of all possible $\Delta W$, say $\mathcal{W}_{LoRA-XS}$, is given as: $\{M\in\mathbb{R}^{m\times n}\mid\text{Col}(M)\subseteq\text{Col}(B)\wedge\text{Row}(M)\subseteq\text{Row}(A)\},$ where for any $M\in\mathbb{R}^{m\times n}$ the column and row spaces are defined as $\text{Col}(M)=\{y\in\mathbb{R}^{m}\mid\exists x\in\mathbb{R}^{n}:\;y=Mx\}$ and $\text{Row}(M)=\{y\in\mathbb{R}^{n}\mid\exists x\in\mathbb{R}^{m}:\;y=M^{\top}x\}$.
+Then, the set of all possible $\Delta W$, say $\mathcal{W}_{LoRA-XS}$, is given as: $\mathcal{W}_{LoRA-XS}=\{M\in\mathbb{R}^{m\times n}|\text{Col}(M)\subseteq\text{Col}(B)\wedge\text{Row}(M)\subseteq\text{Row}(A)\},$ where $\text{Col}(M)$ and $\text{Row}(M)$ are column and row spaces of matrix $M$ respectively.
 Proof.
 Since $\Delta W=BRA$, we have
-$\text{Col}(\Delta W)=\{y\in\mathbb{R}^{m}\mid y=BRAx,\,x\in\mathbb{R}^{n}\}\implies\text{Col}(\Delta W)=\{y\in\mathbb{R}^{m}\mid y=Bz,\,z\in\text{Col}(RA)\}\subseteq\text{Col}(B).$
+$\text{Col}(\Delta W)=\{y\in\mathbb{R}^{m}\mid y=BRAx,\,x\in\mathbb{R}^{n}\}\implies$
+$\text{Col}(\Delta W)=\{y\in\mathbb{R}^{m}\mid y=Bz,\,z\in\text{Col}(RA)\}\subseteq\text{Col}(B).$
 That is, we proved that
@@
 Following similar arguments, one can also show $\text{Row}(\Delta W)\subseteq\text{Row}(A)$. ∎
-A.2 Proof of Lemma 2
+B.2 Proof of Lemma 2
 Lemma.
-The gradient of the loss with respect to matrix $R$ can be expressed in terms of the gradient with respect to the weight matrix $W$ as: $g^{R}_{LoRA-XS}=sB^{\top}gA^{\top}$
+The gradient of the loss with respect to matrix $R$ can be expressed in terms of the gradient with respect to the weight matrix $W$ as: $g^{R}_{LoRA-XS}=sB^{\top}gA^{\top}.$
 Proof.
@@
 Let $sB^{\top}g=y$. We know that when $X=RA$:
-$y\frac{\partial X}{\partial R}=yA^{\top}\implies\frac{\partial L}{\partial R}=yA^{\top}=sB^{\top}gA^{\top}$ (22)
+$y\frac{\partial X}{\partial R}=yA^{\top}\implies\frac{\partial L}{\partial R}=yA^{\top}=sB^{\top}gA^{\top}\$(17)
+$\text{Therefore, \quad}\boxed{g^{R}_{\text{LoRA-XS}}=sB^{\top}gA^{\top}}$ (18)
+B.3 Proof of Theorem 3
+Theorem.
+For full-rank AA and $B$ matrices, the optimal solution for the objective
 […]
-A.3 Proof of Theorem 3
-Theorem.
-The optimal solution for the objective $\text{min}_{g^{R}}||\tilde{g}-g||^{2}_{F}$, such that $\tilde{g}=sBg^{R}A$, is: $g^{R}=\dfrac{1}{s^{2}}(B^{\top}B)^{-1}g^{R}_{LoRA-XS}(AA^{\top})^{-1}$ (24)
 Proof.
@@
 $(B^{\top}B)^{-1}(B^{\top}Bg^{R}AA^{\top})(AA^{\top})^{-1}$ $=(B^{\top}B)^{-1}g^{R}_{\text{LoRA-XS}}(AA^{\top})^{-1}/s^{2}$ (23)
-[…]
-A.4 Proof of Theorem 4
+$\text{Therefore,}\quad\boxed{g^{R}=\dfrac{1}{s^{2}}(B^{\top}B)^{-1}g^{R}_{\text{LoRA-XS}}(AA^{\top})^{-1}}$ (24)
+B.4 Proof of Theorem 4
 Theorem.
 Consider the update for matrix $R$ using the solution derived in Theorem 3: $R\leftarrow R-\eta g^{R}$ where $\eta>0$ is the (sufficiently small) learning rate.
-This update guarantees a reduction in the loss $\Delta L$, given by: $\Delta L\coloneqq L(W_{0}+sB(R-\eta g^{R})A)-L(W_{0}+sBRA)=-\eta\langle g^{R}_{LoRA-XS},g^{R}\rangle_{F}+o(\eta)\leq 0$
+This update guarantees a reduction in the loss $\Delta L$, given by: $\Delta L\coloneqq L(W_{0}+sB(R-\eta g^{R})A)-L(W_{0}+sBRA)=-\eta\langle g^{R}_{LoRA-XS},g^{R}\rangle_{F}+o(\eta)\leq 0.$
 Proof.
@@
 Next, we note that matrices $B^{\top}B\in\mathbb{R}^{r\times r}$ and $AA^{\top}\in\mathbb{R}^{r\times r}$ are positive definite since they are positive semi-definite and matrices $B$ and $A$ are full-rank (i.e., with rank $r$) matrices, which means that $B^{\top}B$ and $AA^{\top}$ have non-zero eigenvalues.
-Therefore, $(B^{\top}B)^{-1}$ and $(AA^{\top})^{-1}$ are also positive definite, implying that there exist matrices $U$ and $V$ such that $(B^{\top}B)^{-1}=VV^{\top}$ and $(AA^{\top})^{-1}=UU^{\top}$ (e.g., one can find such matrices using Cholesky decomposition).
+Therefore, $(B^{\top}B)^{-1}$ and $(AA^{\top})^{-1}$ are also positive definite, implying that there exist matrices $X$ and $Y$ such that $(B^{\top}B)^{-1}=YY^{\top}$ and $(AA^{\top})^{-1}=XX^{\top}$ (e.g., one can find such matrices using Cholesky decomposition).
 Then, we have
-$\langle g^{R}_{\text{LoRA-XS}},(B^{\top}B)^{-1}g^{R}_{\text{LoRA-XS}}(AA^{\top})^{-1}\rangle_{F}$ $=\langle g^{R}_{\text{LoRA-XS}},VV^{\top}g^{R}_{\text{LoRA-XS}}UU^{\top}\rangle_{F}$
-$=\frac{1}{s^{2}}\langle V^{\top}g^{R}_{\text{LoRA-XS}}U,V^{\top}g^{R}_{\text{LoRA-XS}}U\rangle_{F}$
-$=\|V^{\top}g^{R}_{\text{LoRA-XS}}U\|_{F}^{2}\geq 0.$
+$\langle g^{R}_{\text{LoRA-XS}},(B^{\top}B)^{-1}g^{R}_{\text{LoRA-XS}}(AA^{\top})^{-1}\rangle_{F}$ $=\langle g^{R}_{\text{LoRA-XS}},YY^{\top}g^{R}_{\text{LoRA-XS}}XX^{\top}\rangle_{F}$
+$=\langle Y^{\top}g^{R}_{\text{LoRA-XS}}X,Y^{\top}g^{R}_{\text{LoRA-XS}}X\rangle_{F}$
+$=\|Y^{\top}g^{R}_{\text{LoRA-XS}}X\|_{F}^{2}\geq 0.$
 This concludes the proof. ∎
@@
 $\Delta L$ $=-\eta\langle g^{R}_{\text{LoRA-XS}},g^{R}_{\text{LoRA-XS}}\rangle_{F}+o(\eta)\leq 0.$ (27)
-A.5 Proof of Theorem 5
+B.5 Proof of Theorem 5
 Theorem.
-The equivalent gradient $\tilde{g}$ is hyperparameter $s$ independent when $\tilde{g}=sBg^{R}A\quad\text{ but not when }\quad\tilde{g}=sBg^{R}_{LoRA-XS}A$
+The equivalent gradient $\tilde{g}$ is hyperparameter $s$ independent when $\tilde{g}=sBg^{R}A\quad\text{ but not when }\quad\tilde{g}=sBg^{R}_{LoRA-XS}A.$
 Proof.
@@
 which is not $s$-independent. ∎
-A.6 Proof of Theorem 6
+B.6 Proof of Theorem 6
 Theorem.
-If $A_{init}$ and $B_{init}$ are initialized using LoRA-SB for the first step of SGD optimizer, then $\Delta(B_{init}R_{init}A_{init})\approx\Delta W$
+If $A_{init}$ and $B_{init}$ are initialized using LoRA-SB for the first step of SGD optimizer, then the update given by LoRA-SB, $\Delta(B_{init}R_{init}A_{init})$ , is the best low-rank approximation of full fine-tuning update, $\Delta W.$
 Proof.
@@
 Since we also get an identical expression, our solution is optimal. ∎
-Appendix B Inference Overhead vs LoRA
-As discussed in Section 5, LoRA-SB introduces minimal inference cost overhead due to the insertion of the $r\times r$ matrix $R$ between $B$ and $A$, as well as the requirement for higher ranks to match LoRA’s performance.
+Appendix C Simulating the First Step of Full Fine-Tuning Under AdamW
+Our initialization is designed to approximate the first update step that would occur during full fine-tuning using the AdamW optimizer, which is also used in LoRA-SB training.
+AdamW computes the parameter update using both first and second moment estimates of the gradient.
+At the first step, these moments are initialized to zero, so the update becomes:
+$\theta_{1}=\theta_{0}-\alpha\cdot\frac{g_{1}}{\sqrt{g_{1}^{2}+\epsilon}}\approx-\alpha\cdot\text{sign}(g_{1})$
+where $g_{1}$ is the gradient at the first step, $\epsilon$ is a small constant for numerical stability, and $\alpha$ is the learning rate.
+Due to zero-initialization and bias correction, the direction of the update is approximately the element-wise sign of the gradient.
+To simulate this behavior in our low-rank initialization, we use:
+$\Delta W_{\text{avg}}=-\eta\cdot\text{sign}\left(\sum_{i=1}^{n}\nabla_{W}\mathcal{L}(W_{0},x_{i})\right)$
+This reflects the direction of the first AdamW step averaged over a mini-batch.
+By using the sign of the gradient sum, we ensure our initialization aligns with the dynamics of AdamW, leading to a consistent and faithful approximation of full fine-tuning updates within the low-rank subspace.
+Appendix D Algorithm
+We provide a pseudo-code implementation of our method in Algorithm 1.
+Algorithm 1 LoRA‑SB, PyTorch‑like
+1:def initSB(model, D)
+2:   # Estimate gradient with n samples
+3:   $\Delta W_{\mathrm{avg}}\leftarrow$ est_grad(model, D, n)
+4:   # Initialize B, R, A
+5:   $(B,R,A)\leftarrow$ trunc_SVD($\Delta W_{\mathrm{avg}}$)
+6:   # Convert to LoRA‑SB model
+7:   sb_model $\leftarrow$ lora_SB(model, B, R, A)
+8:    return sb_model
+9:
+10:# Load pre‑trained model
+11: model $\leftarrow$ AutoModel(base_model)
+12:# Initialize LoRA‑SB with D
+13:sb_model $\leftarrow$ initSB(model, D)
+14:# Train, only R trainable
+15:trainer $\leftarrow$ Trainer(sb_model,…)
+16:trainer.train()
+Appendix E Optimal Gradient Approximation is Important!
+As discussed in Section 4, optimal gradient approximation plays a key role in the effectiveness of LoRA-SB.
+In Figure 3, we compare the loss curves of models trained with and without this component on Mistral-7B.
+While both variants begin with similar performance due to effective initialization, LoRA-SB with optimal gradient approximation converges to substantially lower loss values, highlighting its contribution to improved optimization.
+Figure 3: Training loss for Mistral-7B, highlighting the impact of optimal gradient approximation.
+Appendix F Training Time Overhead vs LoRA-XS
+As previously mentioned, we compute the update approximation using only $1/1000$ of the total training samples for each dataset.
+Table 6 presents the associated training time overhead for these computations, compared to LoRA-XS.
+The results show that the additional overhead is negligible, adding just 242–4 minutes compared to the total training time of 353–5 hours per epoch ($\approx 1.1\%$ to $1.3\%$).
+Additionally, the update computation is performed only once, at the beginning of the first epoch, prior to training.
+Notably, the initialization step is highly efficient, as we directly compute the truncated SVD using optimized PyTorch libraries (torch.svd_lowrank).
+For reference, this computation takes less than one second for each of the entire LLMs used in our experiments.
+Table 6: Training time overhead due to the initialization for various models on their respective tasks.
+Model Overhead Training Time/Epoch
+Mistral-7B 0:02:01 3:03:57
+Gemma-2 9B 0:03:46 4:13:24
+Llama-3.2 3B 0:03:54 4:54:31
+Appendix G Inference Overhead vs LoRA
+LoRA-SB introduces a minimal inference cost overhead due to the insertion of the $r\times r$ matrix $R$ between $B$ and $A$, and the need for higher ranks to achieve comparable performance to LoRA.
+We benchmark the inference-time FLOPs and MACs across various models and find that the overhead is negligible.
 This comparison is presented in Table 7, showing that the additional overhead of LoRA-SB is negligible.
@@
 Model Method Rank MACs FLOPs
-RoBERTa-large LoRA $8$ $77.86$ G $155.79$ G
+RoBERTa-large LoRA 88 $77.86$ G $155.79$ G
 LoRA-SB $16$ $78.42$ G $156.91$ G
@@
 LoRA-SB $96$ $3.97$ T $7.94$ T
-Appendix C Experiment Details
-We ran our experiments on a single NVIDIA A6000 GPU, averaging results over three independent trials. We trained all models using the AdamW optimizer (Loshchilov & Hutter, 2019).
+Appendix H Experiment Details
+We use PyTorch (33) and the HuggingFace Transformers library (48) for our implementations.
+We run all experiments on a single NVIDIA A6000 GPU and report results as the average of three random seeds.
+To save memory, we initialize base models in torch.bfloat16 precision.
+We trained all models using the AdamW optimizer (28).
+We compute the update approximation using only $\mathbf{1/1000}$ of each dataset’s total number of samples.
+The samples are randomly selected from the training set in each run.
 For arithmetic and commonsense reasoning tasks, we set up Mistral-7B, Gemma-2 9B, and Llama-3.2 3B with hyperparameters and configurations listed in Table 8.
-We adopted most settings from previous studies (Hu et al., 2023) but conducted our own learning rate sweep.
+We adopted most settings from previous studies (18) but conducted our own learning rate sweep.
 Following LoRA-XS guidelines, we set $\alpha=r$ for their baseline configuration.
 For the GLUE benchmark using RoBERTa-large, you can find the hyperparameter details in Table 9.
-We mostly adhered to the original configurations from the LoRA paper (Hu et al., 2021) but adjusted the learning rate through a sweep.
+We mostly adhered to the original configurations from the LoRA paper (17) but adjusted the learning rate through a sweep.
 In line with LoRA-XS settings, we fixed $\alpha$ at $16$ for their baseline.
-For all tasks, we followed the baseline configurations provided in the PiSSA (Meng et al., 2024) and rsLoRA (Kalajdzievski, 2023) papers for our comparisons.
+For all tasks, we followed the baseline configurations provided in the PiSSA (30), rsLoRA (20), DoRA (26), and LoRA-Pro (46) papers for our comparisons.
 Table 8: Hyperparameter settings for training Mistral-7B and Gemma-2 9B on MetaMathQA, and Llama-3.2 3B on Commonsense170K.
@@
 Optimizer AdamW AdamW
-Batch size $1$ $6$
+Batch size 11 66
 Max. Seq. Len $512$ $256$
 Grad Acc. Steps $32$ $24$
-Epochs $1$ $2$
-Dropout 00 $0.05$
+Epochs 11 22
+Dropout 0 $0.05$
 Learning Rate $1\times 10^{-4}$ $2\times 10^{-3}$
@@
 Epochs $30$ $30$ $30$ $15$ $15$ $30$
-Dropout 00
+Dropout 0
 Learning Rate $1\times 10^{-3}$
@@
 Warmup Ratio $0.06$
-Appendix D Dataset Details
-The MetaMathQA dataset (Yu et al., 2024) creates mathematical questions by rephrasing existing ones from different viewpoints, without adding new information.
-We assess this dataset using two benchmarks: GSM8K (Cobbe et al., 2021), which consists of grade-school math problems requiring multi-step reasoning, and MATH (Hendrycks et al., 2021), which presents difficult, competition-level math problems.
+Appendix I Dataset Details
+The MetaMathQA dataset (50) creates mathematical questions by rephrasing existing ones from different viewpoints, without adding new information.
+We assess this dataset using two benchmarks: GSM8K (8), which consists of grade-school math problems requiring multi-step reasoning, and MATH (16), which presents difficult, competition-level math problems.
 Evaluation focuses solely on the final numeric answer.
-CommonSense170K is a comprehensive dataset that consolidates eight commonsense reasoning datasets (Hu et al., 2023).
+CommonSense170K is a comprehensive dataset that consolidates eight commonsense reasoning datasets (18).
 Each example is framed as a multiple-choice question where the model generates the correct answer without explanations.
-We use the prompt template from (Hu et al., 2023).
+We use the prompt template from (18).
 The individual datasets used are described below:
   1. 1.
-    HellaSwag (Zellers et al., 2019) challenges models to select the most plausible continuation of a given scenario from multiple possible endings.
+    HellaSwag (51) challenges models to select the most plausible continuation of a given scenario from multiple possible endings.
   2. 2.
-    ARC Easy (or ARC-e) (Clark et al., 2018) includes basic science questions at a grade-school level, offering simpler tasks to assess fundamental reasoning abilities.
+    ARC Easy (or ARC-e) (7) includes basic science questions at a grade-school level, offering simpler tasks to assess fundamental reasoning abilities.
   3. 3.
-    PIQA (Bisk et al., 2020) evaluates physical commonsense reasoning, where models must choose the best action to take in a hypothetical scenario.
+    PIQA (3) evaluates physical commonsense reasoning, where models must choose the best action to take in a hypothetical scenario.
   4. 4.
-    SIQA (Sap et al., 2019) tests social commonsense reasoning by asking models to predict the social consequences of human actions.
+    SIQA (39) tests social commonsense reasoning by asking models to predict the social consequences of human actions.
   5. 5.
-    WinoGrande (Sakaguchi et al., 2021) presents sentence completion tasks requiring commonsense reasoning to select the correct binary option.
+    WinoGrande (38) presents sentence completion tasks requiring commonsense reasoning to select the correct binary option.
   6. 6.
-    ARC Challenge (or ARC-c) (Clark et al., 2018) consists of more complex science questions designed to challenge models with sophisticated reasoning, beyond simple co-occurrence patterns.
+    ARC Challenge (or ARC-c) (7) consists of more complex science questions designed to challenge models with sophisticated reasoning, beyond simple co-occurrence patterns.
   7. 7.
-    OBQA (Mihaylov et al., 2018) features open-book, knowledge-intensive QA tasks that require multi-hop reasoning across multiple information sources.
+    OBQA (31) features open-book, knowledge-intensive QA tasks that require multi-hop reasoning across multiple information sources.
   8. 8.
-    BoolQ (Clark et al., 2019) involves answering yes/no questions based on real-world, naturally occurring queries.
+    BoolQ (6) involves answering yes/no questions based on real-world, naturally occurring queries.
 The GLUE Benchmark is a comprehensive collection of tasks designed to evaluate natural language understanding (NLU) abilities.
-It included various datasets, including STS-B for measuring semantic textual similarity (Cer et al., 2017), RTE for recognizing textual entailment, MRPC for detecting paraphrases (Dolan & Brockett, 2005), CoLA for assessing linguistic acceptability (Warstadt et al., 2019), SST-2 for sentiment analysis (Socher et al., 2013), and QNLI for question-answer inference (Rajpurkar et al., 2018).
+It included various datasets, including STS-B for measuring semantic textual similarity (5), RTE for recognizing textual entailment, MRPC for detecting paraphrases (11), CoLA for assessing linguistic acceptability (47), SST-2 for sentiment analysis (41), and QNLI for question-answer inference (36).
 GLUE’s broad scope makes it a standard benchmark for evaluating models like RoBERTa.
+Appendix J Use of Large Language Models
+LLMs are only used for small writing improvements, like polishing grammar and smoothing out phrasing.